### v1

# Identifying Reliable Predictions in Detection Transformers

Abstract
DEtection TRansformer (DETR) has emerged as a promising architecture for object detection, offering an end-to-end prediction pipeline. In practice, however, DETR generates hundreds of predictions that far outnumber the actual number of objects present in an image. This raises the question: can we trust and use all of these predictions? Addressing this concern, we present empirical evidence highlighting how different predictions within the same image play distinct roles, resulting in varying reliability levels across those predictions. More specifically, while multiple predictions are often made for a single object, our findings show that most often one such prediction is well-calibrated, and the others are poorly calibrated. Based on these insights, we demonstrate identifying a reliable subset of DETR’s predictions is crucial for accurately assessing the reliability of the model at both object and image levels.
Building on this viewpoint, we first tackle the shortcomings of widely used performance and calibration metrics, such as average precision and various forms of expected calibration error. Specifically, they are inadequate for determining which subset of DETR’s predictions should be trusted and utilized. In response, we present Object-level Calibration Error (OCE), which is capable of assessing the calibration quality both across different models and among various configurations within a specific model. As a final contribution, we introduce a post hoc Uncertainty Quantification (UQ) framework that predicts the accuracy of the model on a per-image basis. By contrasting the average confidence scores of positive (i.e., likely to be matched) and negative predictions determined by OCE, the framework assesses the reliability of the DETR model for each test image. The code is available at https://github.com/azizanlab/reliable-detr.
(a) Low confidence with random boxes
(b) Calibrated confidence with similar boxes
(c) Uncalibrated confidence with similar boxes
Figure 1: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with one being well-calibrated. This figure illustrates how DETR can handle the remaining predictions. In principle, DETR may (a) generate low confidence with random bounding boxes, (b) assign equally calibrated confidence with similar boxes, or (c) poorly calibrated confidence with similar boxes. Our analysis indicates that DETR mostly follows the third scheme; resulting in varying levels of reliability across predictions.
1 Introduction
Object detection is an essential task in computer vision, with applications that span various domains including autonomous driving, warehousing, and medical image analysis. Existing object detection methods predominantly utilize Convolutional Neural Networks (CNNs) [8, 23, 22, 9] to identify and locate objects within images. However, the recent introduction of DEtection TRansformer (DETR) [1] has revolutionized the field by offering an end-to-end prediction pipeline where the model predicts a set of bounding boxes and class probabilities.
The core innovation of DETR lies in the use of a Transformer encoder-decoder architecture, enabling the model to generate predictions in an end-to-end manner and enhancing scalability. This paradigm shift has led to the exploration of various DETR variants—such as Deformable-DETR [32] and DINO [31]—positioning them as potential foundation models for object detection tasks. Despite these advancements, the inner workings of how these predictions are generated and interact within the Transformer decoder layers remain under-explored.
Likewise, there is an ongoing debate within the community about whether DETR truly provides an end-to-end prediction pipeline. The central concern is whether all the hundreds of predictions generated by DETR can be trusted and used. By means of an answer, practitioners often employ heuristic approaches to achieve high precision, such as by setting a user-defined threshold to retain only a small subset of high-confidence outputs, as seen in the official demo. Previous studies exploring model reliability, such as [10, 15, 16, 21, 17], have also applied a user-defined confidence threshold (e.g., 0.3) to retain a subset of predictions for evaluating calibration quality, rather than using the entire set. On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR and 300 out of 900 for DINO) based on confidence scores during post-processing to achieve a high average precision (AP) score. However, the significance of selecting a subset of the predictions as well as the choice of a suitable configuration for doing so remain under-explored. We provide a more in-depth discussion about related works in Appendix A.
This paper examines these issues by addressing the following three research questions: (RQ1) Do all predictions generated for a given image exhibit comparable levels of reliability? (RQ2) If not, what is the appropriate way to identify reliable predictions across the entire set? and (RQ3) How can we accurately assess DETR’s reliability?
Our primary findings reveal that predictions from DETR within the same image are interdependent, leading to varying levels of reliability. Since DETR is trained using gradients from optimally matched predictions, it has flexibility in handling the remaining predictions. In principle, DETR could assign them low confidence with random bounding boxes (Figure 1(a)). Alternatively, it could also generate similar bounding boxes with equally calibrated confidence (Figure 1(b)). Otherwise, it may assign poorly calibrated confidence scores (Figure 1(c)). Our analysis indicates that DETR mostly follows the third scheme; we refer to these well-calibrated predictions as positive predictions and the others as negative predictions.
Building on this observation, this paper investigates the importance of distinguishing between positive and negative predictions, and how this finding can be leveraged into an uncertainty quantification (UQ) framework for DETR. More specifically, our main contributions are:
  • We offer both qualitative and quantitative insights into how positive and negative predictions influence the model’s object-level calibration. Our analysis shows that negative predictions are poorly calibrated. Furthermore, their confidence scores are inversely correlated with image-level reliability, highlighting the need for a separation method to ensure DETR’s reliable use.
  • We raise concerns about existing performance and calibration metrics—including AP [25, 6, 14], detection expected calibration error (D-ECE) [10], and Localisation-aware expected calibration error (LA-ECE) [19, 11]—by noting their inadequacy for identifying the reliable subset of predictions. For instance, AP favors retaining more predictions, while ECE prefers a smaller high-confidence set. Consequently, the subset of predictions that maximizes performance for each metric differs significantly from the optimal positive set.
  • To address this, we introduce a new calibration metric, object-level calibration error (OCE), which calculates calibration error along ground truth objects rather than predictions. We demonstrate that OCE is capable of both ranking different models by their calibration qualities and identifying the reliable subset of predictions to use.
  • Based on our findings, we provide a practical implication: we introduce a novel framework for quantifying image-level reliability by contrasting the average confidence scores between the positive and negative predictions identified by our OCE metric. We conduct numerical experiments, demonstrating the effectiveness of our approach across in-distribution, near out-of-distribution, and far out-of-distribution scenarios.
  • We present a comparative analysis of methods for identifying positive predictions in DETR, from the UQ perspective. Our results show that thresholding (on predictive confidence score) is a more suitable method than widely used techniques such as top-$k$ or non-maximum suppression (NMS) for obtaining a reliable subset of predictions.
2 Problem Statement
2.1 Detection Transformers (DETRs)
We consider a test image $\bm{x}$ and denote $\mathcal{T}_{\bm{x}}$ as the set of ground truth objects present in the image. Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{T}}_{\theta}(\bm{x})$. Each prediction $t\in\hat{\mathcal{T}}_{\theta}(\bm{x})$ is characterized by a bounding box $\hat{\bm{b}}$ and an associated class label with a corresponding probability $\hat{p}$.
The structure of DETR is composed of two main components: the Transformer encoder, which extracts a collection of features from the given image; and the Transformer decoder, which uses these features to make predictions. In addition to the features extracted by the encoder, the decoder’s input consists of $M$ (typically several hundred) learnable embeddings, also known as object queries. Each decoder layer is composed of a self-attention module among object queries and a cross-attention module between each object query and the features. After processing the queries through several decoder layers, the model produces the $M$ final representation vectors that are converted into bounding boxes and class labels via a shared feedforward network, $f_{\phi}$. Together, these predictions form the final outputs, making DETR’s predictions essentially an $M$-element set. We refer to Figure 4 in the Appendix for an illustration.
The encoder follows the common structure of standard computer vision models and is based on pre-trained models, whose reliability has been relatively widely explored [27, 26, 29, 20]. This foundation further enables the use of prominent post hoc UQ techniques, such as Monte Carlo dropout [7, 29] and distance-based out-of-distribution (OOD) detection methods [12, 28, 4]. However, despite the decoder being the predominant component for object detection, there is a gap in understanding and quantifying its reliability due to its unique structural characteristic: set prediction. Therefore, this paper delves into the underlying characteristics of these predictions and presents a methodology to quantify the reliability of the decoder in DETR for object detection tasks.
(a) A prediction in high-reliability image: Horse
(b) A prediction in low-reliability image: Chair
Figure 2: Visualizations of the predictions generated by Cal-DETR. The optimal positive prediction (indexed by 0 and bordered in green) and the five optimal negative predictions (indexed by 1-5 and bordered in red) with the largest IoU are presented. For each prediction and layer, the maximum confidence score and its corresponding label are visualized. When the model is confident, the confidence score of the positive prediction either increases or remains high across the decoder layers, while those of negative predictions decrease or stay low. On the other hand, when the model is uncertain, DETR assigns a confidence score to positive prediction based on its confidence, thereby maintaining good calibration. However, conversely, it slightly increases or maintains the confidence scores for negative predictions.
2.2 Identifying Reliable Predictions in DETR
Notably, the number of predictions generated by DETR, $|\widetilde{\mathcal{T}}_{\theta}(\bm{x})|$, is fixed and often in the hundreds, far exceeding the number of ground truth objects. To address the issue, during model training, a bipartite matching algorithm is used to find the optimal matching prediction for each ground truth object based on the alignment of the class label and bounding box (as detailed in Appendix B). Consequently, the parameters $\theta$ are optimized to enhance the accuracy of these matched queries. In this paper, we refer to the matched queries as optimal positive queries, while the remaining queries are termed optimal negative queries.
At the inference stage in real-world scenarios, however, ground truth annotations are unavailable, meaning the optimal positive queries remain unknown. This raises the question of whether all predictions can be trusted and used without any post-processing, or if only a subset should be selected for the final inference—and, if so, which subset? If all of the model’s predictions, including negative queries, were generated independently and were nearly identical to one of the positive predictions, the large number of predictions would not be a concern. We could simply apply well-known algorithms like NMS to remove duplicates and resolve redundancy. Notably, we have observed that DETR assigns well-calibrated confidence scores to only a single positive query (i.e., prediction) per object, while assigning low scores but similar bounding boxes to negative queries (see Section 3.1 and Figure 2). As a result, the overall prediction-level calibration error becomes significantly high (See Section 3.2) if we include those noisy predictions in the final inference results. Hence, improper handling of negative queries (e.g., by using top-100 predictions) can severely compromise DETR’s reliability, with potentially catastrophic consequences in safety-critical scenarios, such as pedestrian detection in autonomous vehicles. For this reason, addressing this challenge and effectively distinguishing the positives from negatives is a crucial task for ensuring reliable UQ in DETR.
3 Analysis and Key Insights
3.1 Exploring the Anatomy of DETR’s Predictions
We begin by closely examining the underlying dynamics of DETR decoders and visualizing the generated outputs. Since the Transformer decoder outputs only representation vectors, investigating their evolution across layers is not straightforward. We address this by reapplying the final feedforward network that operates on the last layer, $f_{\phi}$, to the intermediate layers. This allows us to transform each representation vector into its associated bounding box and class label. This is feasible due to the alignment of intermediate representations, facilitated by residual connections between decoder layers [2]. Sample visualizations are in Figure 2 along with Figures 5 and 6 in Appendix C.
In the first decoder layer, the model appears to explore the encoded image features, producing varied queries that result in various plausible predictions. In this early stage, the distinction between positive and negative queries can be ambiguous (e.g., Figure 2(a)). However, the self-attentions through the subsequent decoder layers progressively refine these predictions. By the final layer, the model selects a single query (i.e., optimal positive) and assigns a confidence score based on its understanding of the image and the object. In contrast, the confidence scores for neighboring queries (i.e., optimal negatives) do not increase to the same extent as the positives and even decrease in high-reliability images. In contrast, in low-reliability images (e.g., Figure 2(b)), the confidence score of the positive query does not significantly increase, while the scores of the negative ones are either slightly raised or unchanged. Based on this observation, we present our claim:
Claim 1.
Predictions from DETR within a given image exhibit varying levels of reliability. For each object in the image, the optimal positive (i.e., best-matched) prediction is calibrated, while the remaining optimal negative predictions are uncalibrated.
3.2 Analysis of Calibration Error for Positive and Negative Predictions
Primarily, to address the first research question (RQ1) and provide quantitative support for our claim regarding the varying levels of reliability across predictions, we evaluate and compare the calibration errors on the optimal positive predictions and the optimal negative predictions. The optimal positives and negatives are determined using the standard bipartite matching algorithm [1].
We evaluate the object-level calibration error, which is defined later in Section 4.2. From Table 1, we confirm our claim that only the optimal positive predictions are properly-calibrated. Furthermore, as illustrated in Figure 2, the optimal negative predictions exhibit a significantly high calibration error, driven by consistently low confidence scores regardless of their actual accuracy. It is also noteworthy that even Cal-DETR, the state-of-the-art calibrated DETR variant, demonstrates poor calibration quality for optimal negative predictions.
Moreover, as will be explained in more detail later, Table 3 further supports our claim by illustrating that distinguishing between positive and negative predictions becomes particularly crucial when applied to image-level UQ in DETR. In short, when confidence scores are averaged without distinguishing positive queries (see the row of $\mathsf{Conf}^{+}$ using the top-$100$ predictions), the correlation between the confidence scores and image-level reliability (i.e., precision) becomes negative. This is because the average confidence scores for negative queries are shown to be inversely correlated with image-level reliability, and the majority of DETR’s outputs are actually negative. These results further highlight the importance of employing an appropriate separation scheme to accurately identify optimal positive predictions, even when ground truth annotations are unavailable.
Table 1: Object-level calibration error on top-$100$, optimal positive, and negative predictions are reported across various models and datasets. For each model, the smallest and largest errors are highlighted in blue and red, respectively. Notably, negative queries are uncalibrated. Since the top-$100$ predictions often include these negative queries, the resulting calibration error tends to be high.
Model Predictions COCO City Foggy
D-DETR Top-100 0.582 0.552 0.618
Negative 0.803 0.789 0.844
Positive 0.275 0.341 0.419
UP-DETR Top-100 0.660 0.689 0.709
Negative 0.872 0.871 0.883
Positive 0.438 0.494 0.546
Cal-DETR Top-100 0.655 0.668 0.688
Negative 0.867 0.859 0.871
Positive 0.425 0.463 0.513
DINO Top-100 0.635 0.693 0.713
Negative 0.953 0.959 0.962
Positive 0.411 0.480 0.544
4 A Systematic Framework for Identifying Positive Predictions
Identifying the positive predictions is crucial for the reliable use of DETR. Nonetheless, identifying optimal positive predictions during the inference stage is not feasible due to the lack of ground truth labels. Therefore, an alternative systematic framework is essential not only for improving its interpretability but also for ensuring reliability. Given that the optimal matching achieves the smallest calibration error, we use calibration error to measure the quality of a separation scheme, i.e., a separation scheme with a lower calibration error is deemed a better scheme.
(a) Cal-DETR
(b) UP-DETR
Figure 3: Impact of confidence threshold selection on various performance metrics in Cal-DETR and UP-DETR on COCO.
4.1 Limitations of Existing Metrics
To this end, this section analyzes the effectiveness of existing metrics, including AP, D-ECE, LA-ECE0, and LRP, for measuring calibration error. Specifically, we modify a variable (e.g., the confidence score threshold) to generate different subsets of DETR predictions. We then assess the performance of each metric for these subsets and determine the variable value that yields the highest performance. Lastly, we evaluate the extent of separation by comparing the predicted positive set to the optimal positive set. Sample analysis results are illustrated in Figure 3. For additional results and detailed information on these metrics, please refer to Appendix D.
Primarily, as noted in several studies [18, 19, 11] and our empirical findings, the optimal AP is achieved when the threshold is set to $0.0$; AP does not penalize harshly for having low confidence predictions. However, as discussed earlier, using the entire prediction set carries a high risk of including uncalibrated negatives, leading to unreliable decisions in practical applications. Furthermore, using hundreds of predictions harms the interpretability of the model.
In contrast, the optimal ECEs are often achieved when the threshold is set close to $1.0$, meaning ECEs favor retaining fewer predictions with high confidence. This is a structural pitfall that prediction-level ECEs commonly face [11]. Since ECEs do not penalize missed detections (i.e., false negatives), they can achieve near-zero error when the evaluation positive set consists solely of highly accurate and confident predictions. Therefore, unless the model is trained to be excessively overconfident—which is unlikely given that DETR is trained on large datasets using various auxiliary loss functions—ECEs can result in very small error values when using a large confidence threshold.
LRP [18], a localization-aware performance metric, can be used instead. However, LRP is not designed as a calibration metric and does not explicitly consider calibration error. As a result, the ranking scored by LRP across the different models is not necessarily aligned with the model’s calibration qualities. See Section 6.1 for the numerical analysis and more discussions.
4.2 Proposed Metric: Object-Level Calibration Error
Notation.
For each image $\bm{x}_{i}$ having $N_{i}$ objects, we consider the ground truth set $\mathcal{T}_{\bm{x}_{i}}=\{o_{i,j}=(l_{i,j},\bm{b}_{i,j})\}_{j=1}^{N_{i}bu}$, and a set of DETR predictions: $\hat{\mathcal{T}}_{\theta}(\bm{x}_{i})=\{t_{i,q}=(\hat{\bm{p}}_{i,q},\hat{\bm{b}}_{i,q})\}_{q=1}^{M}$ where $M$ represents the number of object queries. Here, $l_{i,j}\in[1,C]$ and $\hat{\bm{p}}_{i,q}\in\{0,1\}^{C}$ represents the ground-truth label and predicted probability distribution over $C$ classes, respectively, while $\bm{b}_{i,j}$ and $\hat{\bm{b}}_{i,q}\in\{0,1\}^{4}$ corresponds to the ground truth and predicted scaled bounding box, respectively.
Definition 1.
Consider a subset of predictions, $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\subseteq\hat{\mathcal{T}}_{\theta}(\bm{x}_{i})$, that is generated by post-processing algorithm $\mathcal{S}$ from the entire prediction set: $\hat{\mathcal{S}}_{\theta}=\mathcal{S}\circ\hat{\mathcal{T}}_{\theta}$. We define an object-level calibration error (OCE) as the average Brier score per object:
$\mathsf{OCE}\big{(}\hat{\mathcal{S}}_{\theta};\mathcal{I}\big{)}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(i,j)\in\mathcal{I}}\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});o_{i,j}\big{)}$ (1)
$\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});o_{i,j}\big{)}=\sum_{c=1}^{C}\big{(}\mathbbm{1}(c=l_{i,j})-\widetilde{p}_{i,j}[c]\big{)}^{2}$ (2)
$\widetilde{p}_{i,j}[c]=\frac{1}{|\mathcal{Q}_{i,j}|}\sum_{q\in\mathcal{Q}_{i,j}}\hat{\bm{p}}_{i,q}[c]$ (3)
where $\mathcal{I}=\cup_{i}\{(i,j)\}_{j=1}^{N_{i}}$ is a set of all objects indices and $\hat{\bm{p}}(\cdot)[c]$ outputs the probability of $c$-th class; $\mathcal{Q}_{i,j}$ is a set of query indices that matches to the ground truth object $o_{i,j}$ and we propose two variants:
$\mathcal{Q}_{i,j}$$\triangleq\{q\mid\mathsf{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\geq\epsilon\}~{}$ $(\mathsf{OCE_{ENS\epsilon}})$ (4)
$\mathcal{Q}_{i,j}$$\triangleq\{q=\text{argmin}_{q}\mathsf{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\}~{}$ $(\mathsf{OCE_{MAX}})$ (5)
The difference is that $\mathsf{OCE_{ENS\epsilon}}$ ensembles the overlapping predictions, while $\mathsf{OCE_{MAX}}$ selects the prediction with the best bounding-box matching. When $|\mathcal{Q}_{i,j}|=0$, we consider the predicted probability to be zero, thus the corresponding Brier score is estimated as $1.0$. Following [14, 11], we use IoU thresholds of $\epsilon=0.5,0.75$ and report the average score as $\mathsf{OCE_{ENS}}$ or simply $\mathsf{OCE}$.
The introduced calibration error has two desirable characteristics. First, the prediction set achieves the lowest calibration error when the predictions are well-calibrated to the respective closest ground truth objects. Second, it penalizes the subset, $\hat{\mathcal{S}}_{\theta}(\cdot)$, that includes missing ground truth objects; this ensures that subsets containing a small set of highly precise predictions are not assigned an artificially low error, unlike D-ECE and LA-ECE metrics. Thus, this metric can effectively assess not only whether the given subsets $\{\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\}_{i}$ are reliable, but also whether they comprehensively capture all ground truth objects, providing richer information compare to existing metrics.
5 Use-Case Implications
5.1 Image-Level Reliability
As a main use-case implication of this research, this paper investigates how effective separation impacts the performance of quantifying image-level reliability within the DETR framework. Specifically, we first introduce a formal definition of image-level reliability by examining the model’s overall object detection performance on the image.
Definition 2.
We define image-level reliability as a measure of how accurately and confidently the predictions match the ground truth objects:
$\mathsf{ImReli}(\bm{x};\theta)\triangleq\mathsf{Perf}\big{(}\hat{\mathcal{T}}_{\theta}(\bm{x}),~{}\mathcal{T}_{\bm{x}}\big{)}.$ (6)
where one can quantify $\mathsf{Perf}$ using any standard performance metrics such as average precision, depending on the specification.
By its definition, image-level reliability directly addresses the model’s applicability to a given test instance. However, since image-level reliability requires ground truth annotations for its determination, obtaining it during inference is not feasible. Therefore, our goal is to employ a proxy method that quantifies uncertainty or confidence, which should closely correlate with image-level reliability.
5.2 Proposed Framework: Quantifying Reliability by Contrasting
As illustrated, positive and negative queries exhibit varying levels of reliability. Interestingly, having queries with low confidence scores does not necessarily imply low reliability. Our empirical observations show that confidence scores in negative queries actually inversely correlate with image-level reliability. More specifically, for a reliable instance, we observe that the confidence of positive queries increases across the decoder layers, while that of negative queries remains low; this results in a large gap between positive and negative queries. In contrast, for unreliable instances (e.g., Figure 2(a)), the confidence score of the positive query does not increase across the layers, whereas the scores of negative queries are either slightly elevated or remain unchanged. Consequently, there is a small gap between positive and negative queries (e.g., Figure 2(b)).
Based on the finding, we propose a post hoc UQ approach by contrasting the confidence scores of positives and negatives:
$\mathsf{ContrastiveConf}(\bm{x})$ $=\mathsf{Conf}^{+}(\bm{x})-\lambda\mathsf{Conf}^{-}(\bm{x})$ (7)
$\mathsf{Conf}^{+}(\bm{x})$ $=\frac{1}{|\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})|}\sum\nolimits_{t\in\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})}\mathsf{c}(t)$ (8)
$\mathsf{Conf}^{-}(\bm{x})$ $=\frac{1}{|\hat{\mathcal{T}}_{\theta}^{-}(\bm{x})|}\sum\nolimits_{t\in\hat{\mathcal{T}}_{\theta}^{-}(\bm{x})}\mathsf{c}(t)$ (9)
where $\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})$ and $\hat{\mathcal{T}}_{\theta}^{-}(\bm{x})$ are predicted sets of positive and negative predictions, respectively, and $\lambda$ is a scaling factor, that can be selected from the validation set. We include an ablation study of the scaling factor in the supplementary material. Our results show that the proposed method is not too sensitive to the choice of scaling factor and consistently outperforms baselines, with a scaling factor of $10.0$ providing adequate performance. $\mathsf{c}(\cdot)$ denotes the confidence estimate for the prediction; in this paper, we use maximum probability, $\mathsf{c}(t=(\hat{\bm{p}},\hat{\bm{b}}))=\max_{c}\hat{\bm{p}}[c]$.
As shown so far, identifying a $\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})$ from $\hat{\mathcal{T}}_{\theta}(\bm{x})$ is a crucial factor in the success of this approach. In practice, however, neither the ground truth (i.e., the queries that DETR intends to use) nor the optimal positives are available during the UQ stage. Instead, we approximate the ground truth separation by applying a post-processing algorithm that minimizes the calibration error on the validation set:
$\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})=\mathcal{S}^{*}\circ\hat{\mathcal{T}}_{\theta}(\bm{x})$ (10)
$\mathcal{S}^{*}=\text{argmin}_{\mathcal{S}}\mathsf{OCE}(\mathcal{S}\circ\hat{\mathcal{T}}_{\theta};\mathcal{I}_{val})$ (11)
where $\mathcal{I}_{val}=\cup_{i\in\mathcal{V}}\{(i,j)\}_{j=1}^{O_{i}}$ is the set of all objects indices in the validation dataset $\mathcal{V}$. Unless specified otherwise, we adopt confidence thresholding for the separation scheme (i.e., post-processing algorithm), where the corresponding threshold is the variable to optimize.
Table 2: Comparison of the effectiveness of the proposed OCE and baseline metrics, focusing on their ranking correlation with the rankings given by the models’ calibration quality on optimal positive predictions (denoted by the superscript *). For each metric, various post-processing schemes (i.e., ways to identify positive predictions) are applied and indicated within parentheses. Correlations exceeding $0.9$ are highlighted in blue, while negative correlations are highlighted in red. Our OCE-based rankings consistently provide strong correlations with optimal ranking and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD) Average
[…]
References D-ECE 1.000 0.995 0.969 1.000 0.996 0.993 1.000 0.873 0.996 0.980
[…]
$\mathsf{OCE_{ENS{}}}^{*}$ 0.969 0.948 1.000 0.993 0.990 1.000 0.996 0.836 1.000 0.970
Baselines AP (Top-100) [6, 14] 0.552 0.583 0.343 0.779 0.784 0.700 0.597 0.843 0.523 0.634
LRP (LRP) [18] 0.103 0.051 0.338 -0.268 -0.272 -0.150 -0.257 -0.610 -0.170 -0.137
D-ECE (Top-100) -0.873 -0.833 -0.966 0.608 0.576 0.511 -0.254 0.250 -0.316 -0.144
D-ECE (0.3) [10] 0.905 0.939 0.827 0.982 0.967 0.990 0.964 0.799 0.981 0.928
D-ECE (LRP) 0.802 0.855 0.659 0.987 0.984 0.999 0.983 0.793 0.996 0.895
D-ECE ($\mathsf{OCE_{ENS}}$) 0.856 0.901 0.727 0.987 0.983 0.999 0.983 0.793 0.996 0.914
LA-ECE0 (Top-100) -0.944 -0.913 -0.994 -0.431 -0.482 -0.517 -0.786 -0.386 -0.822 -0.697
LA-ECE0 (0.3) -0.789 -0.729 -0.842 -0.780 -0.818 -0.830 -0.451 -0.041 -0.468 -0.639
LA-ECE0 (LRP) [11] -0.908 -0.864 -0.948 -0.833 -0.877 -0.846 -0.901 -0.597 -0.913 -0.854
LA-ECE0 ($\mathsf{OCE_{ENS}}$) -0.901 -0.855 -0.943 -0.723 -0.780 -0.739 -0.884 -0.568 -0.897 -0.810
Ours $\mathsf{OCE_{ENS}}$ ($\mathsf{OCE_{ENS}}$) 0.962 0.940 1.000 0.987 0.983 0.999 0.964 0.737 0.984 0.951
$\mathsf{OCE_{MAX}}$ ($\mathsf{OCE_{MAX}}$) 0.954 0.929 0.998 0.954 0.958 0.983 0.955 0.697 0.976 0.934
6 Experiments
To demonstrate the effectiveness of the proposed method, we conducted experiments using four DETR variants: UP-DETR [3], Deformable-DETR (D-DETR), Cal-DETR [17], and DINO. Each model is trained on the COCO (train2017) dataset and we use $1,000$ images (i.e., $20\%$) of COCO (val2017) for the validation set and the remaining images for the test set. The model is tested on three datasets with varying levels of out-of-distribution (OOD) characteristics: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD). Experimental details can be found in Appendix E.
6.1 Effectiveness of OCE
To showcase OCE’s effectiveness in assessing calibration quality, we analyze how different metrics rank calibration error across four DETR models. As baselines, we measure ECEs using the top-$100$ detections (best AP), a standard confidence threshold of 0.3, and the threshold that minimizes LRP on the validation set. In a similar manner, we assess our OCEs using thresholds that minimize each OCE metric on the validation set. For fair comparison, we execute the optimal matching process on each test dataset to obtain the optimal positive set. We then evaluate D-ECE, LA-ECE0, and OCE using these sets as reference scores and calculate the Pearson correlation coefficient to the other scoring schemes across different datasets.
As shown in Table 2, our OCE metrics maintain a strong correlation with the optimal ECE metrics across different datasets. This evidence supports the reliability of our OCEs in ranking models by their overall calibration quality on positive sets—which downstream applications are more likely to utilize instead of the top 100 predictions. Notably, LA-ECE0 with a non-optimal positive set shows an inverse correlation with the rankings estimated on the optimal positive set. Meanwhile, from a model ranking perspective, empirical evidence suggests that solely adopting LRP does not align with calibration qualities. Nonetheless, the ranking by D-ECE and OCE scores measured at the optimal LRP threshold — denoted by D-ECE (LRP) and OCE (LRP), respectively — show a strong correlation to those ranked from the optimal positive set; thus, the method of using LRP to identify a positive set [11] appears acceptable. Experimentally, we confirmed that OCE and LRP provide thresholds in a similar range across different settings.
Another important attribute of OCE is that it can not only be used to compare different models but is also capable of identifying the most reliable configuration of the same model. As illustrated in Figure 3, the confidence vs. OCE curve exhibits a bell shape, unlike ECE metrics, and the optimal OCE is achieved around a threshold of $0.3$. This is particularly notable because it reflects the practical choices in many calibration studies to date [17, 11], affirming the reliability of using OCE.
Using a fixed separation threshold of $0.3$ is often a good approximation for distinguishing between optimal positives and negatives (See further discussion in Appendix G). However, trained or post-calibrated DETR models may have varying optimal confidence thresholds (for example, low-temperature calibrated models are likely to have higher optimal thresholds). Therefore, using a fixed threshold can pose a potential risk.
The biggest advantage of using OCE over other metrics is that OCE measures the model’s calibration quality alongside the post-processing scheme employed. While measuring the calibration error on the optimal positive set would ideally assess quality, this optimal set is not accessible during test-time inference. Consequently, the score measured on the optimal set may not correspond to the quality experienced in real-world applications, such as those employing confidence thresholding schemes.
Table 3: Comparison of the proposed method ($\mathsf{ContrastiveConf}$) with baseline methods on their Pearson correlation coefficient with the image-level reliability in Equation (6), with different models. $\mathsf{Perf}$ is evaluated using AP (Average Precision)[14]. For each model, we apply the optimal matching, the standard post-processing scheme (Top-$100$ and confidence thresholding by $0.3$), and the proposed OCE-based post-processing scheme, as indicated within the parentheses. The strongest correlations are highlighted in bold, while negative correlations appear in red. First, the proposed method achieves the highest correlation with image-level reliability. In addition, it is noteworthy that the average confidence of negative predictions ($\mathsf{Conf}^{-}$) exhibits a negative correlation with image-level reliability, in contrast to the positive correlation observed when using positive predictions ($\mathsf{Conf}^{+}$).
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.503 0.618 0.670 0.635 0.555 0.647 0.649 0.633 0.561 0.642 0.662 0.634
$\mathsf{Conf}^{-}$ (Optimal) -0.608 -0.584 -0.572 -0.586 -0.293 -0.296 -0.330 -0.311 -0.177 -0.235 -0.212 -0.196
$\mathsf{ContrastiveConf}$ (Optimal) 0.700 0.684 0.709 0.695 0.601 0.633 0.662 0.634 0.612 0.657 0.672 0.643
Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.619 -0.603 -0.581 -0.601 -0.409 -0.374 -0.372 -0.391 -0.262 -0.270 -0.239 -0.229
$\mathsf{Conf}^{+}$ (0.3) 0.476 0.464 0.539 0.504 0.229 0.353 0.385 0.399 0.195 0.256 0.258 0.258
Ours $\mathsf{Conf}^{+}$ (OCE) 0.442 0.477 0.547 0.512 0.237 0.355 0.393 0.370 0.228 0.291 0.270 0.274
$\mathsf{Conf}^{-}$ (OCE) -0.633 -0.581 -0.571 -0.585 -0.365 -0.389 -0.379 -0.394 -0.245 -0.325 -0.253 -0.283
$\mathsf{ContrastiveConf}$ (OCE) 0.585 0.570 0.597 0.587 0.334 0.432 0.431 0.438 0.287 0.362 0.302 0.327
6.2 Quantifying Image-Level Reliability
To address the third research question (RQ3), we compare different methods for quantifying per-image reliability using different separation techniques. We measure the Pearson correlation coefficient (PCC) of each method with the $\mathsf{ImReli}$ computed based on AP metrics. Additionally, we evaluate the proposed contrasting approach against methods that utilize only positive or negative queries. Empirical results are presented in Table 3 and further in Appendix F. The key takeaways are as follows:
First, as demonstrated throughout our analysis, distinguishing between positive and negative queries is crucial for uncertainty quantification in DETR. Specifically, when confidence scores are averaged without distinguishing positive queries (i.e., $\mathsf{Conf}^{+}$ w/ top-$100$), the correlation becomes negative. This occurs because the majority of DETR’s predictions are actually negative (i.e., $|\mathcal{T}_{\bm{x}}|\ll|\hat{\mathcal{T}}_{\theta}(\bm{x})|$), and that confidence scores for negative queries are inversely correlated with $\mathsf{ImReli}$.
Second, the proposed $\mathsf{ContrastiveConf}$ consistently achieves the best correlation with image-level reliability. Interestingly, the absolute value of $\mathsf{Conf}^{-}$ occasionally outperforms $\mathsf{ContrastiveConf}$ when a non-optimal separation scheme is applied; however, its performance varies across different settings. This variability underscores the robustness of our contrasting approach, which leverages the strengths of both $\mathsf{Conf}^{+}$ and $\mathsf{Conf}^{-}$.
Table 4: Comparison of different post-processing schemes on OCE (lower is better) and PCC (higher is better) between $\mathsf{ContrastiveConf}$ and image-level reliability in Cal-DETR. The thresholding approach consistently yields the best performance.
Schemes COCO City Foggy
OCE PCC OCE PCC OCE PCC
Optimal 0.425 0.695 0.463 0.634 0.513 0.643
Thr 0.454 0.594 0.448 0.464 0.482 0.325
Top-$k$ 0.502 0.527 0.453 0.411 0.492 0.298
NMS 0.555 0.574 0.582 0.373 0.598 0.239
6.3 Comparative Study on Post-Processing
Lastly, we present a comparative analysis across different post-processing schemes for identifying positive queries in DETR, from the reliability perspective. We assess OCE for each predicted positive to evaluate the impact of various separation schemes on calibration performance.
For the separation methods, we compare the following approaches: (1) applying a threshold on the confidence score, (2) selecting the top-$k$ predictions, and (3) utilizing NMS. For each method, we evaluate OCE by varying the hyperparameter (e.g., confidence threshold, $k$, and IoU threshold, respectively) and compare with the score on the optimal positive. Likewise, we assess the correlation to image-level reliability for each separation scheme. For each scheme, the best hyperparameter is chosen using the validation set. A summary of the results is provided in Table 4, with the comprehensive results available in Appendix G.
The thresholding approach achieves the smallest OCE and the best image-level UQ performance. In addition, the resulting OCE and correlation is the most comparable to the optimal score, especially for the in-distribution test dataset. The top-$k$ approach underperforms because the number of objects in an image varies significantly, making it impractical to determine a single optimal value for $k$. The best $k$ ranges from $10$ to $20$, corresponding to the average number of objects per image in the COCO dataset. On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive queries are included, a substantial number of negative queries are retained, often outnumbering the positive ones by several times. Nonetheless, a performance gap remains compared to optimal matching, especially on out-of-distribution datasets. Addressing this gap is left for future work.
7 Conclusion and Limitations
The main contribution of our work is an in-depth analysis of the importance of post-processing in DETR frameworks. Furthermore, this paper demonstrates the impact of the post-separation method on the reliability of their predictions, and proposes a novel UQ method for quantifying image-level reliability in DETR.
The primary limitation of our work is that our analysis is confined to standard separation schemes, such as thresholding, without proposing a novel separation technique. We conjecture that leveraging information from intermediate layers, rather than solely relying on final predictions, can achieve more robust separation because positive and negative queries follow different evolutionary patterns. One potential approach is to enable the model to explicitly output matching probabilities by introducing an auxiliary loss during training, as proposed by Oksuz et al. [19].
Acknowledgments
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper. This work was supported in part by the MIT Summer Research Program (MSRP), the MIT-Google Program for Computing Innovation, the MIT-Amazon Science Hub, the MIT-IBM Watson AI Lab, and MathWorks.
Appendix A Related Work
Uncertainty Quantification.
Shelmanov et al. [27] and Vazhentsev et al. [29] focus on uncertainty quantification in general Transformer architectures, building upon the MC dropout approach proposed by Gal and Ghahramani [7]. Specifically, they apply a modified MC dropout that prioritizes sample diversity over randomness to improve sample efficiency. However, these methods are not directly applicable to object detection problems because, unlike classification tasks, object detection involves predicting an unordered set. Consequently, it is questionable whether applying MC dropout to each output is appropriate, as the permutation of predictions could be completely shuffled after applying MC dropout.
Several studies focus on out-of-distribution (OOD) identification in object detection models. For example, Li et al. [13] propose a built-in OOD detector to isolate OOD data for human review. This approach can determine OOD scenarios involving both unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty) by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions. Du et al. [5] generate outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data. Similarly, Oksuz et al. [19] employ an auxiliary detection model capable of expressing its confidence. Other works, including Du et al. [4] and Wilson et al. [30], investigate the latent representations generated by object detection models to identify the OOD nature of the input.
To the best of our knowledge, the aforementioned existing uncertainty quantification techniques primarily focus on prediction-level analysis. Moreover, they predominantly address CNN-based models and explore the methodological way to better quantify the uncertainty in object detection model. In contrast, our paper emphasizes the significance of identifying reliable sets within the entire set of predictions for uncertainty quantification, particularly in DETRs. Another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
Calibration.
Previous studies primarily addressed OOD detection to assess the prediction-level trustworthiness of DETR. They aim to introduce a framework that does not rely on the model’s confidence scores, which can be either overconfident or underconfident, especially for OOD data points. Simultaneously, additional research endeavors to determine the model’s overall reliability in its confidence scores. More specifically, these studies assess whether the model’s confidence scores align with their actual accuracy. In other words, low-confidence samples should exhibit low accuracy, and high-confidence samples should demonstrate high accuracy. To evaluate this alignment, they measure the Expected Calibration Error (ECE) by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin. However, in object detection tasks, measuring accuracy is not straightforward because predictions comprise both class probabilities and bounding boxes. Additionally, due to the set prediction nature of object detection models, it is unclear which ground truth object corresponds to each prediction. To address these challenges, D-ECE [10] defines precision as the accuracy metric and matches each prediction to a ground truth object based on an IoU threshold (commonly set at 0.5 or 0.75). On the other hand, LaECE [18] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
However, as demonstrated throughout our paper, DETR exhibits poor calibration on optimal negative predictions, while in practice, users prefer to utilize only predicted positive samples (e.g., high-confidence predictions). Therefore, if we rank the model’s calibration quality based on the entire prediction set, these rankings are highly biased and thus unreliable. To avoid this issue, previous studies have often measured D-ECE only on predictions with confidence scores exceeding $0.3$. As detailed in Appendix G, a threshold of $0.3$ approximates the optimal positive predictions in many DETR variants; thus, using such a fixed threshold may be practically acceptable. However, this threshold is not guaranteed to be $0.3$ for other models like UP-DETR, which introduces potential risks. In response, our paper addresses the importance of identifying a reliable subset of predictions and proposes the Object-level Calibration Error (OCE) metric, which measures the model’s calibration quality alongside the employed post-processing scheme and thus can adaptively identify the reliable subset.
Broad Implications.
The primary objective of this study is to determine how DETR can be used in a trustworthy manner on new downstream images/tasks and to establish how their predictions should be properly used, which has several important implications. For example, when building an auto-labeling system using data-driven detection models like DETR, it is crucial to consider that the model may not always provide accurate predictions. As a result, when the model’s reliability for a specific image is in question, the image might need to be reviewed by a human labeler. Therefore, evaluating reliability at the image level is often necessary to assess the model’s overall understanding of the given test image. Furthermore, by providing a comprehensive analysis for identifying a reliable subset of predictions, our study offers users a systematic and trustworthy method for post-processing the outputs of DETR. We hope our efforts expand the scope of DETR applications by enabling more precise and reliable deployment, sparking further research into the area of positive prediction identification.
Figure 4: A diagram of the DETR architecture. An input image is first processed through a CNN backbone to generate a 2D feature representation. This representation is then passed to the Transformer encoder, which extracts feature vectors. These feature vectors are sent to the decoder, which receives $M$ learned object queries together. The decoder outputs $M$ prediction sets, each containing a bounding box and corresponding class probabilities. Please refer to the original papers for details.
Appendix B Bipartite Matching: Positive & Negative Queries
Since the number of queries in DETR, $M$, is much higher than the number of ground truth objects, DETR matches each ground truth object with the corresponding best model prediction during its training. To compute this optimal (i.e., ground truth) matching for the predictions in a given image, a bipartite matching algorithm is applied. More specifically, a matching cost between each pair of a given prediction and a ground truth object is defined as follows:
$\mathcal{L}_{matching}=\mathcal{L}_{class}+\mathcal{L}_{box}$ (12)
where $\mathcal{L}_{class}$ is the negative prediction confidence of the ground truth class and $\mathcal{L}_{box}$ is the linear combination of the $\ell_{1}$ loss between the corners of the bounding boxes and $\mathcal{L}_{iou}$. $\mathcal{L}_{iou}$ is the Generalized Intersection over Union (GIoU) [24] loss between bounding boxes. After computing this matching cost for every combination of prediction set and ground truth objects, DETR then efficiently calculates the permutation that minimizes the total matching cost using the Hungarian matching algorithm.
In this paper, we then refer to those matched queries as positive predictions and the remaining unmatched queries as negative predictions. More details on the bipartite matching process and Hungarian algorithm can be found in [1].
Appendix C Further Visualizations of How DETR Generates Predictions
The predictions across all six decoder layers for each of the images presented in the main paper are provided in Figure 5. Another noticeable example is provided in Figure 6, where predictions for two different objects, a dog and a laptop, are shown for the same low-reliability image. In Figure 6(a), for the laptop object, all of the queries start off with low confidence and remain that way over the course of the layers. However, the class predictions gradually shift from dining table and person to laptop, and result in an overall low confidence score despite a correct final prediction (laptop).
However, Figure 6(b) indicates that even in a low-reliability image there can still be reliable predictions. For the dog object, the model is very confident in its predictions from the start, with many overlapping prediction sets in the same area and with matching classes. While there are two dominant predictions in the first layer, DETR focuses on a single prediction (query 0) while the confidence for query 1 is gradually reduced; this aligns with the trend observed in the reliable examples shown earlier.
Appendix D Performance Metrics
D.1 Expected Calibration Error (ECE) Metrics
Detection Expected Calibration Error (D-ECE) [10] is a calibration metric that specifically focuses on calibrating the confidence of an object detection model to align with precision. The confidence space is divided into $J$ bins, and then D-ECE is computed as follows:
[…]
where $\hat{\mathcal{D}}$ and $\hat{\mathcal{D}}_{j}$ are the set of all detections and the detections in the $j$-th bin, and $\bar{p}_{j}$ and precision$(j)$ are the average confidence and the precision of the detections in the $j$-th bin. D-ECE identifies true positive (TP) samples by confirming the accuracy of class labels and ensuring that the IoU surpasses the threshold $\tau$. Following the standard protocol [11], this work measures D-ECE using thresholds $\tau=0.5$ and $\tau=0.75$, and then averages the scores.
To account for both localization error and classification precision, LaECE [19, 11] calibrates model confidence to the product of precision and the IoU between the bounding box predictions and ground-truth boxes. LaECE is computed in a class-wise manner to combat class imbalance, as follows:
$\text{LaECE}\triangleq\frac{1}{K}\sum_{c=1}^{K}\sum_{j=1}^{J}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}|\bar{p}_{j}^{c}-\text{precision}^{c}(j)\times\bar{\text{IoU}}^{c}(j)|$ (14)
where $c$ refers to each class and $\bar{\text{IoU}}^{c}(j)$ refers to the average IoU of $\mathcal{\hat{D}}^{c}_{j}$. Similar to D-ECE, LaECE requires an IoU threshold, $\tau$. In contrast, LaECE0 proposes to set this threshold $\tau$ to 00 and instead set $\bar{\text{IoU}}^{c}(\cdot)=0$ for the false positive (FP) samples. This leads to a reduction of the calibration criterion of LaECE as follows:
$\text{LaECE}_{0}=\frac{1}{K}\sum^{K}_{c=1}\sum^{J}_{j=1}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}|\bar{p}^{c}_{j}-\bar{\text{IoU}}^{c}(j)|$ (15)
D.2 Performance Metrics
Average precision (AP) [6] is the most common accuracy metric for object detectors. Before calculating AP, the top-$k$ (e.g., typically $100$) class predictions for a given image are kept and sorted, while the others are discarded. Then, in order of this sorting, the predictions are compared against against the ground-truth bounding boxes, and the predicted label is compared with the ground truth label if the IoU is large enough. The precision/recall curve is calculated over the sorted predictions, and AP averages the precision across a set of 101 evenly spaced recall thresholds:
$\text{AP}=\frac{1}{101}\sum_{r\in[0:0.01:1]}p(r)$ (16)
where $p(r)$ is calculated by the maximum precision at recall level $r$.
[…]
$\text{LRP}_{\tau}=\frac{1}{N_{\text{FP}}+N_{\text{FN}}+N_{\text{TP}}}\left(N_{\text{FP}}+N_{\text{FN}}+\sum_{i=1}^{N_{\text{TP}}}\frac{1-\text{IoU}_{i}}{1-\tau}\right)$ (17)
where $\tau$ is IoU threshold, $i$ denotes the index of each true positive prediction, and IoUi represents the IoU between that prediction and the best-matching ground-truth object. LRP is then computed by averaging LRP0.5 and LRP0.75.
D.3 Impact of Separation Schemes on Metrics
The trends presented in Figure 3 also hold for other datasets and models. Figure 7 shows the performance of the model according to different metrics with various confidence thresholds (i.e. post-processing schemes). As illustrated in Section 4.1, the optimal AP is achieved when the threshold is set to $0.0$. In contrast, the optimal ECEs are often achieved when the threshold is set close to $1.0$, meaning ECEs favor retaining fewer predictions with high confidence. However, as discussed earlier, using the entire or small subset carries a high risk of including uncalibrated negatives or having missing detections, leading to unreliable decisions in practical applications and diminished interpretability.
Appendix E Experiment Setup
For our experiments, we used the Cityscapes and Foggy Cityscapes datasets, which each have 500 images of first-person driving footage in realistic environments. Foggy Cityscapes has the same base images as Cityscapes, but with fog simulated and added to create a further out of distribution set. Since the DETR models were trained on COCO, the Cityscapes and Foggy Cityscapes annotations were converted to correspond to the labels of COCO. More specifically, the person, bicycle, car, motorcycle, bus, train, and truck classes were transferred directly. In addition, the rider class of Cityscapes and Foggy Cityscapes was mapped to the person class of COCO. The other classes present in Cityscapes and Foggy Cityscapes are largely focused on image segmentation, and thus were omitted (e.g. building, sky, sidewalk). The pre-trained model weights were obtained from their respective official implementations.
To compute $\mathsf{ImReli}$ for each image, we use the same COCO evaluator [14] to obtain image-wise AP scores. However, rather than passing in the entire image set, the predictions for each image are passed in individually.
Appendix F Further Experimental Results
F.1 Effectiveness of OCE
Table 5 presents the correlations between the rankings provided by three different calibration metrics on the optimal positive set and various methods. The first three rows demonstrate that the calibration metrics—D-ECE, LA-ECE, and our proposed OCE—–are highly correlated. This indicates that all three metrics effectively capture the notion of calibration quality. Additionally, these metrics show a decent correlation with AP metrics. It is important to note that AP primarily accounts for accuracy rather than calibration quality, which may explain why the correlation, particularly for AP@50 on the out-of-distribution dataset, is not perfectly aligned.
Furthermore, LRP, D-ECE, and LA-ECE0 exhibit negative correlations with calibration quality when used alone. This supports our assertion that these metrics are inadequate for measuring models’ calibration quality in conjunction with the employed post-processing scheme, unlike our OCE.
Finally, D-ECE exhibits a high correlation when paired with either a fixed threshold or a proper adaptive threshold (i.e., using LRP and our OCE). This is because DETR’s optimal positive threshold is empirically found to be around $0.3$ (See Table 7.). Consequently, OCE aligns well with the experimentally approximated threshold and the LRP metric in identifying effective post-processing schemes. This alignment further supports the robustness and effectiveness of our method.
F.2 Quantifying Image-Level Reliability
This section provides an extended version of Table 3 with more methods such as the proposed method but with a fixed threshold. Furthermore, we show more results with different $\mathsf{Perf}$ metrics by evaluating it with AP50 and AP75 (average precision with Iou threshold 50% and 75%). The results are shown in Table 6.
As illustrated, the observation is consistent across different settings. Another notable observation is the comparison between the fixed threshold versus our adaptive threshold selection using OCE. Although the proposed fixed-threshold approach ($0.3$) occasionally outperforms the adaptive method, the performance gap remains marginal. Moreover, the fixed threshold consistently exhibits a lower correlation with UP-DETR, for which the optimal threshold has been empirically determined to be approximately $0.5$ ($0.45$ by OCE and $0.55$ by LRP). This indicates that although the OCE approach may not achieve threshold selection that best approximates the optimal positives, it ensures the robustness of uncertainty quantification performance. In contrast, fixed thresholding can degrade performance when the chosen value deviates from the optimal threshold.
Finally, we perform an ablation study on the scaling factor ($\lambda$), with the results presented in Figure 8. The study shows that the optimal scaling factor lies within the range of $5.0$ to $10.0$. Notably, this range remains effective even for out-of-distribution datasets, such as Cityscapes and Foggy Cityscapes. This consistency suggests that the scaling factor selected using the validation set (e.g., COCO) generalizes well across datasets. Furthermore, it demonstrates that $\mathsf{ContrastiveConf}$ with $\lambda>0.0$ consistently outperforms $\mathsf{Conf}^{+}$ (i.e., $\lambda=0.0$).
Appendix G Comparative Study on Post-Processing
In this section, we provide a comprehensive analysis regarding the impact of post-processing on the model’s overall calibration quality. First, we vary the hyperparameter of the standard post-processing processes. In this first analysis, to exclude the dependency of the final performance on OCE, we don’t use OCE as a selection criterion. Rather, we choose the best hyperparameter on the validation set for each setting and compare their best-possible performances. The results are shown in Table 7 .
As illustrated, we can observe that the confidence thresholding approach outperforms the top-$k$ and NMS approaches. Top-$k$ achieves the minimum OCE with top-$20$. Given that the average and 95th percentile number of objects per image in the COCO dataset are 7 and 22, respectively, these results appear reasonable. Therefore, we confirm that using an excessively large number (e.g., 100) for top-$k$ is inadequate for achieving well-calibrated predictions.
Similarly, we compare the image-level uncertainty quantification performance, and the results are shown in Table 8. We reconfirm the effectiveness of confidence thresholding; for Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds exceed $0.5$, highlighting the potential limitation of using a fixed threshold in image-level UQ applications. Moreover, the correlation is negative especially when NMS is applied along with the $\mathsf{Conf}^{+}$ framework. This is because, while most of the optimal positive queries are likely to be included with NMS, a substantial number of optimal negative queries remain within the final subset, often outnumbering the positive ones by several times. As a result, applying NMS leads to an inaccurate reliability assessment. However, many schemes, including NMS, achieve significant improvement when used with the proposed contrastive framework. This is because even if the post-processing scheme is inaccurate, the framework can robustly perform by leveraging the negative queries.
Lastly, we plot the OCE and image-level UQ performance with different hyperparameter selections to show the sensitivity of each scheme to the choice of hyperparameter. As depicted in Figure 10, the empirical results highlight that carefully identifying a reliable subset is crucial for achieving both high object-level calibration quality and effective image-level uncertainty quantification performance. We also highlight this claim with an exemplary visualization in Figure 9.
Table 5: Comparison of the effectiveness of the proposed OCE and baseline metrics, focusing on their ranking correlation with the rankings given by the models’ calibration quality on optimal positive predictions (denoted by the superscript *). For each metric, various post-processing schemes (i.e., ways to identify positive predictions) are applied and indicated within parentheses. Correlations exceeding $0.9$ are highlighted in blue, while negative correlations are highlighted in red. Our OCE-based rankings consistently provide strong correlations with optimal ranking and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD) Average
[…]
References D-ECE 1.000 0.995 0.969 1.000 0.996 0.993 1.000 0.873 0.996 0.980 $\pm$ 0.039
[…]
$\mathsf{OCE_{ENS{}}}^{*}$ 0.969 0.948 1.000 0.993 0.990 1.000 0.996 0.836 1.000 0.970 $\pm$ 0.050
Baselines AP (Top-100) [6, 14] 0.552 0.583 0.343 0.779 0.784 0.700 0.597 0.843 0.523 0.634 $\pm$ 0.149
AP@50 (Top-100) [14] 0.380 0.409 0.164 0.687 0.710 0.606 -0.360 0.079 -0.440 0.249 $\pm$ 0.402
AP@75 (Top-100) [14] 0.652 0.680 0.458 0.830 0.829 0.757 0.765 0.930 0.705 0.734 $\pm$ 0.127
LRP (LRP) [18] 0.103 0.051 0.338 -0.268 -0.272 -0.150 -0.257 -0.610 -0.170 -0.137 $\pm$ 0.257
D-ECE (D-ECE) -0.757 -0.700 -0.772 0.365 0.284 0.337 0.309 0.436 0.336 -0.018 $\pm$ 0.514
D-ECE (Top-100) -0.873 -0.833 -0.966 0.608 0.576 0.511 -0.254 0.250 -0.316 -0.144 $\pm$ 0.614
D-ECE (0.3) [10] 0.905 0.939 0.827 0.982 0.967 0.990 0.964 0.799 0.981 0.928 $\pm$ 0.067
D-ECE (LRP) 0.802 0.855 0.659 0.987 0.984 0.999 0.983 0.793 0.996 0.895 $\pm$ 0.116
D-ECE ($\mathsf{OCE_{ENS}}$) 0.856 0.901 0.727 0.987 0.983 0.999 0.983 0.793 0.996 0.914 $\pm$ 0.095
LA-ECE0 (LA-ECE0) -0.960 -0.930 -0.968 -0.913 -0.874 -0.907 -0.826 -0.839 -0.835 -0.895 $\pm$ 0.051
LA-ECE0 (Top-100) -0.944 -0.913 -0.994 -0.431 -0.482 -0.517 -0.786 -0.386 -0.822 -0.697 $\pm$ 0.228
LA-ECE0 (0.3) -0.789 -0.729 -0.842 -0.780 -0.818 -0.830 -0.451 -0.041 -0.468 -0.639 $\pm$ 0.254
LA-ECE0 (LRP) [11] -0.908 -0.864 -0.948 -0.833 -0.877 -0.846 -0.901 -0.597 -0.913 -0.854 $\pm$ 0.097
LA-ECE0 ($\mathsf{OCE_{ENS}}$) -0.901 -0.855 -0.943 -0.723 -0.780 -0.739 -0.884 -0.568 -0.897 -0.810 $\pm$ 0.112
Ours $\mathsf{OCE_{ENS}}$ ($\mathsf{OCE_{ENS}}$) 0.962 0.940 1.000 0.987 0.983 0.999 0.964 0.737 0.984 0.951 $\pm$ 0.078
$\mathsf{OCE_{MAX}}$ ($\mathsf{OCE_{MAX}}$) 0.954 0.929 0.998 0.954 0.958 0.983 0.955 0.697 0.976 0.934 $\pm$ 0.086
Table 6: Comparison of the proposed method ($\mathsf{ContrastiveConf}$) with baseline methods on their Pearson correlation coefficient with the image-level reliability in Equation (6), with different models. $\mathsf{Perf}$ is evaluated using AP (top), AP50 (middle), and AP75 (bottom) metrics. For each model, we apply the optimal matching, the standard post-processing scheme (Top-$100$ and confidence thresholding by $0.3$), and the proposed OCE-based post-processing scheme, as indicated within the parentheses. The strongest correlations are highlighted in bold, while negative correlations appear in red. First, the proposed method achieves the highest correlation with image-level reliability. In addition, it is noteworthy that the average confidence of negative predictions ($\mathsf{Conf}^{-}$) exhibits a negative correlation with image-level reliability, in contrast to the positive correlation observed when using positive predictions ($\mathsf{Conf}^{+}$). Although the proposed fixed-threshold approach (0.3) occasionally outperforms the adaptive method, the performance gap remains marginal. Moreover, the fixed threshold consistently exhibits lower correlation with UP-DETR, for which the optimal threshold has been empirically determined to be approximately 0.5.
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.503 0.618 0.670 0.635 0.555 0.647 0.649 0.633 0.561 0.642 0.662 0.634
$\mathsf{Conf}^{-}$ (Optimal) -0.608 -0.584 -0.572 -0.586 -0.293 -0.296 -0.330 -0.311 -0.177 -0.235 -0.212 -0.196
$\mathsf{ContrastiveConf}$ (Optimal) 0.700 0.684 0.709 0.695 0.601 0.633 0.662 0.634 0.612 0.657 0.672 0.643
Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.619 -0.603 -0.581 -0.601 -0.409 -0.374 -0.372 -0.391 -0.262 -0.270 -0.239 -0.229
$\mathsf{Conf}^{+}$ (0.3) 0.476 0.464 0.539 0.504 0.229 0.353 0.385 0.399 0.195 0.256 0.258 0.258
Ours $\mathsf{Conf}^{+}$ (OCE) 0.442 0.477 0.547 0.512 0.237 0.355 0.393 0.370 0.228 0.291 0.270 0.274
$\mathsf{Conf}^{-}$ (OCE) -0.633 -0.581 -0.571 -0.585 -0.365 -0.389 -0.379 -0.394 -0.245 -0.325 -0.253 -0.283
$\mathsf{ContrastiveConf}$ (OCE) 0.585 0.570 0.597 0.587 0.334 0.432 0.431 0.438 0.287 0.362 0.302 0.327
$\mathsf{ContrastiveConf}$ (0.3) 0.568 0.574 0.599 0.594 0.297 0.437 0.430 0.464 0.248 0.346 0.296 0.325
$\mathsf{ContrastiveConf}$ (Top-100) -0.619 -0.253 -0.490 -0.321 -0.409 -0.342 -0.358 -0.360 -0.262 -0.175 -0.215 -0.139
(a) AP
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.488 0.591 0.637 0.611 0.584 0.638 0.616 0.642 0.598 0.640 0.624 0.639
$\mathsf{Conf}^{-}$ (Optimal) -0.583 -0.580 -0.555 -0.575 -0.248 -0.305 -0.316 -0.324 -0.137 -0.185 -0.175 -0.156
$\mathsf{ContrastiveConf}$ (Optimal) 0.675 0.664 0.677 0.673 0.594 0.630 0.629 0.646 0.629 0.634 0.629 0.632
Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.586 -0.581 -0.546 -0.571 -0.375 -0.369 -0.351 -0.394 -0.216 -0.217 -0.196 -0.204
$\mathsf{Conf}^{+}$ (0.3) 0.445 0.435 0.494 0.465 0.229 0.321 0.342 0.379 0.203 0.223 0.179 0.194
Ours $\mathsf{Conf}^{+}$ (OCE) 0.426 0.443 0.501 0.471 0.242 0.331 0.349 0.364 0.256 0.239 0.198 0.211
$\mathsf{Conf}^{-}$ (OCE) -0.600 -0.576 -0.554 -0.573 -0.317 -0.394 -0.357 -0.406 -0.194 -0.277 -0.216 -0.247
$\mathsf{ContrastiveConf}$ (OCE) 0.560 0.545 0.554 0.552 0.314 0.416 0.388 0.438 0.285 0.302 0.229 0.263
$\mathsf{ContrastiveConf}$ (0.3) 0.532 0.553 0.557 0.561 0.277 0.418 0.388 0.456 0.237 0.298 0.216 0.259
$\mathsf{ContrastiveConf}$ (Top-100) -0.586 -0.189 -0.435 -0.250 -0.375 -0.315 -0.332 -0.340 -0.216 -0.122 -0.171 -0.121
(b) AP@50
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.490 0.593 0.629 0.598 0.486 0.586 0.608 0.563 0.488 0.607 0.635 0.574
$\mathsf{Conf}^{-}$ (Optimal) -0.572 -0.541 -0.521 -0.538 -0.280 -0.244 -0.284 -0.274 -0.154 -0.234 -0.182 -0.189
$\mathsf{ContrastiveConf}$ (Optimal) 0.672 0.649 0.662 0.649 0.541 0.562 0.614 0.563 0.534 0.627 0.641 0.587
Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.580 -0.555 -0.523 -0.549 -0.373 -0.314 -0.325 -0.348 -0.226 -0.266 -0.206 -0.216
$\mathsf{Conf}^{+}$ (0.3) 0.459 0.434 0.497 0.467 0.214 0.320 0.348 0.351 0.188 0.231 0.290 0.241
Ours $\mathsf{Conf}^{+}$ (OCE) 0.423 0.447 0.503 0.478 0.234 0.330 0.361 0.322 0.206 0.272 0.293 0.262
$\mathsf{Conf}^{-}$ (OCE) -0.597 -0.539 -0.520 -0.537 -0.339 -0.325 -0.327 -0.342 -0.220 -0.321 -0.217 -0.269
$\mathsf{ContrastiveConf}$ (OCE) 0.557 0.532 0.548 0.545 0.319 0.386 0.390 0.381 0.258 0.346 0.314 0.312
$\mathsf{ContrastiveConf}$ (0.3) 0.543 0.535 0.551 0.548 0.279 0.384 0.385 0.406 0.233 0.324 0.316 0.306
$\mathsf{ContrastiveConf}$ (Top-100) -0.580 -0.225 -0.435 -0.291 -0.373 -0.295 -0.320 -0.334 -0.226 -0.170 -0.188 -0.137
(c) AP@75
Table 7: Comparison of post-processing schemes based on calibration quality (i.e., OCE). We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding). For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses, and is applied on the test set. the strongest correlations are highlighted in bold. Confidence thresholding achieves the lowest OCE, demonstrating its better efficacy compared to the other schemes. The optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies.
Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
Thresholding 0.276 (0.35) 0.450 (0.20) 0.416 (0.25) 0.436 (0.25) 0.313 (0.45) 0.430 (0.30) 0.413 (0.30) 0.402 (0.35) 0.387 (0.40) 0.488 (0.30) 0.469 (0.30) 0.458 (0.30)
Top-$k$ 0.358 (20.00) 0.485 (20.00) 0.457 (20.00) 0.479 (20.00) 0.352 (20.00) 0.451 (20.00) 0.436 (20.00) 0.422 (20.00) 0.416 (20.00) 0.500 (20.00) 0.490 (20.00) 0.473 (20.00)
NMS 0.358 (0.90) 0.535 (0.90) 0.624 (0.90) 0.521 (0.90) 0.426 (0.90) 0.625 (0.90) 0.784 (0.90) 0.583 (0.90) 0.495 (0.90) 0.650 (0.90) 0.794 (0.90) 0.609 (0.90)
Table 8: Comparison of post-processing schemes and baseline methods based on their Pearson correlation coefficients with image-level reliability when $\mathsf{Conf}^{+}$ and $\mathsf{ContrastiveConf}$ are applied, respectively. We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding). For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses, and is applied on the test set. the strongest correlations are highlighted in bold. Confidence thresholding achieves the lowest OCE, demonstrating its better efficacy compared to the other schemes. For Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds exceed $0.5$, highlighting the potential limitation of using a fixed threshold in image-level uncertainty quantification (UQ) applications.
UQ Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
$\mathsf{Conf}^{+}$ Thresholding 0.479 (0.25) 0.478 (0.20) 0.547 (0.25) 0.512 (0.25) 0.333 (0.80) 0.355 (0.25) 0.411 (0.20) 0.403 (0.35) 0.229 (0.50) 0.291 (0.25) 0.281 (0.15) 0.282 (0.35)
Top-$k$ 0.067 (1.00) 0.132 (1.00) 0.252 (1.00) 0.155 (1.00) -0.037 (2.00) 0.019 (1.00) 0.078 (1.00) 0.021 (1.00) 0.080 (1.00) -0.006 (1.00) 0.026 (1.00) 0.090 (1.00)
NMS -0.588 (0.10) -0.560 (0.10) -0.548 (0.10) -0.562 (0.10) -0.402 (0.10) -0.346 (0.10) -0.332 (0.10) -0.365 (0.10) -0.240 (0.10) -0.265 (0.10) -0.218 (0.10) -0.224 (0.10)
$\mathsf{ContrastiveConf}$ Thresholding 0.597 (0.55) 0.574 (0.30) 0.599 (0.30) 0.594 (0.30) 0.408 (0.80) 0.437 (0.30) 0.443 (0.20) 0.472 (0.35) 0.312 (0.75) 0.362 (0.25) 0.315 (0.40) 0.352 (0.35)
Top-$k$ 0.629 (1.00) 0.506 (1.00) 0.474 (1.00) 0.527 (1.00) 0.409 (1.00) 0.377 (1.00) 0.324 (1.00) 0.411 (1.00) 0.284 (1.00) 0.254 (1.00) 0.156 (1.00) 0.298 (1.00)
NMS -0.408 (0.10) 0.585 (0.10) 0.535 (0.10) 0.574 (0.10) -0.101 (0.10) 0.368 (0.10) 0.347 (0.10) 0.373 (0.10) -0.001 (0.10) 0.262 (0.10) 0.228 (0.10) 0.239 (0.10)
(a) Predictions for a High-Reliability Prediction: Horse
(b) Predictions for a High-Reliability Prediction: Chair
(c) Predictions for a Low-Reliability Prediction: Chair
Figure 5: The evolution of predictions made by Cal-DETR across decoder layers for a high-reliability and low-reliability image. The positive query (indexed by 0 and bordered with green) and the five negative queries with the largest IoU (indexed by 1-5 and bordered with red) are presented.
(a) Predictions for a Low-Reliability Object: Laptop
(b) Predictions for a High-Reliability Object: Dog
Figure 6: The evolution of predictions made by Cal-DETR across decoder layers for two different objects (one reliable and one unreliable) in a low-reliability image. The positive query (indexed by 0 and bordered with green) and the five negative queries with the largest IoU (indexed by 1-5 and bordered with red) are presented.
(a) UP-DETR (COCO)
(b) D-DETR (COCO)
(c) Cal-DETR (COCO)
(d) DINO (COCO)
(e) UP-DETR (Cityscapes)
(f) D-DETR (Cityscapes)
(g) Cal-DETR (Cityscapes)
(h) DINO (Cityscapes)
(i) UP-DETR (Foggy)
(j) D-DETR (Foggy)
(k) Cal-DETR (Foggy)
(l) DINO (Foggy)
Figure 7: Impact of confidence threshold selection on various performance metrics in UP-DETR, Deformable-DETR, Cal-DETR, and DINO on COCO, Cityscapes, and Foggy Cityscapes.
(a) COCO
(b) Cityscapes
(c) Foggy Cityscapes
Figure 8: Impact of the scaling factor ($\lambda$) on image-level UQ performance of $\mathsf{ContrastiveConf}$. Pearson correlation coefficient (PCC) using various scaling factors is reported. The optimal scaling factor lies within the range of $5.0$ to $10.0$, while this range generalizes well across out-of-distribution datasets. Furthermore, it shows the efficacy of $\mathsf{ContrastiveConf}$ over $\mathsf{Conf}^{+}$ (i.e., $\mathsf{ContrastiveConf}$ with $\lambda=0.0$).
(a) Thresholding (0.1)
(b) Thresholding (0.3)
(c) Thresholding (0.9)
(d) Top-$1$
(e) Top-$20$
(f) Top-$100$
(g) NMS ($\tau=0.1$)
(h) NMS ($\tau=0.5$)
(i) NMS ($\tau=0.9$)
Figure 9: Exemplary visualization demonstrating the impact of parameter selection on the final subset of predictions in Cal-DETR for different post-processing schemes. Optimal positive and negative predictions are highlighted with green and red boxes, respectively. As shown, the top-$k$ and NMS approaches often include too many negative predictions, degrading the calibration quality. Confidence thresholding with too low of a threshold faces a similar issue, while too high of a threshold risks omitting positive predictions. Therefore, accurately identifying a reliable set of predictions significantly affects the reliability of DETR for downstream applications.
(a) Thresholding (UP-DETR)
(b) Top-$k$ (UP-DETR)
(c) NMS (UP-DETR)
(d) Thresholding (D-DETR)
(e) Top-$k$ (D-DETR)
(f) NMS (D-DETR)
(g) Thresholding (Cal-DETR)
(h) Top-$k$ (Cal-DETR)
(i) NMS (Cal-DETR)
(j) Thresholding (DINO)
(k) Top-$k$ (DINO)
(l) NMS (DINO)
Figure 10: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{ContrastiveConf}$ and image-level reliability in Cal-DETR on COCO for different post-processing schemes.

### v2

@@
-Identifying Reliable Predictions in Detection Transformers
+Quantifying the Reliability of Predictions in Detection Transformers: Object-Level Calibration and Image-Level Uncertainty
 Abstract
@@
 More specifically, while multiple predictions are often made for a single object, our findings show that most often one such prediction is well-calibrated, and the others are poorly calibrated.
-Based on these insights, we demonstrate identifying a reliable subset of DETR’s predictions is crucial for accurately assessing the reliability of the model at both object and image levels.
-Building on this viewpoint, we first tackle the shortcomings of widely used performance and calibration metrics, such as average precision and various forms of expected calibration error.
+Based on these insights, we demonstrate that identifying a reliable subset of DETR’s predictions is crucial for accurately assessing the reliability of the model at both object and image levels.
+Building on this viewpoint, we first address the shortcomings of widely used performance and calibration metrics, such as average precision and various forms of expected calibration error.
 Specifically, they are inadequate for determining which subset of DETR’s predictions should be trusted and utilized.
-In response, we present Object-level Calibration Error (OCE), which is capable of assessing the calibration quality both across different models and among various configurations within a specific model.
+In response, we present Object-level Calibration Error (OCE), which assesses the calibration quality more effectively and is suitable for both ranking different models and identifying the most reliable predictions within a specific model.
 As a final contribution, we introduce a post hoc uncertainty quantification (UQ) framework that predicts the accuracy of the model on a per-image basis.
-By contrasting the average confidence scores of positive (i.e., likely to be matched) and negative predictions determined by OCE, the framework assesses the reliability of the DETR model for each test image.
-The code is available at https://github.com/azizanlab/reliable-detr.
-(a) Low confidence with random boxes
-(b) Calibrated confidence with similar boxes
-(c) Uncalibrated confidence with similar boxes
-Figure 1: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with one being well-calibrated.
-This figure illustrates how DETR can handle the remaining predictions.
-In principle, DETR may (a) generate low confidence with random bounding boxes, (b) assign equally calibrated confidence with similar boxes, or (c) poorly calibrated confidence with similar boxes.
+By contrasting the average confidence scores of positive (i.e., likely to be matched) and negative predictions determined by OCE, our framework assesses the reliability of the DETR model for each test image.
+(a) Random boxes. Low Confidence.
+(b) Accurate boxes. Calibrated.
+(c) Accurate boxes. Uncalibrated.
+Figure 1: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with at least one (i.e., blue boxes) being well-calibrated.
+This figure illustrates how DETR can handle the remaining predictions (i.e., red and gray boxes).
+In principle, DETR may (a) generate low confidence with random bounding boxes, (b) assign equally calibrated confidence with accurate prediction for the bounding boxes, or (c) poorly calibrated confidence with accurate boxes.
 Our analysis indicates that DETR mostly follows the third scheme; resulting in varying levels of reliability across predictions.
-1 Introduction
+I Introduction
 Object detection is an essential task in computer vision, with applications that span various domains including autonomous driving, warehousing, and medical image analysis.
-Existing object detection methods predominantly utilize Convolutional Neural Networks (CNNs) [8, 23, 22, 9] to identify and locate objects within images.
-However, the recent introduction of DEtection TRansformer (DETR) [1] has revolutionized the field by offering an end-to-end prediction pipeline where the model predicts a set of bounding boxes and class probabilities.
-The core innovation of DETR lies in the use of a Transformer encoder-decoder architecture, enabling the model to generate predictions in an end-to-end manner and enhancing scalability.
-This paradigm shift has led to the exploration of various DETR variants—such as Deformable-DETR [32] and DINO [31]—positioning them as potential foundation models for object detection tasks.
-Despite these advancements, the inner workings of how these predictions are generated and interact within the Transformer decoder layers remain under-explored.
-Likewise, there is an ongoing debate within the community about whether DETR truly provides an end-to-end prediction pipeline.
-The central concern is whether all the hundreds of predictions generated by DETR can be trusted and used.
-By means of an answer, practitioners often employ heuristic approaches to achieve high precision, such as by setting a user-defined threshold to retain only a small subset of high-confidence outputs, as seen in the official demo.
-Previous studies exploring model reliability, such as [10, 15, 16, 21, 17], have also applied a user-defined confidence threshold (e.g., 0.3) to retain a subset of predictions for evaluating calibration quality, rather than using the entire set.
-On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR and 300 out of 900 for DINO) based on confidence scores during post-processing to achieve a high average precision (AP) score.
-However, the significance of selecting a subset of the predictions as well as the choice of a suitable configuration for doing so remain under-explored.
-We provide a more in-depth discussion about related works in Appendix A.
-This paper examines these issues by addressing the following three research questions: (RQ1) Do all predictions generated for a given image exhibit comparable levels of reliability?
+Existing object detection methods predominantly utilize Convolutional Neural Networks (CNNs) [1, 2, 3, 4, 5, 6] to identify and locate objects within images.
+More recently, DEtection TRansformer (DETR) [7] has revolutionized the field by utilizing a Transformer encoder-decoder architecture to offer a scalable end-to-end prediction pipeline, where the model predicts a set of bounding boxes and class probabilities.
+This paradigm shift has led to the exploration of various DETR variants, positioning them as potential foundation models for object detection tasks.
+While notable progress has been made, the reliability of these predictions remains insufficiently investigated.
+The primary objective of this study is to determine how DETR can be used in a trustworthy manner on new downstream images/tasks and to establish how their predictions should be properly used.
+For example, when building an auto-labeling system using DETRs, it is crucial to consider that the model may not always provide accurate predictions.
+As a result, when the model’s reliability for a specific image is in question, the image might need to be reviewed by a human labeler.
+Therefore, evaluating reliability at the image level is often necessary to assess the model’s overall understanding of the given test image.
+However, owing to DETR’s unique set-prediction mechanism, quantifying its image-level reliability is not straightforward and remains largely underexplored.
+Moreover, there is an ongoing debate within the community about whether DETR can be considered an entirely end-to-end solution, free from any post-processing requirements in practice.
+In particular, DETR outputs a fixed number of predictions, typically in the hundreds; consequently, the central concern is which predictions can be trusted and used.
+Practitioners often employ heuristic approaches to determine this subset, such as by setting a user-defined threshold to retain only a small subset of high-confidence (e.g., $>0.7$) outputs, as seen in the official demo.
+Similarly, previous studies exploring model reliability, such as [8, 9, 10, 11, 12, 13], have also applied a user-defined confidence threshold (e.g., 0.3) to retain a subset of predictions for evaluating calibration quality, rather than using the entire set.
+On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR [14] and 300 out of 900 for DINO [15]) based on confidence score.
+Nonetheless, how different subset selection schemes affect model reliability, as well as the process of choosing appropriate configurations for each scheme, remain under-explored, and their significance has yet to be fully recognized.
+Our primary findings reveal that DETR predictions within the same image are interdependent, leading to significantly varying levels of reliability.
+Since DETR is trained using gradients derived exclusively from optimally matched predictions (i.e., the predictions most closely corresponding to each annotated object), it will output at least one well-calibrated prediction for each perceived object while having flexibility in handling the remaining unmatched predictions.
+Throughout the paper, we refer to those matched predictions as positive predictions and the others as negative predictions.
+In principle, to avoid an inaccurate match, DETR could assign those unmatched predictions low confidence with random bounding boxes (Figure 1a).
+Alternatively, it could generate similarly accurate bounding boxes with (nearly) equally calibrated confidence (Figure 1b).
+Otherwise, it may assign poorly calibrated confidence scores (Figure 1c).
+Our analysis indicates that a prediction generated by DETR mostly follows the third scheme.
+Building on this observation, this paper investigates the importance of distinguishing between positive and negative predictions, and addresses the following three research questions: (RQ1) Do all predictions generated for a given image exhibit comparable levels of reliability?
 (RQ2) If not, what is the appropriate way to identify reliable predictions across the entire set?
-and (RQ3) How can we accurately assess DETR’s reliability?
-Our primary findings reveal that predictions from DETR within the same image are interdependent, leading to varying levels of reliability.
-Since DETR is trained using gradients from optimally matched predictions, it has flexibility in handling the remaining predictions.
-In principle, DETR could assign them low confidence with random bounding boxes (Figure 1(a)).
-Alternatively, it could also generate similar bounding boxes with equally calibrated confidence (Figure 1(b)).
-Otherwise, it may assign poorly calibrated confidence scores (Figure 1(c)).
-Our analysis indicates that DETR mostly follows the third scheme; we refer to these well-calibrated predictions as positive predictions and the others as negative predictions.
-Building on this observation, this paper investigates the importance of distinguishing between positive and negative predictions, and how this finding can be leveraged into an uncertainty quantification (UQ) framework for DETR.
-More specifically, our main contributions are:
-  • We offer both qualitative and quantitative insights into how positive and negative predictions influence the model’s object-level calibration.
-Our analysis shows that negative predictions are poorly calibrated.
-Furthermore, their confidence scores are inversely correlated with image-level reliability, highlighting the need for a separation method to ensure DETR’s reliable use.
-  • We raise concerns about existing performance and calibration metrics—including AP [25, 6, 14], detection expected calibration error (D-ECE) [10], and Localisation-aware expected calibration error (LA-ECE) [19, 11]—by noting their inadequacy for identifying the reliable subset of predictions.
-For instance, AP favors retaining more predictions, while ECE prefers a smaller high-confidence set.
-Consequently, the subset of predictions that maximizes performance for each metric differs significantly from the optimal positive set.
-  • To address this, we introduce a new calibration metric, object-level calibration error (OCE), which calculates calibration error along ground truth objects rather than predictions.
-We demonstrate that OCE is capable of both ranking different models by their calibration qualities and identifying the reliable subset of predictions to use.
-  • Based on our findings, we provide a practical implication: we introduce a novel framework for quantifying image-level reliability by contrasting the average confidence scores between the positive and negative predictions identified by our OCE metric.
-We conduct numerical experiments, demonstrating the effectiveness of our approach across in-distribution, near out-of-distribution, and far out-of-distribution scenarios.
-  • We present a comparative analysis of methods for identifying positive predictions in DETR, from the UQ perspective.
-Our results show that thresholding (on predictive confidence score) is a more suitable method than widely used techniques such as top-$k$ or non-maximum suppression (NMS) for obtaining a reliable subset of predictions.
-2 Problem Statement
-2.1 Detection Transformers (DETRs)
-We consider a test image $\bm{x}$ and denote $\mathcal{T}_{\bm{x}}$ as the set of ground truth objects present in the image.
-Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{T}}_{\theta}(\bm{x})$.
-Each prediction $t\in\hat{\mathcal{T}}_{\theta}(\bm{x})$ is characterized by a bounding box $\hat{\bm{b}}$ and an associated class label with a corresponding probability $\hat{p}$.
+and (RQ3) How can we accurately assess DETR’s image-level reliability?
+In addressing these questions, our specific contributions include:
+  1. 1.
+    We provide both qualitative and quantitative insights into how positive and negative predictions influence the model’s reliability.
+Our analysis reveals that, unlike positives, negative predictions are often poorly calibrated and therefore should be handled separately.
+Furthermore, their confidence scores are inversely correlated with image-level reliability, highlighting the need for a separation method to ensure DETR’s reliable use (Section IV).
+  2. 2.
+    Identifying ground-truth positives and negatives ideally requires a matching process based on ground-truth annotations, which are unavailable at test time.
+Consequently, an alternative framework for identifying positive predictions is needed.
+However, we show that existing performance and calibration metrics—such as average precision [16, 17, 18] and expected calibration error [8, 19, 20]—are inadequate for this purpose.
+To address this, we introduce a new calibration metric, object-level calibration error (OCE), which assesses calibration error along ground-truth objects rather than predictions.
+We demonstrate that OCE is more suitable for both ranking different models by their calibration qualities and identifying the positive predictions (Section V).
+  3. 3.
+    Based on our findings, we introduce a novel framework for quantifying image-level reliability by contrasting the average confidence scores between the positive and negative predictions identified by our OCE metric.
+We conduct numerical experiments, demonstrating the effectiveness of our approach across in-distribution, near out-of-distribution, and far out-of-distribution scenarios (Section VI).
+In addition to addressing the three primary research questions, we present a comparative analysis of standard post-processing methods for identifying positive predictions in DETR from an uncertainty quantification (UQ) perspective (Section VII).
+Our results indicate that thresholding on the predictive confidence score is more effective than widely used techniques such as top-$k$ selection or non-maximum suppression (NMS) for obtaining a reliable subset of predictions.
+II Related Work
+Calibration.
+Model calibration refers to how well a model’s predicted confidence scores align with the actual likelihood of those predictions being correct.
+In other words, low-confidence samples should exhibit low accuracy, and high-confidence samples should demonstrate high accuracy.
+For instance, if a model predicts an event with 80% confidence, it should be correct around 80% of the time for that event to be considered well-calibrated.
+Achieving good calibration is crucial because it improves trust in the model’s predictions, guides more informed decision-making, and enables better risk assessment in real-world applications.
+To evaluate the alignment, they measure the expected calibration error by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin.
+However, in object detection tasks, measuring accuracy is not straightforward because predictions comprise both class probabilities and bounding boxes.
+Additionally, due to the set prediction nature of object detection models, it is unclear which ground truth object corresponds to each prediction.
+To address these challenges, D-ECE [8] defines precision as the accuracy metric and matches each prediction to a ground truth object based on an intersection over union (IoU) threshold (commonly set at $0.5$ or $0.75$).
+On the other hand, LaECE [21] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
+However, as we demonstrate, DETR exhibits poor calibration on negative predictions, regardless of its actual reliability.
+In practice, users often focus only on positive samples.
+Yet, when model calibration is evaluated across the entire set of predictions, including the negatives, the resulting evaluation becomes highly biased and therefore unreliable.
+To avoid this issue, previous studies have often measured D-ECE only on predictions with confidence scores exceeding $0.3$.
+As detailed in later sections, a threshold of $0.3$ approximates the optimal positive predictions in some DETR variants; thus, using such a fixed threshold may occasionally be acceptable.
+However, the optimal threshold is not guaranteed to be always $0.3$ for other models like UP-DETR, which introduces potential risks.
+In response, our paper proposes OCE, which measures the model’s calibration quality alongside the employed post-processing scheme and thus can adaptively identify the reliable subset.
+Uncertainty Quantification.
+Several studies focus on out-of-distribution (OOD) identification in object detection models.
+For example, [22] propose a built-in OOD detector to isolate OOD data for human review, including those of unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty), by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions.
+[23] generate outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data.
+Similarly, [19] employ an auxiliary detection model capable of expressing its confidence.
+Other works, including [24] and [25], investigate the latent representations generated by object detection models to identify the OOD nature of the input.
+To the best of our knowledge, the aforementioned existing UQ techniques primarily focus on prediction-level analysis.
+Moreover, they predominantly address CNN-based models and explore the methodological way to better quantify the uncertainty in object detection models.
+In contrast, our paper emphasizes the significance of identifying reliable sets within the entire set of predictions for uncertainty quantification, particularly in DETRs.
+Another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
+III Preliminaries
+Figure 2: A diagram of the DETR architecture.
+An input image is first processed through a CNN backbone to generate a 2D feature representation.
+This representation is then passed to the Transformer encoder, which extracts feature vectors.
+These feature vectors are sent to the decoder, which receives $M$ learned object queries together.
+The decoder outputs $M$ prediction sets, each containing a bounding box and corresponding class probabilities.
+III-A Detection Transformer (DETR)
+We consider a test image $\bm{x}$ and denote $\mathcal{D}_{\bm{x}}$ as the set of ground truth objects present in the image.
+Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{D}}_{\theta}(\bm{x})$.
+Each prediction $d\in\hat{\mathcal{D}}_{\theta}(\bm{x})$ is characterized by a bounding box $\hat{\bm{b}}$ and an associated class label with a corresponding probability $\hat{p}$.
 The structure of DETR is composed of two main components: the Transformer encoder, which extracts a collection of features from the given image; and the Transformer decoder, which uses these features to make predictions.
@@
 Together, these predictions form the final outputs, making DETR’s predictions essentially an $M$-element set.
-We refer to Figure 4 in the Appendix for an illustration.
-The encoder follows the common structure of standard computer vision models and is based on pre-trained models, whose reliability has been relatively widely explored [27, 26, 29, 20].
-This foundation further enables the use of prominent post hoc UQ techniques, such as Monte Carlo dropout [7, 29] and distance-based out-of-distribution (OOD) detection methods [12, 28, 4].
+We refer to Figure 2 for an illustration.
+It is noteworthy that the encoder follows the common structure of standard computer vision models, whose reliability has been relatively widely explored [26, 27, 28, 29].
+This foundation further enables the use of prominent post hoc UQ techniques, such as Monte Carlo dropout [30].
 However, despite the decoder being the predominant component for object detection, there is a gap in understanding and quantifying its reliability due to its unique structural characteristic: set prediction.
-Therefore, this paper delves into the underlying characteristics of these predictions and presents a methodology to quantify the reliability of the decoder in DETR for object detection tasks.
-(a) A prediction in high-reliability image: Horse
-(b) A prediction in low-reliability image: Chair
+Therefore, this paper delves into the underlying characteristics of these predictions and presents a methodology to quantify the reliability of DETR.
+III-B Bipartite Matching
+Since the number of queries in DETR, $M$, is much higher than the number of annotated objects, DETR matches each object with the corresponding best model prediction during its training.
+To compute this optimal matching for the predictions in a given image, a bipartite matching algorithm is applied.
+More specifically, a matching cost between each pair of a given prediction and an object is defined as follows:
+$\mathcal{L}_{matching}=\mathcal{L}_{class}+\mathcal{L}_{box}$ (1)
+where $\mathcal{L}_{class}$ is the negative prediction confidence of the ground truth class and $\mathcal{L}_{box}$ is the linear combination of the $\ell_{1}$ loss between the corners of the bounding boxes and $\mathcal{L}_{iou}$.
+$\mathcal{L}_{iou}$ is the Generalized Intersection over Union (GIoU) [31] loss between bounding boxes.
+After computing this matching cost for every combination of prediction set and ground truth objects, DETR then efficiently calculates the permutation that minimizes the total matching cost using the Hungarian matching algorithm [32, 7].
+III-C Expected Calibration Error (ECE) Metrics
+D-ECE.
+For a given object detection model, detection expected calibration error (D-ECE) [8] quantifies how closely the model’s predicted confidences align with its observed precision.
+Specifically, let $\hat{\mathcal{D}}$ be the set of all detections that the model produces on the validation set.
+Each detection has an associated predicted confidence score $\hat{p}\in[0,1]$.
+To compute D-ECE, the confidence space $[0,1]$ is partitioned into $J$ bins, for instance via uniform intervals $\big{[}0,\tfrac{1}{J}\big{)},\big{[}\tfrac{1}{J},\tfrac{2}{J}\big{)},\dots,\big{[}\tfrac{J-1}{J},1\big{]}$.
+Let $\widehat{\mathcal{D}}_{j}\subseteq\widehat{\mathcal{D}}$ be the subset of detections falling into the $j$-th bin.
+Denote by $\bar{p}_{j}$ the average confidence of detections in bin $j$, and by $\mathrm{precision}(j)$ the empirical precision in bin $j$.
+A detection is considered a true positive if it has the correct predicted class label and its IoU with the ground-truth box exceeds a given threshold $\tau$.
+The empirical precision $\mathrm{precision}(j)$ is thus the fraction of detections in $\widehat{\mathcal{D}}_{j}$ that are true positives.
+D-ECE then aggregates the absolute difference between each bin’s average confidence and observed precision, weighted by bin size, and is defined as:
+[…]
+Following the standard protocol [20], two IoU thresholds ($\tau=0.5$ and $\tau=0.75$) are used, and the D-ECE scores obtained under these thresholds are averaged to yield the final metric.
+Because this process is carried out on a validation set, it provides an overall measure of the model’s calibration performance on that set.
+A lower D-ECE indicates that the confidence scores are better aligned with actual precision, while a higher D-ECE reveals a potential mismatch between confidence and true detection performance.
+LaECE.
+To account for both localization error and classification precision, localisation-aware expected calibration error (LaECE) [19] aligns model confidence with the product of the predicted bounding box’s precision and its IoU with the ground-truth box.
+Specifically, LaECE is computed in a class-wise manner to mitigate class imbalance.
+Let $c$ index the class, and let $\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$.
+Partition the confidence scores into $J$ bins, and let $\widehat{\mathcal{D}}_{j}^{c}\subseteq\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$ in bin $j$.
+Define $\bar{p}^{c}_{j}$ to be the average confidence in $\widehat{\mathcal{D}}_{j}^{c}$, $\mathrm{precision}^{c}(j)$ to be the corresponding precision, and $\bar{\mathrm{IoU}}^{c}(j)$ to be the average IoU of $\widehat{\mathcal{D}}_{j}^{c}$.
+The LaECE metric is then given by:
+$\text{LaECE}\triangleq\frac{1}{K}\sum_{c=1}^{K}\sum_{j=1}^{J}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}_{j}^{c}-\text{precision}^{c}(j)\times\bar{\text{IoU}}^{c}(j)\Bigr{|}$ (3)
+where $K$ is the total number of classes, and an IoU threshold $\tau$ is employed to determine whether a detection is considered a true positive.
+LaECE0 [20] sets the IoU threshold $\tau$ to zero and assigns an average IoU of zero to false positive detections, thereby reducing the original formulation.
+In particular, if a detection is not a true positive under $\tau=0$, its IoU contribution is taken to be zero.
+The resulting metric becomes:
+$\text{LaECE}_{0}=\frac{1}{K}\sum^{K}_{c=1}\sum^{J}_{j=1}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}^{c}_{j}-\bar{\text{IoU}}^{c}(j)\Bigr{|}.$ (4)
+By setting $\tau=0$, false positives have $\bar{\mathrm{IoU}}=0$, and thus the calibration criterion focuses more explicitly on the contribution of bounding box localization to the overall calibration error.
+III-D Image-Level Reliability
+We introduce a formal definition of image-level reliability by examining the model’s overall object detection performance on the image, extending [33, 29].
+Definition 1.
+We define image-level reliability as a measure of how accurately and confidently the predictions match the ground truth objects:
+$\mathsf{ImReli}(\bm{x};\theta)\triangleq\mathsf{Perf}\big{(}\hat{\mathcal{D}}_{\theta}(\bm{x}),~{}\mathcal{D}_{\bm{x}}\big{)}$ (5)
+where one can use any standard performance metrics, such as average precision and recall, for $\mathsf{Perf}$ depending on the user’s requirements.
+By its definition, image-level reliability directly addresses the model’s applicability to a given test instance.
+However, since image-level reliability requires ground truth annotations for its determination, obtaining it during inference is not feasible.
+Therefore, this paper’s objective is to develop a method that assigns a quantitative score to each image instance, closely aligning with the model’s image-level reliability.
+(a) Predictions within High-Reliability Image
+(b) Predictions within Low-Reliability Image
 Figure 3: Visualizations of the predictions generated by Cal-DETR.
-The optimal positive prediction (indexed by 0 and bordered in green) and the five optimal negative predictions (indexed by 1-5 and bordered in red) with the largest IoU are presented.
+The optimal positive prediction (indexed by 0 and bordered in blue) and the five optimal negative predictions (indexed by 1-5 and bordered in red) with the largest IoU are presented.
 For each prediction and layer, the maximum confidence score and its corresponding label are visualized.
-When the model is confident, the confidence score of the positive prediction either increases or remains high across the decoder layers, while those of negative predictions decrease or stay low.
+When the model is confident, the confidence score of the positive prediction either increases or remains high across the decoder layers, while those of negative predictions decrease or stay low despite its accurate bounding box predictions.
 On the other hand, when the model is uncertain, DETR assigns a confidence score to positive prediction based on its confidence, thereby maintaining good calibration.
 However, conversely, it slightly increases or maintains the confidence scores for negative predictions.
-2.2 Identifying Reliable Predictions in DETR
-Notably, the number of predictions generated by DETR, $|\widetilde{\mathcal{T}}_{\theta}(\bm{x})|$, is fixed and often in the hundreds, far exceeding the number of ground truth objects.
-To address the issue, during model training, a bipartite matching algorithm is used to find the optimal matching prediction for each ground truth object based on the alignment of the class label and bounding box (as detailed in Appendix B).
-Consequently, the parameters $\theta$ are optimized to enhance the accuracy of these matched queries.
-In this paper, we refer to the matched queries as optimal positive queries, while the remaining queries are termed optimal negative queries.
-At the inference stage in real-world scenarios, however, ground truth annotations are unavailable, meaning the optimal positive queries remain unknown.
-This raises the question of whether all predictions can be trusted and used without any post-processing, or if only a subset should be selected for the final inference—and, if so, which subset?
-If all of the model’s predictions, including negative queries, were generated independently and were nearly identical to one of the positive predictions, the large number of predictions would not be a concern.
+IV Impact of Negative Predictions on DETR’s Reliability
+IV-A Motivation and Scope
+Notably, the number of predictions generated by DETR, $|\widetilde{\mathcal{D}}_{\theta}(\bm{x})|$, is fixed and often in the hundreds, far exceeding the number of ground truth objects.
+To address this issue during model training, a bipartite matching algorithm is used to find the optimal matching prediction [7] for each ground truth object based on the alignment of the class label and bounding box (as detailed in Section III-B).
+The training loss depends on the quality of this matching, and as a result, the parameters $\theta$ are primarily optimized to enhance the accuracy of these matched predictions.
+In this paper, we refer to the matched predictions as optimal positive predictions, while the remaining predictions are termed optimal negative predictions.
+At the inference stage in real-world scenarios, however, ground truth annotations are unavailable, meaning the optimal positive predictions remain unknown.
+This raises the question of whether all predictions are trustworthy, or should only a specific subset be chosen—and if so, which subset?
+If all of the model’s predictions were generated independently and were well-calibrated (e.g., Figure 1b), the large number of predictions would not be a concern.
 We could simply apply well-known algorithms like NMS to remove duplicates and resolve redundancy.
-Notably, we have observed that DETR assigns well-calibrated confidence scores to only a single positive query (i.e., prediction) per object, while assigning low scores but similar bounding boxes to negative queries (see Section 3.1 and Figure 2).
-As a result, the overall prediction-level calibration error becomes significantly high (See Section 3.2) if we include those noisy predictions in the final inference results.
-Hence, improper handling of negative queries (e.g., by using top-100 predictions) can severely compromise DETR’s reliability, with potentially catastrophic consequences in safety-critical scenarios, such as pedestrian detection in autonomous vehicles.
-For this reason, addressing this challenge and effectively distinguishing the positives from negatives is a crucial task for ensuring reliable UQ in DETR.
-3 Analysis and Key Insights
-3.1 Exploring the Anatomy of DETR’s Predictions
-We begin by closely examining the underlying dynamics of DETR decoders and visualizing the generated outputs.
+Notably, we have observed that DETR assigns well-calibrated confidence scores to only a single positive prediction per object.
+Meanwhile, unmatched negative predictions, which often have accurate bounding boxes, receive uncalibrated low scores (e.g., Figure 1c).
+As a result, the overall calibration error becomes significantly high if we include those noisy predictions in the final inference results.
+Hence, as detailed in the following sections, improper handling of negative predictions can severely compromise DETR’s reliability, with potentially catastrophic consequences in safety-critical scenarios.
+For this reason, properly distinguishing the positives from the negatives is a crucial task for ensuring the reliable use of DETR.
+IV-B Exploring the Anatomy of DETR’s Predictions
+We begin our analysis by visualizing and examining the outputs generated from the DETR decoders.
 Since the Transformer decoder outputs only representation vectors, investigating their evolution across layers is not straightforward.
@@
 This is feasible due to the alignment of intermediate representations, facilitated by residual connections between decoder layers [34].
-Sample visualizations are in Figure 2 along with Figures 5 and 6 in Appendix C.
+Sample visualizations are in Figure 3.
 In the first decoder layer, the model appears to explore the encoded image features, producing varied queries that result in various plausible predictions.
-In this early stage, the distinction between positive and negative queries can be ambiguous (e.g., Figure 2(a)).
+In this early stage, the distinction between positive and negative queries can be ambiguous.
 However, the self-attentions through the subsequent decoder layers progressively refine these predictions.
-By the final layer, the model selects a single query (i.e., optimal positive) and assigns a confidence score based on its understanding of the image and the object.
-In contrast, the confidence scores for neighboring queries (i.e., optimal negatives) do not increase to the same extent as the positives and even decrease in high-reliability images.
-In contrast, in low-reliability images (e.g., Figure 2(b)), the confidence score of the positive query does not significantly increase, while the scores of the negative ones are either slightly raised or unchanged.
+By the final layer, the model selects a single query (i.e., optimal positive prediction) and assigns a confidence score based on its understanding of the image and the object.
+In contrast, the confidence scores for neighboring queries (i.e., optimal negative predictions) do not increase to the same extent as the positives and even decrease in high-reliability images.
+In contrast, in low-reliability images, the confidence score of the positive query does not significantly increase, while the scores of the negative ones are either slightly raised or unchanged.
 Based on this observation, we present our claim:
@@
 Predictions from DETR within a given image exhibit varying levels of reliability.
-For each object in the image, the optimal positive (i.e., best-matched) prediction is calibrated, while the remaining optimal negative predictions are uncalibrated.
-3.2 Analysis of Calibration Error for Positive and Negative Predictions
-Primarily, to address the first research question (RQ1) and provide quantitative support for our claim regarding the varying levels of reliability across predictions, we evaluate and compare the calibration errors on the optimal positive predictions and the optimal negative predictions.
-The optimal positives and negatives are determined using the standard bipartite matching algorithm [1].
-We evaluate the object-level calibration error, which is defined later in Section 4.2.
-From Table 1, we confirm our claim that only the optimal positive predictions are properly-calibrated.
-Furthermore, as illustrated in Figure 2, the optimal negative predictions exhibit a significantly high calibration error, driven by consistently low confidence scores regardless of their actual accuracy.
-It is also noteworthy that even Cal-DETR, the state-of-the-art calibrated DETR variant, demonstrates poor calibration quality for optimal negative predictions.
-Moreover, as will be explained in more detail later, Table 3 further supports our claim by illustrating that distinguishing between positive and negative predictions becomes particularly crucial when applied to image-level UQ in DETR.
-In short, when confidence scores are averaged without distinguishing positive queries (see the row of $\mathsf{Conf}^{+}$ using the top-$100$ predictions), the correlation between the confidence scores and image-level reliability (i.e., precision) becomes negative.
-This is because the average confidence scores for negative queries are shown to be inversely correlated with image-level reliability, and the majority of DETR’s outputs are actually negative.
-These results further highlight the importance of employing an appropriate separation scheme to accurately identify optimal positive predictions, even when ground truth annotations are unavailable.
-Table 1: Object-level calibration error on top-$100$, optimal positive, and negative predictions are reported across various models and datasets.
-For each model, the smallest and largest errors are highlighted in blue and red, respectively.
-Notably, negative queries are uncalibrated.
-Since the top-$100$ predictions often include these negative queries, the resulting calibration error tends to be high.
-Model Predictions COCO City Foggy
-D-DETR Top-100 0.582 0.552 0.618
-Negative 0.803 0.789 0.844
-Positive 0.275 0.341 0.419
-UP-DETR Top-100 0.660 0.689 0.709
-Negative 0.872 0.871 0.883
-Positive 0.438 0.494 0.546
-Cal-DETR Top-100 0.655 0.668 0.688
-Negative 0.867 0.859 0.871
-Positive 0.425 0.463 0.513
-DINO Top-100 0.635 0.693 0.713
-Negative 0.953 0.959 0.962
-Positive 0.411 0.480 0.544
-4 A Systematic Framework for Identifying Positive Predictions
-Identifying the positive predictions is crucial for the reliable use of DETR.
-Nonetheless, identifying optimal positive predictions during the inference stage is not feasible due to the lack of ground truth labels.
+For each object in the image, the optimal positive prediction is calibrated (i.e., reliable), while the remaining optimal negative predictions are not.
+IV-C Numerical Analysis
+(a) Object-level Calibration Error
+(b) Correlation to Image-level Reliability
+Figure 4: A visualization of the difference in calibration between positive and negative predictions on the COCO dataset.
+In Figure 4a, Object-Level Calibration Error (OCE) values of their confidence scores are shown, where a lower score represents better-calibrated predictions.
+In Figure 4b, Pearson Correlation Coefficient (PCC) values between the ground truth $\mathsf{ImReli}$ and the average confidence scores are shown, where a higher value is better.
+Setup.
+To provide quantitative support for our claim regarding the varying levels of reliability across predictions, we conducted experiments using four DETR variants: UP-DETR [35], Deformable-DETR (D-DETR), Cal-DETR [12], and DINO.
+Each model is trained on the COCO (train2017) dataset and we use $1,000$ images (i.e., $20\%$) of COCO (val2017) for the validation set and the remaining images for the test set.
+The model is tested on three datasets with varying levels of out-of-distribution (OOD) characteristics: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD).
+Further details can be found in the Appendix.
+Object-level Calibration Error.
+We evaluate and compare the object-level calibration errors on the optimal positive predictions and the optimal negative predictions.
+The optimal positives and negatives are determined using the previously mentioned bipartite matching algorithm.
+As illustrated in Figure 3, the optimal negative predictions exhibit a significantly high calibration error, driven by consistently low confidence scores regardless of their actual accuracy.
+Furthermore, as shown in Figure 4a, the optimal positive predictions are significantly better calibrated than the negative ones.
+This trend is evident across different state-of-the-art DETR variants and datasets, all of which exhibit poor calibration quality for optimal negative predictions.
+Correlation to Image-Level Reliability.
+We compute $\mathsf{ImReli}$ for each image by using the same COCO evaluator [18] to obtain the image-wise AP score for $\mathsf{Perf}$ in Equation 1, passing the predictions and annotations for each image individually rather than the entire image set.
+Then, we measure the Pearson correlation coefficient (PCC) between the $\mathsf{ImReli}$ and the average confidence score of the optimal positives and the optimal negatives, respectively.
+As shown in Figure 4b, the average confidence scores of the optimal positive predictions exhibit a moderately strong positive correlation with image-level reliability (i.e., the model’s actual average precision on that image) across different models.
+Notably, the average confidence scores for negative predictions are inversely correlated with image-level reliability.
+This finding reinforces our claim, highlighting the importance of distinguishing between positive and negative predictions, especially when conducting image-level UQ in DETR.
+Further visualizations for the other datasets are provided in the Appendix.
+V A Systematic Framework for Identifying Positive Predictions
+As illustrated so far, using positive predictions is crucial for the reliable use of DETR.
+Nonetheless, due to the lack of annotations, it is not feasible to identify either the ground truth or the optimal positive predictions during the inference stage.
 Therefore, an alternative systematic framework is essential not only for improving its interpretability but also for ensuring reliability.
-Given that the optimal matching achieves the smallest calibration error, we use calibration error to measure the quality of a separation scheme, i.e., a separation scheme with a lower calibration error is deemed a better scheme.
-(a) Cal-DETR
-(b) UP-DETR
-Figure 3: Impact of confidence threshold selection on various performance metrics in Cal-DETR and UP-DETR on COCO.
-4.1 Limitations of Existing Metrics
-To this end, this section analyzes the effectiveness of existing metrics, including AP, D-ECE, LA-ECE0, and LRP, for measuring calibration error.
-Specifically, we modify a variable (e.g., the confidence score threshold) to generate different subsets of DETR predictions.
-We then assess the performance of each metric for these subsets and determine the variable value that yields the highest performance.
-Lastly, we evaluate the extent of separation by comparing the predicted positive set to the optimal positive set.
-Sample analysis results are illustrated in Figure 3.
-For additional results and detailed information on these metrics, please refer to Appendix D.
-Primarily, as noted in several studies [18, 19, 11] and our empirical findings, the optimal AP is achieved when the threshold is set to $0.0$; AP does not penalize harshly for having low confidence predictions.
+Given that the optimal matching achieves a small calibration error, we use calibration error to measure the quality of a separation scheme, i.e., a separation scheme with a lower calibration error is deemed a better scheme.
+Figure 5: Impact of confidence threshold selection on various performance and error metrics in Cal-DETR evaluated on COCO.
+Higher scores are preferable ($\uparrow$), while lower scores are preferable ($\downarrow$).
+V-A Limitations of Existing Metrics
+To this end, this section analyzes the effectiveness of existing metrics, including AP, D-ECE, LA-ECE0, and localization recall precision (LRP) [21, 36], for identifying positive predictions.
+Specifically, we sweep the threshold for the confidence score to generate different subsets of DETR predictions (i.e., the predicted positives).
+We then assess the performance of each metric for these subsets and determine the threshold value that yields the highest performance.
+Analytical results using Cal-DETR are illustrated in Figure 5.
+For definitions of AP and LRP and additional results on the other DETR models, please refer to the Appendix.
+Primarily, as noted in several studies [21, 19, 20, 36] and our empirical findings, the optimal AP is achieved when the threshold is set to $0.0$ because AP does not penalize harshly for having low confidence predictions.
 However, as discussed earlier, using the entire prediction set carries a high risk of including uncalibrated negatives, leading to unreliable decisions in practical applications.
-Furthermore, using hundreds of predictions harms the interpretability of the model.
+Furthermore, using hundreds of predictions diminishes the interpretability of the model.
 In contrast, the optimal ECEs are often achieved when the threshold is set close to $1.0$, meaning ECEs favor retaining fewer predictions with high confidence.
 This is a structural pitfall that prediction-level ECEs commonly face [20].
-Since ECEs do not penalize missed detections (i.e., false negatives), they can achieve near-zero error when the evaluation positive set consists solely of highly accurate and confident predictions.
-Therefore, unless the model is trained to be excessively overconfident—which is unlikely given that DETR is trained on large datasets using various auxiliary loss functions—ECEs can result in very small error values when using a large confidence threshold.
-LRP [18], a localization-aware performance metric, can be used instead.
-However, LRP is not designed as a calibration metric and does not explicitly consider calibration error.
-As a result, the ranking scored by LRP across the different models is not necessarily aligned with the model’s calibration qualities.
-See Section 6.1 for the numerical analysis and more discussions.
-4.2 Proposed Metric: Object-Level Calibration Error
+Since ECEs do not penalize missed detections (i.e., false negatives), they could achieve near-zero error when the evaluation positive set consists solely of highly accurate and confident predictions.
+Therefore, unless the model is trained to be excessively overconfident—which is unlikely given that DETR is trained on large datasets using various auxiliary loss functions—ECEs can result in very small error values when paired with a large confidence threshold.
+On the other hand, LRP, a localization-focused performance metric, could be used instead.
+However, LRP is not designed as a calibration metric so does not explicitly consider calibration error.
+As a result, the model ranking scored by LRP across the different models is not necessarily aligned with the model’s calibration qualities, as shown in the following numerical analysis.
+V-B Proposed Metric: Object-Level Calibration Error
 Notation.
-For each image $\bm{x}_{i}$ having $N_{i}$ objects, we consider the ground truth set $\mathcal{T}_{\bm{x}_{i}}=\{o_{i,j}=(l_{i,j},\bm{b}_{i,j})\}_{j=1}^{N_{i}bu}$, and a set of DETR predictions: $\hat{\mathcal{T}}_{\theta}(\bm{x}_{i})=\{t_{i,q}=(\hat{\bm{p}}_{i,q},\hat{\bm{b}}_{i,q})\}_{q=1}^{M}$ where $M$ represents the number of object queries.
+For each image $\bm{x}_{i}$ having $N_{i}$ objects, we consider a set of ground truth object annotations $\mathcal{D}_{i}=\{d_{i,j}=(l_{i,j},\bm{b}_{i,j})\}_{j=1}^{N_{i}}$, and a set of DETR predictions: $\hat{\mathcal{D}}_{\theta}(\bm{x}_{i})=\{\hat{d}_{i,q}=(\hat{\bm{p}}_{i,q},\hat{\bm{b}}_{i,q})\}_{q=1}^{M}$ where $M$ represents the number of object queries.
 Here, $l_{i,j}\in[1,C]$ and $\hat{\bm{p}}_{i,q}\in\{0,1\}^{C}$ represents the ground-truth label and predicted probability distribution over $C$ classes, respectively, while $\bm{b}_{i,j}$ and $\hat{\bm{b}}_{i,q}\in\{0,1\}^{4}$ corresponds to the ground truth and predicted scaled bounding box, respectively.
 Definition 2.
-Consider a subset of predictions, $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\subseteq\hat{\mathcal{T}}_{\theta}(\bm{x}_{i})$, that is generated by post-processing algorithm $\mathcal{S}$ from the entire prediction set: $\hat{\mathcal{S}}_{\theta}=\mathcal{S}\circ\hat{\mathcal{T}}_{\theta}$.
+Consider a subset of predictions, $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\subseteq\hat{\mathcal{D}}_{\theta}(\bm{x}_{i})$, that is generated by post-processing algorithm $\mathcal{A}$ from the entire prediction set: $\hat{\mathcal{S}}_{\theta}=\mathcal{A}\circ\hat{\mathcal{D}}_{\theta}$.
 We define an object-level calibration error (OCE) as the average Brier score per object:
-$\mathsf{OCE}\big{(}\hat{\mathcal{S}}_{\theta};\mathcal{I}\big{)}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(i,j)\in\mathcal{I}}\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});o_{i,j}\big{)}$ (1)
-$\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});o_{i,j}\big{)}=\sum_{c=1}^{C}\big{(}\mathbbm{1}(c=l_{i,j})-\widetilde{p}_{i,j}[c]\big{)}^{2}$ (2)
-$\widetilde{p}_{i,j}[c]=\frac{1}{|\mathcal{Q}_{i,j}|}\sum_{q\in\mathcal{Q}_{i,j}}\hat{\bm{p}}_{i,q}[c]$ (3)
-where $\mathcal{I}=\cup_{i}\{(i,j)\}_{j=1}^{N_{i}}$ is a set of all objects indices and $\hat{\bm{p}}(\cdot)[c]$ outputs the probability of $c$-th class; $\mathcal{Q}_{i,j}$ is a set of query indices that matches to the ground truth object $o_{i,j}$ and we propose two variants:
-$\mathcal{Q}_{i,j}$$\triangleq\{q\mid\mathsf{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\geq\epsilon\}~{}$ $(\mathsf{OCE_{ENS\epsilon}})$ (4)
-$\mathcal{Q}_{i,j}$$\triangleq\{q=\text{argmin}_{q}\mathsf{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\}~{}$ $(\mathsf{OCE_{MAX}})$ (5)
-The difference is that $\mathsf{OCE_{ENS\epsilon}}$ ensembles the overlapping predictions, while $\mathsf{OCE_{MAX}}$ selects the prediction with the best bounding-box matching.
+$\mathsf{OCE}\big{(}\hat{\mathcal{S}}_{\theta};\mathcal{I}\big{)}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(i,j)\in\mathcal{I}}\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});d_{i,j}\big{)}$ (6)
+$\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});d_{i,j}\big{)}=\sum_{c=1}^{C}\big{(}\mathbbm{1}(c=l_{i,j})-\bar{p}_{i,j}[c]\big{)}^{2}$ (7)
+$\bar{p}_{i,j}[c]=\frac{1}{|\mathcal{Q}_{i,j}|}\sum_{q\in\mathcal{Q}_{i,j}}\hat{\bm{p}}_{i,q}[c]$ (8)
+where $\mathcal{I}=\cup_{i}\{(i,j)\}_{j=1}^{N_{i}}$ is a set of all objects indices and $\hat{\bm{p}}(\cdot)[c]$ outputs the probability of $c$-th class; $\mathcal{Q}_{i,j}$ is a set of query indices that matches to the ground truth object $d_{i,j}$ and we propose two variants:
+$\mathcal{Q}_{i,j}$$\triangleq\{q\mid\text{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\geq\delta\}~{}$ $(\mathsf{OCE_{ENS,\delta}})$ (9)
+$\mathcal{Q}_{i,j}$$\triangleq\{q=\text{argmax}_{q}\text{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\}~{}$ $(\mathsf{OCE_{MAX}})$ (10)
+The difference is that $\mathsf{OCE_{ENS,\delta}}$ ensembles the overlapping predictions, while $\mathsf{OCE_{MAX}}$ selects the prediction with the best bounding-box matching.
 When $|\mathcal{Q}_{i,j}|=0$, we consider the predicted probability to be zero, thus the corresponding Brier score is estimated as $1.0$.
-Following [14, 11], we use IoU thresholds of $\epsilon=0.5,0.75$ and report the average score as $\mathsf{OCE_{ENS}}$ or simply $\mathsf{OCE}$.
+Following [18, 20], we use IoU thresholds of $\delta=0.5,0.75$ and report the average score as $\mathsf{OCE_{ENS}}$ or simply $\mathsf{OCE}$.
 The introduced calibration error has two desirable characteristics.
 First, the prediction set achieves the lowest calibration error when the predictions are well-calibrated to the respective closest ground truth objects.
-Second, it penalizes the subset, $\hat{\mathcal{S}}_{\theta}(\cdot)$, that includes missing ground truth objects; this ensures that subsets containing a small set of highly precise predictions are not assigned an artificially low error, unlike D-ECE and LA-ECE metrics.
-Thus, this metric can effectively assess not only whether the given subsets $\{\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\}_{i}$ are reliable, but also whether they comprehensively capture all ground truth objects, providing richer information compare to existing metrics.
-5 Use-Case Implications
-5.1 Image-Level Reliability
-As a main use-case implication of this research, this paper investigates how effective separation impacts the performance of quantifying image-level reliability within the DETR framework.
-Specifically, we first introduce a formal definition of image-level reliability by examining the model’s overall object detection performance on the image.
-Definition 2.
-We define image-level reliability as a measure of how accurately and confidently the predictions match the ground truth objects:
-$\mathsf{ImReli}(\bm{x};\theta)\triangleq\mathsf{Perf}\big{(}\hat{\mathcal{T}}_{\theta}(\bm{x}),~{}\mathcal{T}_{\bm{x}}\big{)}.$ (6)
-where one can quantify $\mathsf{Perf}$ using any standard performance metrics such as average precision, depending on the specification.
-By its definition, image-level reliability directly addresses the model’s applicability to a given test instance.
-However, since image-level reliability requires ground truth annotations for its determination, obtaining it during inference is not feasible.
-Therefore, our goal is to employ a proxy method that quantifies uncertainty or confidence, which should closely correlate with image-level reliability.
-5.2 Proposed Framework: Quantifying Reliability by Contrasting
-As illustrated, positive and negative queries exhibit varying levels of reliability.
-Interestingly, having queries with low confidence scores does not necessarily imply low reliability.
-Our empirical observations show that confidence scores in negative queries actually inversely correlate with image-level reliability.
-More specifically, for a reliable instance, we observe that the confidence of positive queries increases across the decoder layers, while that of negative queries remains low; this results in a large gap between positive and negative queries.
-In contrast, for unreliable instances (e.g., Figure 2(a)), the confidence score of the positive query does not increase across the layers, whereas the scores of negative queries are either slightly elevated or remain unchanged.
-Consequently, there is a small gap between positive and negative queries (e.g., Figure 2(b)).
-Based on the finding, we propose a post hoc UQ approach by contrasting the confidence scores of positives and negatives:
-$\mathsf{ContrastiveConf}(\bm{x})$ $=\mathsf{Conf}^{+}(\bm{x})-\lambda\mathsf{Conf}^{-}(\bm{x})$ (7)
-$\mathsf{Conf}^{+}(\bm{x})$ $=\frac{1}{|\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})|}\sum\nolimits_{t\in\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})}\mathsf{c}(t)$ (8)
-$\mathsf{Conf}^{-}(\bm{x})$ $=\frac{1}{|\hat{\mathcal{T}}_{\theta}^{-}(\bm{x})|}\sum\nolimits_{t\in\hat{\mathcal{T}}_{\theta}^{-}(\bm{x})}\mathsf{c}(t)$ (9)
-where $\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})$ and $\hat{\mathcal{T}}_{\theta}^{-}(\bm{x})$ are predicted sets of positive and negative predictions, respectively, and $\lambda$ is a scaling factor, that can be selected from the validation set.
-We include an ablation study of the scaling factor in the supplementary material.
-Our results show that the proposed method is not too sensitive to the choice of scaling factor and consistently outperforms baselines, with a scaling factor of $10.0$ providing adequate performance.
-$\mathsf{c}(\cdot)$ denotes the confidence estimate for the prediction; in this paper, we use maximum probability, $\mathsf{c}(t=(\hat{\bm{p}},\hat{\bm{b}}))=\max_{c}\hat{\bm{p}}[c]$.
-As shown so far, identifying a $\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})$ from $\hat{\mathcal{T}}_{\theta}(\bm{x})$ is a crucial factor in the success of this approach.
-In practice, however, neither the ground truth (i.e., the queries that DETR intends to use) nor the optimal positives are available during the UQ stage.
-Instead, we approximate the ground truth separation by applying a post-processing algorithm that minimizes the calibration error on the validation set:
-$\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})=\mathcal{S}^{*}\circ\hat{\mathcal{T}}_{\theta}(\bm{x})$ (10)
-$\mathcal{S}^{*}=\text{argmin}_{\mathcal{S}}\mathsf{OCE}(\mathcal{S}\circ\hat{\mathcal{T}}_{\theta};\mathcal{I}_{val})$ (11)
-where $\mathcal{I}_{val}=\cup_{i\in\mathcal{V}}\{(i,j)\}_{j=1}^{O_{i}}$ is the set of all objects indices in the validation dataset $\mathcal{V}$.
-Unless specified otherwise, we adopt confidence thresholding for the separation scheme (i.e., post-processing algorithm), where the corresponding threshold is the variable to optimize.
-Table 2: Comparison of the effectiveness of the proposed OCE and baseline metrics, focusing on their ranking correlation with the rankings given by the models’ calibration quality on optimal positive predictions (denoted by the superscript *).
-For each metric, various post-processing schemes (i.e., ways to identify positive predictions) are applied and indicated within parentheses.
+Second, it penalizes the subset, $\hat{\mathcal{S}}_{\theta}(\cdot)$, that includes missing ground truth objects.
+This is achievable due to the primary difference between our OCE and ECEs: while ECEs are evaluated based on selected predictions, OCE is assessed along the ground truth objects.
+This ensures that subsets containing a small set of highly precise predictions are not assigned an artificially low error, unlike D-ECE and LA-ECE metrics.
+Thus, this metric can effectively assess not only whether the given subsets $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})$ are reliable, but also whether they comprehensively capture all ground truth objects, providing richer information compared to existing metrics.
+V-C Numerical Analysis: Effectiveness of OCE
+Table 1: Comparison of the effectiveness of the proposed OCE and exisiting metrics, focusing on their correlation with the models’ overall calibration quality on optimal positive predictions (denoted by the superscript *).
+For each metric, different post-processing schemes (i.e., methods for identifying positive predictions) are applied and specified in parentheses.
+These include using the top 100 predictions or choosing predictions with confidence scores above the threshold that optimizes the respective metric.
 Correlations exceeding $0.9$ are highlighted in blue, while negative correlations are highlighted in red.
-Our OCE-based rankings consistently provide strong correlations with optimal ranking and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.
-Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD) Average
-[…]
-References D-ECE 1.000 0.995 0.969 1.000 0.996 0.993 1.000 0.873 0.996 0.980
-[…]
-$\mathsf{OCE_{ENS{}}}^{*}$ 0.969 0.948 1.000 0.993 0.990 1.000 0.996 0.836 1.000 0.970
-Baselines AP (Top-100) [6, 14] 0.552 0.583 0.343 0.779 0.784 0.700 0.597 0.843 0.523 0.634
-LRP (LRP) [18] 0.103 0.051 0.338 -0.268 -0.272 -0.150 -0.257 -0.610 -0.170 -0.137
-D-ECE (Top-100) -0.873 -0.833 -0.966 0.608 0.576 0.511 -0.254 0.250 -0.316 -0.144
-D-ECE (0.3) [10] 0.905 0.939 0.827 0.982 0.967 0.990 0.964 0.799 0.981 0.928
-D-ECE (LRP) 0.802 0.855 0.659 0.987 0.984 0.999 0.983 0.793 0.996 0.895
-D-ECE ($\mathsf{OCE_{ENS}}$) 0.856 0.901 0.727 0.987 0.983 0.999 0.983 0.793 0.996 0.914
-LA-ECE0 (Top-100) -0.944 -0.913 -0.994 -0.431 -0.482 -0.517 -0.786 -0.386 -0.822 -0.697
-LA-ECE0 (0.3) -0.789 -0.729 -0.842 -0.780 -0.818 -0.830 -0.451 -0.041 -0.468 -0.639
-LA-ECE0 (LRP) [11] -0.908 -0.864 -0.948 -0.833 -0.877 -0.846 -0.901 -0.597 -0.913 -0.854
-LA-ECE0 ($\mathsf{OCE_{ENS}}$) -0.901 -0.855 -0.943 -0.723 -0.780 -0.739 -0.884 -0.568 -0.897 -0.810
-Ours $\mathsf{OCE_{ENS}}$ ($\mathsf{OCE_{ENS}}$) 0.962 0.940 1.000 0.987 0.983 0.999 0.964 0.737 0.984 0.951
-$\mathsf{OCE_{MAX}}$ ($\mathsf{OCE_{MAX}}$) 0.954 0.929 0.998 0.954 0.958 0.983 0.955 0.697 0.976 0.934
-6 Experiments
-To demonstrate the effectiveness of the proposed method, we conducted experiments using four DETR variants: UP-DETR [3], Deformable-DETR (D-DETR), Cal-DETR [17], and DINO.
-Each model is trained on the COCO (train2017) dataset and we use $1,000$ images (i.e., $20\%$) of COCO (val2017) for the validation set and the remaining images for the test set.
-The model is tested on three datasets with varying levels of out-of-distribution (OOD) characteristics: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD).
-Experimental details can be found in Appendix E.
-6.1 Effectiveness of OCE
-To showcase OCE’s effectiveness in assessing calibration quality, we analyze how different metrics rank calibration error across four DETR models.
-As baselines, we measure ECEs using the top-$100$ detections (best AP), a standard confidence threshold of 0.3, and the threshold that minimizes LRP on the validation set.
-In a similar manner, we assess our OCEs using thresholds that minimize each OCE metric on the validation set.
-For fair comparison, we execute the optimal matching process on each test dataset to obtain the optimal positive set.
-We then evaluate D-ECE, LA-ECE0, and OCE using these sets as reference scores and calculate the Pearson correlation coefficient to the other scoring schemes across different datasets.
-As shown in Table 2, our OCE metrics maintain a strong correlation with the optimal ECE metrics across different datasets.
-This evidence supports the reliability of our OCEs in ranking models by their overall calibration quality on positive sets—which downstream applications are more likely to utilize instead of the top 100 predictions.
-Notably, LA-ECE0 with a non-optimal positive set shows an inverse correlation with the rankings estimated on the optimal positive set.
-Meanwhile, from a model ranking perspective, empirical evidence suggests that solely adopting LRP does not align with calibration qualities.
-Nonetheless, the ranking by D-ECE and OCE scores measured at the optimal LRP threshold — denoted by D-ECE (LRP) and OCE (LRP), respectively — show a strong correlation to those ranked from the optimal positive set; thus, the method of using LRP to identify a positive set [11] appears acceptable.
-Experimentally, we confirmed that OCE and LRP provide thresholds in a similar range across different settings.
-Another important attribute of OCE is that it can not only be used to compare different models but is also capable of identifying the most reliable configuration of the same model.
-As illustrated in Figure 3, the confidence vs.
-OCE curve exhibits a bell shape, unlike ECE metrics, and the optimal OCE is achieved around a threshold of $0.3$.
-This is particularly notable because it reflects the practical choices in many calibration studies to date [17, 11], affirming the reliability of using OCE.
-Using a fixed separation threshold of $0.3$ is often a good approximation for distinguishing between optimal positives and negatives (See further discussion in Appendix G).
-However, trained or post-calibrated DETR models may have varying optimal confidence thresholds (for example, low-temperature calibrated models are likely to have higher optimal thresholds).
-Therefore, using a fixed threshold can pose a potential risk.
-The biggest advantage of using OCE over other metrics is that OCE measures the model’s calibration quality alongside the post-processing scheme employed.
-While measuring the calibration error on the optimal positive set would ideally assess quality, this optimal set is not accessible during test-time inference.
-Consequently, the score measured on the optimal set may not correspond to the quality experienced in real-world applications, such as those employing confidence thresholding schemes.
-Table 3: Comparison of the proposed method ($\mathsf{ContrastiveConf}$) with baseline methods on their Pearson correlation coefficient with the image-level reliability in Equation (6), with different models.
-$\mathsf{Perf}$ is evaluated using AP (Average Precision)[14].
-For each model, we apply the optimal matching, the standard post-processing scheme (Top-$100$ and confidence thresholding by $0.3$), and the proposed OCE-based post-processing scheme, as indicated within the parentheses.
-The strongest correlations are highlighted in bold, while negative correlations appear in red.
-First, the proposed method achieves the highest correlation with image-level reliability.
-In addition, it is noteworthy that the average confidence of negative predictions ($\mathsf{Conf}^{-}$) exhibits a negative correlation with image-level reliability, in contrast to the positive correlation observed when using positive predictions ($\mathsf{Conf}^{+}$).
-Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
-UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
-Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.503 0.618 0.670 0.635 0.555 0.647 0.649 0.633 0.561 0.642 0.662 0.634
-$\mathsf{Conf}^{-}$ (Optimal) -0.608 -0.584 -0.572 -0.586 -0.293 -0.296 -0.330 -0.311 -0.177 -0.235 -0.212 -0.196
-$\mathsf{ContrastiveConf}$ (Optimal) 0.700 0.684 0.709 0.695 0.601 0.633 0.662 0.634 0.612 0.657 0.672 0.643
-Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.619 -0.603 -0.581 -0.601 -0.409 -0.374 -0.372 -0.391 -0.262 -0.270 -0.239 -0.229
-$\mathsf{Conf}^{+}$ (0.3) 0.476 0.464 0.539 0.504 0.229 0.353 0.385 0.399 0.195 0.256 0.258 0.258
-Ours $\mathsf{Conf}^{+}$ (OCE) 0.442 0.477 0.547 0.512 0.237 0.355 0.393 0.370 0.228 0.291 0.270 0.274
-$\mathsf{Conf}^{-}$ (OCE) -0.633 -0.581 -0.571 -0.585 -0.365 -0.389 -0.379 -0.394 -0.245 -0.325 -0.253 -0.283
-$\mathsf{ContrastiveConf}$ (OCE) 0.585 0.570 0.597 0.587 0.334 0.432 0.431 0.438 0.287 0.362 0.302 0.327
-6.2 Quantifying Image-Level Reliability
-To address the third research question (RQ3), we compare different methods for quantifying per-image reliability using different separation techniques.
-We measure the Pearson correlation coefficient (PCC) of each method with the $\mathsf{ImReli}$ computed based on AP metrics.
-Additionally, we evaluate the proposed contrasting approach against methods that utilize only positive or negative queries.
-Empirical results are presented in Table 3 and further in Appendix F.
-The key takeaways are as follows:
-First, as demonstrated throughout our analysis, distinguishing between positive and negative queries is crucial for uncertainty quantification in DETR.
-Specifically, when confidence scores are averaged without distinguishing positive queries (i.e., $\mathsf{Conf}^{+}$ w/ top-$100$), the correlation becomes negative.
-This occurs because the majority of DETR’s predictions are actually negative (i.e., $|\mathcal{T}_{\bm{x}}|\ll|\hat{\mathcal{T}}_{\theta}(\bm{x})|$), and that confidence scores for negative queries are inversely correlated with $\mathsf{ImReli}$.
-Second, the proposed $\mathsf{ContrastiveConf}$ consistently achieves the best correlation with image-level reliability.
-Interestingly, the absolute value of $\mathsf{Conf}^{-}$ occasionally outperforms $\mathsf{ContrastiveConf}$ when a non-optimal separation scheme is applied; however, its performance varies across different settings.
-This variability underscores the robustness of our contrasting approach, which leverages the strengths of both $\mathsf{Conf}^{+}$ and $\mathsf{Conf}^{-}$.
-Table 4: Comparison of different post-processing schemes on OCE (lower is better) and PCC (higher is better) between $\mathsf{ContrastiveConf}$ and image-level reliability in Cal-DETR.
-The thresholding approach consistently yields the best performance.
-Schemes COCO City Foggy
-OCE PCC OCE PCC OCE PCC
-Optimal 0.425 0.695 0.463 0.634 0.513 0.643
-Thr 0.454 0.594 0.448 0.464 0.482 0.325
-Top-$k$ 0.502 0.527 0.453 0.411 0.492 0.298
-NMS 0.555 0.574 0.582 0.373 0.598 0.239
-6.3 Comparative Study on Post-Processing
-Lastly, we present a comparative analysis across different post-processing schemes for identifying positive queries in DETR, from the reliability perspective.
-We assess OCE for each predicted positive to evaluate the impact of various separation schemes on calibration performance.
-For the separation methods, we compare the following approaches: (1) applying a threshold on the confidence score, (2) selecting the top-$k$ predictions, and (3) utilizing NMS.
-For each method, we evaluate OCE by varying the hyperparameter (e.g., confidence threshold, $k$, and IoU threshold, respectively) and compare with the score on the optimal positive.
-Likewise, we assess the correlation to image-level reliability for each separation scheme.
-For each scheme, the best hyperparameter is chosen using the validation set.
-A summary of the results is provided in Table 4, with the comprehensive results available in Appendix G.
-The thresholding approach achieves the smallest OCE and the best image-level UQ performance.
-In addition, the resulting OCE and correlation is the most comparable to the optimal score, especially for the in-distribution test dataset.
-The top-$k$ approach underperforms because the number of objects in an image varies significantly, making it impractical to determine a single optimal value for $k$.
-The best $k$ ranges from $10$ to $20$, corresponding to the average number of objects per image in the COCO dataset.
-On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive queries are included, a substantial number of negative queries are retained, often outnumbering the positive ones by several times.
-Nonetheless, a performance gap remains compared to optimal matching, especially on out-of-distribution datasets.
-Addressing this gap is left for future work.
-7 Conclusion and Limitations
-The main contribution of our work is an in-depth analysis of the importance of post-processing in DETR frameworks.
-Furthermore, this paper demonstrates the impact of the post-separation method on the reliability of their predictions, and proposes a novel UQ method for quantifying image-level reliability in DETR.
-The primary limitation of our work is that our analysis is confined to standard separation schemes, such as thresholding, without proposing a novel separation technique.
-We conjecture that leveraging information from intermediate layers, rather than solely relying on final predictions, can achieve more robust separation because positive and negative queries follow different evolutionary patterns.
-One potential approach is to enable the model to explicitly output matching probabilities by introducing an auxiliary loss during training, as proposed by Oksuz et al.
-[19].
-Acknowledgments
-The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper.
-This work was supported in part by the MIT Summer Research Program (MSRP), the MIT-Google Program for Computing Innovation, the MIT-Amazon Science Hub, the MIT-IBM Watson AI Lab, and MathWorks.
-Appendix A Related Work
-Uncertainty Quantification.
-Shelmanov et al.
-[27] and Vazhentsev et al.
-[29] focus on uncertainty quantification in general Transformer architectures, building upon the MC dropout approach proposed by Gal and Ghahramani [7].
-Specifically, they apply a modified MC dropout that prioritizes sample diversity over randomness to improve sample efficiency.
-However, these methods are not directly applicable to object detection problems because, unlike classification tasks, object detection involves predicting an unordered set.
-Consequently, it is questionable whether applying MC dropout to each output is appropriate, as the permutation of predictions could be completely shuffled after applying MC dropout.
-Several studies focus on out-of-distribution (OOD) identification in object detection models.
-For example, Li et al.
-[13] propose a built-in OOD detector to isolate OOD data for human review.
-This approach can determine OOD scenarios involving both unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty) by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions.
-Du et al.
-[5] generate outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data.
-Similarly, Oksuz et al.
-[19] employ an auxiliary detection model capable of expressing its confidence.
-Other works, including Du et al.
-[4] and Wilson et al.
-[30], investigate the latent representations generated by object detection models to identify the OOD nature of the input.
-To the best of our knowledge, the aforementioned existing uncertainty quantification techniques primarily focus on prediction-level analysis.
-Moreover, they predominantly address CNN-based models and explore the methodological way to better quantify the uncertainty in object detection model.
-In contrast, our paper emphasizes the significance of identifying reliable sets within the entire set of predictions for uncertainty quantification, particularly in DETRs.
-Another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
-Calibration.
-Previous studies primarily addressed OOD detection to assess the prediction-level trustworthiness of DETR.
-They aim to introduce a framework that does not rely on the model’s confidence scores, which can be either overconfident or underconfident, especially for OOD data points.
-Simultaneously, additional research endeavors to determine the model’s overall reliability in its confidence scores.
-More specifically, these studies assess whether the model’s confidence scores align with their actual accuracy.
-In other words, low-confidence samples should exhibit low accuracy, and high-confidence samples should demonstrate high accuracy.
-To evaluate this alignment, they measure the Expected Calibration Error (ECE) by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin.
-However, in object detection tasks, measuring accuracy is not straightforward because predictions comprise both class probabilities and bounding boxes.
-Additionally, due to the set prediction nature of object detection models, it is unclear which ground truth object corresponds to each prediction.
-To address these challenges, D-ECE [10] defines precision as the accuracy metric and matches each prediction to a ground truth object based on an IoU threshold (commonly set at 0.5 or 0.75).
-On the other hand, LaECE [18] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
-However, as demonstrated throughout our paper, DETR exhibits poor calibration on optimal negative predictions, while in practice, users prefer to utilize only predicted positive samples (e.g., high-confidence predictions).
-Therefore, if we rank the model’s calibration quality based on the entire prediction set, these rankings are highly biased and thus unreliable.
-To avoid this issue, previous studies have often measured D-ECE only on predictions with confidence scores exceeding $0.3$.
-As detailed in Appendix G, a threshold of $0.3$ approximates the optimal positive predictions in many DETR variants; thus, using such a fixed threshold may be practically acceptable.
-However, this threshold is not guaranteed to be $0.3$ for other models like UP-DETR, which introduces potential risks.
-In response, our paper addresses the importance of identifying a reliable subset of predictions and proposes the Object-level Calibration Error (OCE) metric, which measures the model’s calibration quality alongside the employed post-processing scheme and thus can adaptively identify the reliable subset.
-Broad Implications.
-The primary objective of this study is to determine how DETR can be used in a trustworthy manner on new downstream images/tasks and to establish how their predictions should be properly used, which has several important implications.
-For example, when building an auto-labeling system using data-driven detection models like DETR, it is crucial to consider that the model may not always provide accurate predictions.
-As a result, when the model’s reliability for a specific image is in question, the image might need to be reviewed by a human labeler.
-Therefore, evaluating reliability at the image level is often necessary to assess the model’s overall understanding of the given test image.
-Furthermore, by providing a comprehensive analysis for identifying a reliable subset of predictions, our study offers users a systematic and trustworthy method for post-processing the outputs of DETR.
-We hope our efforts expand the scope of DETR applications by enabling more precise and reliable deployment, sparking further research into the area of positive prediction identification.
-Figure 4: A diagram of the DETR architecture.
-An input image is first processed through a CNN backbone to generate a 2D feature representation.
-This representation is then passed to the Transformer encoder, which extracts feature vectors.
-These feature vectors are sent to the decoder, which receives $M$ learned object queries together.
-The decoder outputs $M$ prediction sets, each containing a bounding box and corresponding class probabilities.
-Please refer to the original papers for details.
-Appendix B Bipartite Matching: Positive & Negative Queries
-Since the number of queries in DETR, $M$, is much higher than the number of ground truth objects, DETR matches each ground truth object with the corresponding best model prediction during its training.
-To compute this optimal (i.e., ground truth) matching for the predictions in a given image, a bipartite matching algorithm is applied.
-More specifically, a matching cost between each pair of a given prediction and a ground truth object is defined as follows:
-$\mathcal{L}_{matching}=\mathcal{L}_{class}+\mathcal{L}_{box}$ (12)
-where $\mathcal{L}_{class}$ is the negative prediction confidence of the ground truth class and $\mathcal{L}_{box}$ is the linear combination of the $\ell_{1}$ loss between the corners of the bounding boxes and $\mathcal{L}_{iou}$.
-$\mathcal{L}_{iou}$ is the Generalized Intersection over Union (GIoU) [24] loss between bounding boxes.
-After computing this matching cost for every combination of prediction set and ground truth objects, DETR then efficiently calculates the permutation that minimizes the total matching cost using the Hungarian matching algorithm.
-In this paper, we then refer to those matched queries as positive predictions and the remaining unmatched queries as negative predictions.
-More details on the bipartite matching process and Hungarian algorithm can be found in [1].
-Appendix C Further Visualizations of How DETR Generates Predictions
-The predictions across all six decoder layers for each of the images presented in the main paper are provided in Figure 5.
-Another noticeable example is provided in Figure 6, where predictions for two different objects, a dog and a laptop, are shown for the same low-reliability image.
-In Figure 6(a), for the laptop object, all of the queries start off with low confidence and remain that way over the course of the layers.
-However, the class predictions gradually shift from dining table and person to laptop, and result in an overall low confidence score despite a correct final prediction (laptop).
-However, Figure 6(b) indicates that even in a low-reliability image there can still be reliable predictions.
-For the dog object, the model is very confident in its predictions from the start, with many overlapping prediction sets in the same area and with matching classes.
-While there are two dominant predictions in the first layer, DETR focuses on a single prediction (query 0) while the confidence for query 1 is gradually reduced; this aligns with the trend observed in the reliable examples shown earlier.
-Appendix D Performance Metrics
-D.1 Expected Calibration Error (ECE) Metrics
-Detection Expected Calibration Error (D-ECE) [10] is a calibration metric that specifically focuses on calibrating the confidence of an object detection model to align with precision.
-The confidence space is divided into $J$ bins, and then D-ECE is computed as follows:
-[…]
-where $\hat{\mathcal{D}}$ and $\hat{\mathcal{D}}_{j}$ are the set of all detections and the detections in the $j$-th bin, and $\bar{p}_{j}$ and precision$(j)$ are the average confidence and the precision of the detections in the $j$-th bin.
-D-ECE identifies true positive (TP) samples by confirming the accuracy of class labels and ensuring that the IoU surpasses the threshold $\tau$.
-Following the standard protocol [11], this work measures D-ECE using thresholds $\tau=0.5$ and $\tau=0.75$, and then averages the scores.
-To account for both localization error and classification precision, LaECE [19, 11] calibrates model confidence to the product of precision and the IoU between the bounding box predictions and ground-truth boxes.
-LaECE is computed in a class-wise manner to combat class imbalance, as follows:
-$\text{LaECE}\triangleq\frac{1}{K}\sum_{c=1}^{K}\sum_{j=1}^{J}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}|\bar{p}_{j}^{c}-\text{precision}^{c}(j)\times\bar{\text{IoU}}^{c}(j)|$ (14)
-where $c$ refers to each class and $\bar{\text{IoU}}^{c}(j)$ refers to the average IoU of $\mathcal{\hat{D}}^{c}_{j}$.
-Similar to D-ECE, LaECE requires an IoU threshold, $\tau$.
-In contrast, LaECE0 proposes to set this threshold $\tau$ to 00 and instead set $\bar{\text{IoU}}^{c}(\cdot)=0$ for the false positive (FP) samples.
-This leads to a reduction of the calibration criterion of LaECE as follows:
-$\text{LaECE}_{0}=\frac{1}{K}\sum^{K}_{c=1}\sum^{J}_{j=1}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}|\bar{p}^{c}_{j}-\bar{\text{IoU}}^{c}(j)|$ (15)
-D.2 Performance Metrics
-Average precision (AP) [6] is the most common accuracy metric for object detectors.
-Before calculating AP, the top-$k$ (e.g., typically $100$) class predictions for a given image are kept and sorted, while the others are discarded.
-Then, in order of this sorting, the predictions are compared against against the ground-truth bounding boxes, and the predicted label is compared with the ground truth label if the IoU is large enough.
-The precision/recall curve is calculated over the sorted predictions, and AP averages the precision across a set of 101 evenly spaced recall thresholds:
-$\text{AP}=\frac{1}{101}\sum_{r\in[0:0.01:1]}p(r)$ (16)
-where $p(r)$ is calculated by the maximum precision at recall level $r$.
-[…]
-$\text{LRP}_{\tau}=\frac{1}{N_{\text{FP}}+N_{\text{FN}}+N_{\text{TP}}}\left(N_{\text{FP}}+N_{\text{FN}}+\sum_{i=1}^{N_{\text{TP}}}\frac{1-\text{IoU}_{i}}{1-\tau}\right)$ (17)
-where $\tau$ is IoU threshold, $i$ denotes the index of each true positive prediction, and IoUi represents the IoU between that prediction and the best-matching ground-truth object.
-LRP is then computed by averaging LRP0.5 and LRP0.75.
-D.3 Impact of Separation Schemes on Metrics
-The trends presented in Figure 3 also hold for other datasets and models.
-Figure 7 shows the performance of the model according to different metrics with various confidence thresholds (i.e.
-post-processing schemes).
-As illustrated in Section 4.1, the optimal AP is achieved when the threshold is set to $0.0$.
-In contrast, the optimal ECEs are often achieved when the threshold is set close to $1.0$, meaning ECEs favor retaining fewer predictions with high confidence.
-However, as discussed earlier, using the entire or small subset carries a high risk of including uncalibrated negatives or having missing detections, leading to unreliable decisions in practical applications and diminished interpretability.
-Appendix E Experiment Setup
-For our experiments, we used the Cityscapes and Foggy Cityscapes datasets, which each have 500 images of first-person driving footage in realistic environments.
-Foggy Cityscapes has the same base images as Cityscapes, but with fog simulated and added to create a further out of distribution set.
-Since the DETR models were trained on COCO, the Cityscapes and Foggy Cityscapes annotations were converted to correspond to the labels of COCO.
-More specifically, the person, bicycle, car, motorcycle, bus, train, and truck classes were transferred directly.
-In addition, the rider class of Cityscapes and Foggy Cityscapes was mapped to the person class of COCO.
-The other classes present in Cityscapes and Foggy Cityscapes are largely focused on image segmentation, and thus were omitted (e.g.
-building, sky, sidewalk).
-The pre-trained model weights were obtained from their respective official implementations.
-To compute $\mathsf{ImReli}$ for each image, we use the same COCO evaluator [14] to obtain image-wise AP scores.
-However, rather than passing in the entire image set, the predictions for each image are passed in individually.
-Appendix F Further Experimental Results
-F.1 Effectiveness of OCE
-Table 5 presents the correlations between the rankings provided by three different calibration metrics on the optimal positive set and various methods.
-The first three rows demonstrate that the calibration metrics—D-ECE, LA-ECE, and our proposed OCE—–are highly correlated.
-This indicates that all three metrics effectively capture the notion of calibration quality.
-Additionally, these metrics show a decent correlation with AP metrics.
-It is important to note that AP primarily accounts for accuracy rather than calibration quality, which may explain why the correlation, particularly for AP@50 on the out-of-distribution dataset, is not perfectly aligned.
-Furthermore, LRP, D-ECE, and LA-ECE0 exhibit negative correlations with calibration quality when used alone.
-This supports our assertion that these metrics are inadequate for measuring models’ calibration quality in conjunction with the employed post-processing scheme, unlike our OCE.
-Finally, D-ECE exhibits a high correlation when paired with either a fixed threshold or a proper adaptive threshold (i.e., using LRP and our OCE).
-This is because DETR’s optimal positive threshold is empirically found to be around $0.3$ (See Table 7.).
-Consequently, OCE aligns well with the experimentally approximated threshold and the LRP metric in identifying effective post-processing schemes.
-This alignment further supports the robustness and effectiveness of our method.
-F.2 Quantifying Image-Level Reliability
-This section provides an extended version of Table 3 with more methods such as the proposed method but with a fixed threshold.
-Furthermore, we show more results with different $\mathsf{Perf}$ metrics by evaluating it with AP50 and AP75 (average precision with Iou threshold 50% and 75%).
-The results are shown in Table 6.
-As illustrated, the observation is consistent across different settings.
-Another notable observation is the comparison between the fixed threshold versus our adaptive threshold selection using OCE.
-Although the proposed fixed-threshold approach ($0.3$) occasionally outperforms the adaptive method, the performance gap remains marginal.
-Moreover, the fixed threshold consistently exhibits a lower correlation with UP-DETR, for which the optimal threshold has been empirically determined to be approximately $0.5$ ($0.45$ by OCE and $0.55$ by LRP).
-This indicates that although the OCE approach may not achieve threshold selection that best approximates the optimal positives, it ensures the robustness of uncertainty quantification performance.
-In contrast, fixed thresholding can degrade performance when the chosen value deviates from the optimal threshold.
-Finally, we perform an ablation study on the scaling factor ($\lambda$), with the results presented in Figure 8.
-The study shows that the optimal scaling factor lies within the range of $5.0$ to $10.0$.
-Notably, this range remains effective even for out-of-distribution datasets, such as Cityscapes and Foggy Cityscapes.
-This consistency suggests that the scaling factor selected using the validation set (e.g., COCO) generalizes well across datasets.
-Furthermore, it demonstrates that $\mathsf{ContrastiveConf}$ with $\lambda>0.0$ consistently outperforms $\mathsf{Conf}^{+}$ (i.e., $\lambda=0.0$).
-Appendix G Comparative Study on Post-Processing
-In this section, we provide a comprehensive analysis regarding the impact of post-processing on the model’s overall calibration quality.
-First, we vary the hyperparameter of the standard post-processing processes.
-In this first analysis, to exclude the dependency of the final performance on OCE, we don’t use OCE as a selection criterion.
-Rather, we choose the best hyperparameter on the validation set for each setting and compare their best-possible performances.
-The results are shown in Table 7 .
-As illustrated, we can observe that the confidence thresholding approach outperforms the top-$k$ and NMS approaches.
-Top-$k$ achieves the minimum OCE with top-$20$.
-Given that the average and 95th percentile number of objects per image in the COCO dataset are 7 and 22, respectively, these results appear reasonable.
-Therefore, we confirm that using an excessively large number (e.g., 100) for top-$k$ is inadequate for achieving well-calibrated predictions.
-Similarly, we compare the image-level uncertainty quantification performance, and the results are shown in Table 8.
-We reconfirm the effectiveness of confidence thresholding; for Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies.
-For UP-DETR, the optimal thresholds exceed $0.5$, highlighting the potential limitation of using a fixed threshold in image-level UQ applications.
-Moreover, the correlation is negative especially when NMS is applied along with the $\mathsf{Conf}^{+}$ framework.
-This is because, while most of the optimal positive queries are likely to be included with NMS, a substantial number of optimal negative queries remain within the final subset, often outnumbering the positive ones by several times.
-As a result, applying NMS leads to an inaccurate reliability assessment.
-However, many schemes, including NMS, achieve significant improvement when used with the proposed contrastive framework.
-This is because even if the post-processing scheme is inaccurate, the framework can robustly perform by leveraging the negative queries.
-Lastly, we plot the OCE and image-level UQ performance with different hyperparameter selections to show the sensitivity of each scheme to the choice of hyperparameter.
-As depicted in Figure 10, the empirical results highlight that carefully identifying a reliable subset is crucial for achieving both high object-level calibration quality and effective image-level uncertainty quantification performance.
-We also highlight this claim with an exemplary visualization in Figure 9.
-Table 5: Comparison of the effectiveness of the proposed OCE and baseline metrics, focusing on their ranking correlation with the rankings given by the models’ calibration quality on optimal positive predictions (denoted by the superscript *).
-For each metric, various post-processing schemes (i.e., ways to identify positive predictions) are applied and indicated within parentheses.
-Correlations exceeding $0.9$ are highlighted in blue, while negative correlations are highlighted in red.
-Our OCE-based rankings consistently provide strong correlations with optimal ranking and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.
+Our OCE consistently provides strong correlations with optimal calibration errors and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.
 Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD) Average
@@
 $\mathsf{OCE_{MAX}}$ ($\mathsf{OCE_{MAX}}$) 0.954 0.929 0.998 0.954 0.958 0.983 0.955 0.697 0.976 0.934 $\pm$ 0.086
+To showcase OCE’s effectiveness in assessing calibration quality, we analyze how different metrics rank calibration errors across four DETR models.
+As baselines, we measure ECEs using the top $100$ detections  a standard confidence threshold of 0.3, and the threshold that minimizes the different metrics on the validation set.
+In a similar manner, we assess our OCEs using thresholds that minimize each OCE metric on the validation set.
+For a fair comparison, we execute the optimal matching process on each test dataset to obtain the optimal positive set.
+We then evaluate D-ECE, LA-ECE0, and OCE using these sets as reference scores and calculate the PCC correlation to the other scoring schemes across different datasets.
+Table 1 presents the correlations between the three calibration metrics on the optimal positive set and various methods.
+The first three rows demonstrate that the calibration metrics—D-ECE, LA-ECE, and our proposed OCE—are highly correlated.
+This indicates that all three metrics effectively capture the notion of calibration quality.
+Additionally, these metrics show a decent correlation with AP metrics.
+It is important to note that AP primarily accounts for accuracy rather than calibration quality, which may explain why the correlation, particularly for AP@50 on the out-of-distribution dataset, is not perfectly aligned.
+LRP, D-ECE, and LA-ECE0 exhibit inverse correlations with calibration quality when used alone.
+Furthermore, LA-ECE0 with a non-optimal positive set also shows an inverse correlation with the overall calibration errors estimated on the optimal positive set.
+This supports our assertion that these metrics are inadequate for measuring models’ calibration quality in conjunction with the employed post-processing scheme, unlike our OCE.
+Another important attribute of OCE is that it can not only compare different models but also identify the most reliable configuration of the same model.
+As illustrated in Figure 5, the confidence threshold vs.
+OCE curve exhibits a bell shape, unlike ECE metrics, and the optimal OCE is achieved around a threshold of $0.3$.
+This is particularly notable because it reflects the practical choices in many calibration studies to date [12, 20], affirming the reliability of using OCE.
+The D-ECE and OCE scores measured at the optimal LRP threshold—denoted by D-ECE (LRP) and OCE (LRP), respectively—show a strong correlation to those computed from the optimal positive set; thus, the method of using LRP to identify a positive set [20] appears acceptable.
+Experimentally, we confirmed that OCE and LRP provide thresholds in a similar range across different settings.
+Nonetheless, the biggest advantage of using OCE over other metrics is that OCE measures the model’s calibration quality alongside the post-processing scheme employed.
+While measuring the calibration error on the optimal positive set would ideally assess quality, this optimal set is not accessible during test-time inference.
+Consequently, the score measured on the optimal set may not correspond to the quality experienced in real-world applications, such as those employing confidence thresholding schemes.
+Using a fixed separation threshold of $0.3$ might serve as a reasonable approximation for distinguishing between optimal positives and negatives.
+However, trained or post-calibrated DETR models often have varying optimal confidence thresholds.
+For example, UP-DETR has an optimal threshold of approximately $0.5$, and low-temperature calibrated models are likely to require even higher thresholds.
+Consequently, employing a fixed threshold poses potential risks.
+Similarly, using a fixed number of top confident samples (e.g., Top-$100$) is inadequate, as it may include more predictions than the actual number of ground truth positives.
+VI Quantifying Image-level Reliability
+As demonstrated, positive and negative predictions exhibit varying levels of reliability.
+Interestingly, having predictions with low confidence scores does not necessarily imply low reliability.
+Our empirical observations show that confidence scores in negative predictions are actually inversely correlated with image-level reliability.
+More specifically, for a reliable instance, we observe that the confidence of positive predictions increases across the decoder layers, while that of negative predictions remains low; this results in a large gap between positive and negative predictions.
+In contrast, for unreliable instances (e.g., Figure 3a), the confidence score of the positive prediction does not increase across the layers, whereas the scores of negative predictions are either slightly elevated or remain unchanged.
+Consequently, there is a small gap between positive and negative predictions (e.g., Figure 3b).
+VI-A Proposed Method: Quantifying Reliability by Contrasting
+Based on the finding, we propose a post hoc UQ approach by contrasting the confidence scores of positives and negatives:
+$\mathsf{ContrastiveConf}(\bm{x})=\mathsf{Conf}^{+}(\bm{x})-\mathsf{Conf}^{-}(\bm{x})$ (11)
+$\mathsf{Conf}^{+}(\bm{x})=\frac{1}{|\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})|}\sum\nolimits_{(\hat{\bm{p}},\hat{\bm{b}})\in\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})}\max_{c}\hat{\bm{p}}[c]$ (12)
+$\mathsf{Conf}^{-}(\bm{x})=\frac{1}{|\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})|}\sum\nolimits_{(\hat{\bm{p}},\hat{\bm{b}})\in\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})}\max_{c}\hat{\bm{p}}[c]$ (13)
+where $\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})$ and $\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})$ are predicted sets of positive and negative predictions, respectively, and $\lambda$ is a scaling factor that can be determined on the validation set.
+We include an ablation study of the scaling factor in the subsequent section.
+Our results demonstrate that the proposed method is robust to the choice of scaling factor and consistently outperforms baseline approaches, achieving the robust performance with a scaling factor of $5.0$ $10.0$.
+Identifying a $\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})$ from $\hat{\mathcal{D}}_{\theta}(\bm{x})$ is a crucial factor in the success of this approach.
+In practice, however, neither the ground truth nor the optimal positives are available during the test time.
+Instead, we approximate the ground truth separation by applying a post-processing algorithm $\mathcal{A}^{*}$ that minimizes the calibration error on the validation set:
+$\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})=\mathcal{A}^{*}\circ\hat{\mathcal{D}}_{\theta}(\bm{x})$ (14)
+$\mathcal{A}^{*}=\text{argmin}_{\mathcal{A}}\mathsf{OCE}(\mathcal{A}\circ\hat{\mathcal{D}}_{\theta};\mathcal{I}_{val})$ (15)
+where $\mathcal{I}_{val}=\cup_{i\in\mathcal{V}}\{(i,j)\}_{j=1}^{N_{i}}$ is the set of all objects indices in the validation dataset $\mathcal{V}$.
 Table 2: Comparison of the proposed method ($\mathsf{ContrastiveConf}$) with baseline methods on their Pearson correlation coefficient with the image-level reliability in Equation (5), with different models.
-$\mathsf{Perf}$ is evaluated using AP (top), AP50 (middle), and AP75 (bottom) metrics.
-For each model, we apply the optimal matching, the standard post-processing scheme (Top-$100$ and confidence thresholding by $0.3$), and the proposed OCE-based post-processing scheme, as indicated within the parentheses.
+$\mathsf{Perf}$ is evaluated using AP.
+For each model, we apply the optimal matching, the standard post-processing scheme (Top-$100$ and confidence thresholding by $0.3$), or the proposed OCE-based post-processing scheme, as indicated within the parentheses.
 The strongest correlations are highlighted in bold, while negative correlations appear in red.
-First, the proposed method achieves the highest correlation with image-level reliability.
+First, the proposed contrasting method, $\mathsf{ContrastiveConf}$, achieves the highest correlation with image-level reliability.
 In addition, it is noteworthy that the average confidence of negative predictions ($\mathsf{Conf}^{-}$) exhibits a negative correlation with image-level reliability, in contrast to the positive correlation observed when using positive predictions ($\mathsf{Conf}^{+}$).
-Although the proposed fixed-threshold approach (0.3) occasionally outperforms the adaptive method, the performance gap remains marginal.
-Moreover, the fixed threshold consistently exhibits lower correlation with UP-DETR, for which the optimal threshold has been empirically determined to be approximately 0.5.
 Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
@@
 $\mathsf{Conf}^{-}$ (Optimal) -0.608 -0.584 -0.572 -0.586 -0.293 -0.296 -0.330 -0.311 -0.177 -0.235 -0.212 -0.196
-$\mathsf{ContrastiveConf}$ (Optimal) 0.700 0.684 0.709 0.695 0.601 0.633 0.662 0.634 0.612 0.657 0.672 0.643
+$\mathsf{ContrastiveConf}$ (Optimal) 0.700 0.648 0.684 0.662 0.601 0.659 0.656 0.644 0.612 0.660 0.667 0.647
 Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.619 -0.603 -0.581 -0.601 -0.409 -0.374 -0.372 -0.391 -0.262 -0.270 -0.239 -0.229
 $\mathsf{Conf}^{+}$ (0.3) 0.476 0.464 0.539 0.504 0.229 0.353 0.385 0.399 0.195 0.256 0.258 0.258
-Ours $\mathsf{Conf}^{+}$ (OCE) 0.442 0.477 0.547 0.512 0.237 0.355 0.393 0.370 0.228 0.291 0.270 0.274
-$\mathsf{Conf}^{-}$ (OCE) -0.633 -0.581 -0.571 -0.585 -0.365 -0.389 -0.379 -0.394 -0.245 -0.325 -0.253 -0.283
-$\mathsf{ContrastiveConf}$ (OCE) 0.585 0.570 0.597 0.587 0.334 0.432 0.431 0.438 0.287 0.362 0.302 0.327
-$\mathsf{ContrastiveConf}$ (0.3) 0.568 0.574 0.599 0.594 0.297 0.437 0.430 0.464 0.248 0.346 0.296 0.325
-$\mathsf{ContrastiveConf}$ (Top-100) -0.619 -0.253 -0.490 -0.321 -0.409 -0.342 -0.358 -0.360 -0.262 -0.175 -0.215 -0.139
-(a) AP
-Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
-UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
-Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.488 0.591 0.637 0.611 0.584 0.638 0.616 0.642 0.598 0.640 0.624 0.639
-$\mathsf{Conf}^{-}$ (Optimal) -0.583 -0.580 -0.555 -0.575 -0.248 -0.305 -0.316 -0.324 -0.137 -0.185 -0.175 -0.156
-$\mathsf{ContrastiveConf}$ (Optimal) 0.675 0.664 0.677 0.673 0.594 0.630 0.629 0.646 0.629 0.634 0.629 0.632
-Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.586 -0.581 -0.546 -0.571 -0.375 -0.369 -0.351 -0.394 -0.216 -0.217 -0.196 -0.204
-$\mathsf{Conf}^{+}$ (0.3) 0.445 0.435 0.494 0.465 0.229 0.321 0.342 0.379 0.203 0.223 0.179 0.194
-Ours $\mathsf{Conf}^{+}$ (OCE) 0.426 0.443 0.501 0.471 0.242 0.331 0.349 0.364 0.256 0.239 0.198 0.211
-$\mathsf{Conf}^{-}$ (OCE) -0.600 -0.576 -0.554 -0.573 -0.317 -0.394 -0.357 -0.406 -0.194 -0.277 -0.216 -0.247
-$\mathsf{ContrastiveConf}$ (OCE) 0.560 0.545 0.554 0.552 0.314 0.416 0.388 0.438 0.285 0.302 0.229 0.263
-$\mathsf{ContrastiveConf}$ (0.3) 0.532 0.553 0.557 0.561 0.277 0.418 0.388 0.456 0.237 0.298 0.216 0.259
-$\mathsf{ContrastiveConf}$ (Top-100) -0.586 -0.189 -0.435 -0.250 -0.375 -0.315 -0.332 -0.340 -0.216 -0.122 -0.171 -0.121
-(b) AP@50
-Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
-UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
-Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.490 0.593 0.629 0.598 0.486 0.586 0.608 0.563 0.488 0.607 0.635 0.574
-$\mathsf{Conf}^{-}$ (Optimal) -0.572 -0.541 -0.521 -0.538 -0.280 -0.244 -0.284 -0.274 -0.154 -0.234 -0.182 -0.189
-$\mathsf{ContrastiveConf}$ (Optimal) 0.672 0.649 0.662 0.649 0.541 0.562 0.614 0.563 0.534 0.627 0.641 0.587
-Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.580 -0.555 -0.523 -0.549 -0.373 -0.314 -0.325 -0.348 -0.226 -0.266 -0.206 -0.216
-$\mathsf{Conf}^{+}$ (0.3) 0.459 0.434 0.497 0.467 0.214 0.320 0.348 0.351 0.188 0.231 0.290 0.241
-Ours $\mathsf{Conf}^{+}$ (OCE) 0.423 0.447 0.503 0.478 0.234 0.330 0.361 0.322 0.206 0.272 0.293 0.262
-$\mathsf{Conf}^{-}$ (OCE) -0.597 -0.539 -0.520 -0.537 -0.339 -0.325 -0.327 -0.342 -0.220 -0.321 -0.217 -0.269
-$\mathsf{ContrastiveConf}$ (OCE) 0.557 0.532 0.548 0.545 0.319 0.386 0.390 0.381 0.258 0.346 0.314 0.312
-$\mathsf{ContrastiveConf}$ (0.3) 0.543 0.535 0.551 0.548 0.279 0.384 0.385 0.406 0.233 0.324 0.316 0.306
-$\mathsf{ContrastiveConf}$ (Top-100) -0.580 -0.225 -0.435 -0.291 -0.373 -0.295 -0.320 -0.334 -0.226 -0.170 -0.188 -0.137
-(c) AP@75
+Ours $\mathsf{Conf}^{+}$ (OCE) 0.477 0.478 0.547 0.512 0.252 0.330 0.393 0.370 0.191 0.276 0.270 0.274
+$\mathsf{Conf}^{-}$ (OCE) -0.640 -0.575 -0.571 -0.585 -0.349 -0.382 -0.379 -0.394 -0.256 -0.322 -0.253 -0.283
+$\mathsf{ContrastiveConf}$ (OCE) 0.656 0.588 0.628 0.613 0.359 0.407 0.441 0.440 0.275 0.350 0.317 0.327
+VI-B Numerical Analysis
+We compare different methods for quantifying per-image reliability based on different separation methods.
+For separation methods, we apply the optimal matching (as a reference) as well as the practical post-processing schemes such as fixed and adaptive confidence thresholding.
+We measure the Pearson correlation coefficient of each method with the $\mathsf{ImReli}$ computed based on AP metric (see Appendix for the detail).
+Table 2 shows that the proposed $\mathsf{ContrastiveConf}$ consistently achieves the best correlation with image-level reliability.
+Interestingly, the absolute value of $\mathsf{Conf}^{-}$ often surpasses that of $\mathsf{Conf}^{+}$ when a non-optimal separation scheme is applied.
+This observation highlights how our contrasting approach, which leverages the strengths of both $\mathsf{Conf}^{+}$ and $\mathsf{Conf}^{-}$, achieves robust performance across diverse settings.
+(a) COCO
+(b) Cityscapes
+(c) Foggy Cityscapes
+Figure 6: Impact of the scaling factor ($\lambda$) on image-level UQ performance of $\mathsf{ContrastiveConf}$ (OCE).
+Pearson correlation coefficient (PCC) using various scaling factors is reported.
+The optimal scaling factor lies within the range of $5.0$ to $10.0$, while this range generalizes well across out-of-distribution datasets.
+Furthermore, it shows the efficacy of $\mathsf{ContrastiveConf}$ over $\mathsf{Conf}^{+}$ (i.e., $\mathsf{ContrastiveConf}$ with $\lambda=0.0$).
+In addition, we conduct an ablation study on the scaling factor ($\lambda$), with results presented in Figure 6.
+The study reveals that the best performing scaling factor for COCO dataset lies between $5.0$ and $10.0$.
+Notably, this range remains effective even for out-of-distribution datasets such as Cityscapes and Foggy Cityscapes.
+Furthermore, the results demonstrate that $\mathsf{ContrastiveConf}$ with $1.0\leq\lambda\leq 10.0$ consistently outperforms $\mathsf{Conf}^{+}$ (i.e., $\lambda=0.0$, the leftmost point on each line).
+However, we also observe a rapid drop in performance when the scaling factor is excessively large, particularly on the OOD dataset, Foggy Cityscapes.
+Thus, caution is needed when applying to substantially different datasets.
 Table 3: Comparison of post-processing schemes based on calibration quality (i.e., OCE).
@@
 NMS 0.358 (0.90) 0.535 (0.90) 0.624 (0.90) 0.521 (0.90) 0.426 (0.90) 0.625 (0.90) 0.784 (0.90) 0.583 (0.90) 0.495 (0.90) 0.650 (0.90) 0.794 (0.90) 0.609 (0.90)
-Table 8: Comparison of post-processing schemes and baseline methods based on their Pearson correlation coefficients with image-level reliability when $\mathsf{Conf}^{+}$ and $\mathsf{ContrastiveConf}$ are applied, respectively.
+Table 4: Comparison of post-processing schemes and baseline methods based on their Pearson correlation coefficients with image-level reliability when $\mathsf{Conf}^{+}$ and $\mathsf{ContrastiveConf}(\lambda=1.0)$ are applied, respectively.
 We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding).
@@
 For Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies.
-For UP-DETR, the optimal thresholds exceed $0.5$, highlighting the potential limitation of using a fixed threshold in image-level uncertainty quantification (UQ) applications.
+For UP-DETR, the optimal thresholds often exceed $0.5$, highlighting the potential limitation of using a fixed threshold for image-level UQ.
 UQ Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
@@
 Top-$k$ 0.067 (1.00) 0.132 (1.00) 0.252 (1.00) 0.155 (1.00) -0.037 (2.00) 0.019 (1.00) 0.078 (1.00) 0.021 (1.00) 0.080 (1.00) -0.006 (1.00) 0.026 (1.00) 0.090 (1.00)
-NMS -0.588 (0.10) -0.560 (0.10) -0.548 (0.10) -0.562 (0.10) -0.402 (0.10) -0.346 (0.10) -0.332 (0.10) -0.365 (0.10) -0.240 (0.10) -0.265 (0.10) -0.218 (0.10) -0.224 (0.10)
-$\mathsf{ContrastiveConf}$ Thresholding 0.597 (0.55) 0.574 (0.30) 0.599 (0.30) 0.594 (0.30) 0.408 (0.80) 0.437 (0.30) 0.443 (0.20) 0.472 (0.35) 0.312 (0.75) 0.362 (0.25) 0.315 (0.40) 0.352 (0.35)
-Top-$k$ 0.629 (1.00) 0.506 (1.00) 0.474 (1.00) 0.527 (1.00) 0.409 (1.00) 0.377 (1.00) 0.324 (1.00) 0.411 (1.00) 0.284 (1.00) 0.254 (1.00) 0.156 (1.00) 0.298 (1.00)
-NMS -0.408 (0.10) 0.585 (0.10) 0.535 (0.10) 0.574 (0.10) -0.101 (0.10) 0.368 (0.10) 0.347 (0.10) 0.373 (0.10) -0.001 (0.10) 0.262 (0.10) 0.228 (0.10) 0.239 (0.10)
-(a) Predictions for a High-Reliability Prediction: Horse
-(b) Predictions for a High-Reliability Prediction: Chair
-(c) Predictions for a Low-Reliability Prediction: Chair
-Figure 5: The evolution of predictions made by Cal-DETR across decoder layers for a high-reliability and low-reliability image.
-The positive query (indexed by 0 and bordered with green) and the five negative queries with the largest IoU (indexed by 1-5 and bordered with red) are presented.
-(a) Predictions for a Low-Reliability Object: Laptop
-(b) Predictions for a High-Reliability Object: Dog
-Figure 6: The evolution of predictions made by Cal-DETR across decoder layers for two different objects (one reliable and one unreliable) in a low-reliability image.
-The positive query (indexed by 0 and bordered with green) and the five negative queries with the largest IoU (indexed by 1-5 and bordered with red) are presented.
-(a) UP-DETR (COCO)
-(b) D-DETR (COCO)
-(c) Cal-DETR (COCO)
-(d) DINO (COCO)
-(e) UP-DETR (Cityscapes)
-(f) D-DETR (Cityscapes)
-(g) Cal-DETR (Cityscapes)
-(h) DINO (Cityscapes)
-(i) UP-DETR (Foggy)
-(j) D-DETR (Foggy)
-(k) Cal-DETR (Foggy)
-(l) DINO (Foggy)
-Figure 7: Impact of confidence threshold selection on various performance metrics in UP-DETR, Deformable-DETR, Cal-DETR, and DINO on COCO, Cityscapes, and Foggy Cityscapes.
-(a) COCO
-(b) Cityscapes
-(c) Foggy Cityscapes
-Figure 8: Impact of the scaling factor ($\lambda$) on image-level UQ performance of $\mathsf{ContrastiveConf}$.
-Pearson correlation coefficient (PCC) using various scaling factors is reported.
-The optimal scaling factor lies within the range of $5.0$ to $10.0$, while this range generalizes well across out-of-distribution datasets.
-Furthermore, it shows the efficacy of $\mathsf{ContrastiveConf}$ over $\mathsf{Conf}^{+}$ (i.e., $\mathsf{ContrastiveConf}$ with $\lambda=0.0$).
+NMS -0.524 (0.10) -0.508 (0.10) -0.488 (0.10) -0.510 (0.10) -0.370 (0.10) -0.324 (0.10) -0.273 (0.10) -0.338 (0.10) -0.205 (0.10) -0.251 (0.10) -0.202 (0.15) -0.204 (0.10)
+$\mathsf{ContrastiveConf}$ Thresholding 0.597 (0.55) 0.513 (0.25) 0.563 (0.25) 0.540 (0.25) 0.408 (0.80) 0.389 (0.30) 0.422 (0.20) 0.437 (0.35) 0.312 (0.75) 0.317 (0.25) 0.288 (0.15) 0.311 (0.35)
+Top-$k$ 0.629 (1.00) 0.268 (1.00) 0.320 (1.00) 0.290 (1.00) 0.409 (1.00) 0.169 (1.00) 0.151 (1.00) 0.220 (1.00) 0.284 (1.00) 0.077 (1.00) 0.060 (1.00) 0.174 (1.00)
+NMS -0.227 (0.10) -0.157 (0.90) -0.402 (0.10) -0.140 (0.90) -0.044 (0.80) -0.034 (0.75) -0.203 (0.10) -0.080 (0.80) 0.035 (0.75) -0.035 (0.75) -0.149 (0.90) -0.075 (0.70)
+(a) Thresholding
+(b) Top-$k$
+(c) NMS
+Figure 7: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{Conf}^{+}$ and image-level reliability in Cal-DETR on COCO for different post-processing schemes.
+VII Comparative Study on Post-Processing
+In this section, we provide a comprehensive analysis regarding the impact of post-processing on the model’s overall calibration quality and UQ performance.
+For the post-processing (i.e., separation) methods, we compare the following approaches: (1) applying a threshold on the confidence score, (2) selecting the Top-$k$ predictions, and (3) utilizing NMS.
+Primarily, we evaluate OCE by varying the hyperparameter for each method (i.e., confidence threshold, $k$, and IoU threshold, respectively) and compare the resulting performance.
+In this first analysis, to exclude the dependency of the final performance on OCE, we choose the best hyperparameter on the validation set for each setting and compare their best possible performances.
+Table 3 shows that the confidence thresholding approach outperforms the top-$k$ and NMS approaches.
+The top-$k$ approach underperforms because the number of objects in an image varies significantly, making it impractical to determine a single optimal value for $k$.
+Top-$k$ becomes optimal when $k=20$.
+Given that the average and 95th percentile number of objects per image in the COCO dataset are 7 and 22, respectively, these results appear reasonable.
+Therefore, we confirm that using an excessively large number (e.g., 100) for top-$k$ is inadequate for achieving well-calibrated predictions.
+On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive queries are included, a substantial number of negative queries are retained, often outnumbering the positive ones by several times.
+Similarly, we compare the image-level uncertainty quantification performance, and the results are shown in Table 4.
+In this experiment, we fixed the scaling factor $\lambda$ at $1.0$ to minimize its influence on selecting the optimal hyperparameter.
+We reconfirm the effectiveness of confidence thresholding; for Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies.
+For UP-DETR, the optimal thresholds exceed $0.5$, highlighting the potential limitation of using a fixed threshold in image-level UQ applications.
+Moreover, the correlation is often negative when NMS is applied along with the $\mathsf{Conf}^{+}$ framework.
+This is because, while most of the optimal positive queries are likely to be included with NMS, a substantial number of optimal negative queries remain within the final subset, often outnumbering the positive ones by several times.
+As a result, applying NMS leads to an inaccurate reliability assessment.
+However, many schemes, including NMS, achieve significant improvement when used with the proposed contrastive framework; even if the post-processing scheme is inaccurate, the framework can robustly perform by leveraging the negative predictions.
+Lastly, we plot the OCE and image-level UQ performance with different hyperparameter selections to show the sensitivity of each scheme to the choice of hyperparameter.
+Figure 7 highlight that carefully identifying a reliable subset is crucial for achieving both high object-level calibration quality and effective image-level uncertainty quantification performance.
+We also present this with exemplary visualizations in Figure 8 and in the Appendix.
 (a) Thresholding (0.1)
@@
 Therefore, accurately identifying a reliable set of predictions significantly affects the reliability of DETR for downstream applications.
+VIII Conclusion
+The main contribution of our work is an in-depth analysis of the reliability of DETR frameworks.
+In particular, we reveal that DETR’s predictions for a given image exhibit varying calibration quality, highlighting the importance of identifying well-calibrated positive predictions.
+To address this challenge, we introduced a systematic framework that leverages our object-level calibration error metric to discern these positive predictions effectively.
+Furthermore, we proposed a novel uncertainty quantification method for estimating image-level reliability in DETR and conducted comparative studies of various post-processing schemes regarding their impact on DETR’s reliability.
+We hope our efforts expand the scope of DETR applications by enabling more precise and reliable deployment, sparking further research into the area of positive prediction identification.
+Acknowledgments
+The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper.
+This work was supported in part by the MIT Summer Research Program (MSRP), the MIT-Google Program for Computing Innovation, the MIT-Amazon Science Hub, the MIT-IBM Watson AI Lab, and MathWorks.
+Appendix A Dataset
+For our experiments, we used the Cityscapes and Foggy Cityscapes datasets, which each have 500 images of first-person driving footage in realistic environments.
+Foggy Cityscapes has the same base images as Cityscapes, but with fog simulated and added to create a further out of distribution set.
+Since the DETR models were trained on COCO, the Cityscapes and Foggy Cityscapes annotations were converted to correspond to the labels of COCO.
+More specifically, the person, bicycle, car, motorcycle, bus, train, and truck classes were transferred directly.
+In addition, the rider class of Cityscapes and Foggy Cityscapes was mapped to the person class of COCO.
+The other classes present in Cityscapes and Foggy Cityscapes are largely focused on image segmentation, and thus were omitted (e.g.
+building, sky, sidewalk).
+The pre-trained model weights were obtained from their respective official implementations.
+Appendix B Performance Metrics
+Average precision (AP) [17] is the most common accuracy metric for object detectors.
+Before calculating AP, the top-$k$ (e.g., typically $100$) class predictions for a given image are kept and sorted, while the others are discarded.
+Then, in order of this sorting, the predictions are compared against against the ground-truth bounding boxes, and the predicted label is compared with the ground truth label if the IoU is large enough.
+The precision/recall curve is calculated over the sorted predictions, and AP averages the precision across a set of 101 evenly spaced recall thresholds:
+$\text{AP}=\frac{1}{101}\sum_{r\in[0:0.01:1]}p(r)$ (16)
+where $p(r)$ is calculated by the maximum precision at recall level $r$.
+[…]
+$\text{LRP}_{\tau}=\frac{1}{N_{\text{FP}}+N_{\text{FN}}+N_{\text{TP}}}\left(N_{\text{FP}}+N_{\text{FN}}+\sum_{i=1}^{N_{\text{TP}}}\frac{1-\text{IoU}_{i}}{1-\tau}\right)$ (17)
+where $\tau$ is IoU threshold, $i$ denotes the index of each true positive prediction, and IoUi represents the IoU between that prediction and the best-matching ground-truth object.
+LRP is then computed by averaging LRP0.5 and LRP0.75.
+Appendix C Omitted Experimental Results
+This section provides the omitted experimental results.
+Figure 10 extends Figure 5 in Section V-A across different DETR variants and datasets.
+Figure 11 extends Figure 8 in Section VII across different DETR variants and post-processing schemes.
+Appendix D Contrasting with Confounding Negatives for Image-level UQ
+In Table 2, we observe an interesting empirical result: $\mathsf{Conf}^{+}$ using the top 100 predictions (i.e., the fourth row) exhibits a stronger negative correlation compared to $\mathsf{Conf}^{-}$ using the optimal negative predictions (i.e., the second row).
+While a stronger negative correlation might not be desirable when used alone for UQ, this phenomenon can be leveraged within our contrasting framework, similar to how negative predictions are utilized.
+For instance, rather than contrasting the average confidence scores of positive predictions against those of negative (non-positive) predictions, we can instead contrast them against the top 100 predictions.
+However, since positive predictions are included within this top 100 set, an alternative and potentially clearer approach is to contrast the positive predictions specifically against the top 100 non-positive predictions.
+Nevertheless, employing a fixed value of $k=100$ might introduce unintended consequences, particularly depending on the number of ground truth objects present in an image.
+We hypothesize that among non-positive predictions, those with bounding boxes significantly overlapping the positive predictions play a more meaningful role.
+Consequently, negative predictions that have random bounding boxes and very low confidence scores (as illustrated in Figure 1a) can be safely disregarded during the contrastive evaluation.
+To formalize this intuition, we introduce the concept of confounding negatives:
+Definition 3.
+Confounding negative predictions are defined as non-positive predictions whose bounding boxes significantly overlap with positive predictions, specifically exceeding a threshold overlap $\delta$:
+$\hat{\mathcal{D}}_{\theta}^{-\delta}(\bm{x})=\Big{\{}d^{-}\in\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})\mid\max_{d^{+}\in\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})}\mathrm{IoU}(d^{-},d^{+})\geq\delta\Big{\}}.$ (18)
+We conduct an ablation study comparing the original approaches, which use the entire set of negative predictions, against variants that employ different subsets of negatives described above.
+Our findings are summarized in Table 5.
+Overall, we see that restricting attention to negatives with significant overlap can sharpen the negative correlation for $\mathsf{Conf}^{-}$, especially on in-distribution data (COCO), supporting our hypothesis that misleading or “confouding” boxes are more informative for contrastive scoring.
+On the other hand, using entire negatives sometimes proves more robust in near-OOD and OOD scenarios, particularly for $\mathsf{ContrastiveConf}$.
+Table 5: Comparison of correlation results under $\mathsf{Conf}^{-}$ (rows 1–5) and $\mathsf{ContrastiveConf}$ (rows 6–10) when using different negative sets on three datasets: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD).
+“Entire Negatives” refers to all non-positive boxes, while “Top-100 Negatives” and “Top-100 Positives” limit the selection to the 100 highest-confidence predictions in each category.
+Confounding negatives restricts the negative set to boxes overlapping positive boxes by at least $\delta$.
+Stronger (more negative) correlation values under $\mathsf{Conf}^{-}$ appear when negatives more closely resemble the positive boxes, and $\mathsf{ContrastiveConf}$ tends to benefit from using such confounding negatives, especially in in-distribution scenarios.
+Methods Negatives COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
+UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
+$\mathsf{Conf}^{-}$ Entire Negatives -0.640 -0.575 -0.571 -0.585 -0.349 -0.382 -0.379 -0.394 -0.256 -0.322 -0.253 -0.283
+Top-100 Positives -0.619 -0.603 -0.581 -0.601 -0.409 -0.374 -0.372 -0.391 -0.262 -0.270 -0.239 -0.229
+Top-100 Negatives -0.640 -0.607 -0.607 -0.620 -0.349 -0.385 -0.415 -0.404 -0.256 -0.315 -0.272 -0.279
+Confounding Negatives ($\delta=0.5$) -0.641 -0.640 -0.574 -0.647 -0.330 -0.351 -0.364 -0.386 -0.250 -0.260 -0.254 -0.261
+Confounding Negatives ($\delta=0.75$) -0.635 -0.628 -0.561 -0.602 -0.385 -0.388 -0.384 -0.390 -0.285 -0.279 -0.269 -0.254
+$\mathsf{ContrastiveConf}$ Entire Negatives 0.656 0.588 0.628 0.613 0.359 0.407 0.441 0.440 0.275 0.350 0.317 0.327
+Top-100 Positives 0.632 0.610 0.604 0.619 0.412 0.379 0.388 0.404 0.270 0.277 0.257 0.247
+Top-100 Negatives 0.656 0.612 0.629 0.633 0.359 0.401 0.439 0.431 0.275 0.332 0.299 0.308
+Confounding Negatives ($\delta=0.5$) 0.653 0.641 0.623 0.654 0.340 0.371 0.434 0.418 0.268 0.284 0.314 0.296
+Confounding Negatives ($\delta=0.75$) 0.647 0.630 0.613 0.619 0.388 0.406 0.448 0.418 0.298 0.301 0.326 0.286
+(a) Object-level Calibration Error
+(b) Object-level Calibration Error
+(Foggy Cityscapes)
+(c) Corr. to Image-level Reliability
+(d) Corr. to Image-level Reliability
+(Foggy Cityscapes)
+Figure 9: A visualization of the difference in calibration between positive and negative predictions on the Cityscapes and Foggy Cityscapes datasets.
+In Figures 9a and 9b, Object-Level Calibration Error (OCE) values of their confidence scores are shown, where a lower score represents better-calibrated predictions.
+In Figures 9c and 9d, Pearson Correlation Coefficient (PCC) values between the ground truth $\mathsf{ImReli}$ and the average confidence scores are shown, where a higher value is better.
+(a) UP-DETR (COCO)
+(b) D-DETR (COCO)
+(c) Cal-DETR (COCO)
+(d) DINO (COCO)
+(e) UP-DETR (City)
+(f) D-DETR (City)
+(g) Cal-DETR (City)
+(h) DINO (City)
+(i) UP-DETR (Foggy)
+(j) D-DETR (Foggy)
+(k) Cal-DETR (Foggy)
+(l) DINO (Foggy)
+Figure 10: Impact of confidence threshold selection on various performance metrics in UP-DETR, Deformable-DETR, Cal-DETR, and DINO on COCO, Cityscapes, and Foggy Cityscapes.
+Higher scores are preferable ($\uparrow$), while lower scores are preferable ($\downarrow$).
 (a) Thresholding (UP-DETR)
@@
 (l) NMS (DINO)
-Figure 10: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{ContrastiveConf}$ and image-level reliability in Cal-DETR on COCO for different post-processing schemes.
+Figure 11: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{Conf}^{+}$ and image-level reliability across different DETR models on COCO for different post-processing schemes.

### v3

@@
 Abstract
-DEtection TRansformer (DETR) has emerged as a promising architecture for object detection, offering an end-to-end prediction pipeline.
-In practice, however, DETR generates hundreds of predictions that far outnumber the actual number of objects present in an image.
-This raises the question: can we trust and use all of these predictions?
-Addressing this concern, we present empirical evidence highlighting how different predictions within the same image play distinct roles, resulting in varying reliability levels across those predictions.
-More specifically, while multiple predictions are often made for a single object, our findings show that most often one such prediction is well-calibrated, and the others are poorly calibrated.
-Based on these insights, we demonstrate that identifying a reliable subset of DETR’s predictions is crucial for accurately assessing the reliability of the model at both object and image levels.
-Building on this viewpoint, we first address the shortcomings of widely used performance and calibration metrics, such as average precision and various forms of expected calibration error.
-Specifically, they are inadequate for determining which subset of DETR’s predictions should be trusted and utilized.
-In response, we present Object-level Calibration Error (OCE), which assesses the calibration quality more effectively and is suitable for both ranking different models and identifying the most reliable predictions within a specific model.
-As a final contribution, we introduce a post hoc uncertainty quantification (UQ) framework that predicts the accuracy of the model on a per-image basis.
-By contrasting the average confidence scores of positive (i.e., likely to be matched) and negative predictions determined by OCE, our framework assesses the reliability of the DETR model for each test image.
-(a) Random boxes. Low Confidence.
-(b) Accurate boxes. Calibrated.
-(c) Accurate boxes. Uncalibrated.
-Figure 1: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with at least one (i.e., blue boxes) being well-calibrated.
-This figure illustrates how DETR can handle the remaining predictions (i.e., red and gray boxes).
-In principle, DETR may (a) generate low confidence with random bounding boxes, (b) assign equally calibrated confidence with accurate prediction for the bounding boxes, or (c) poorly calibrated confidence with accurate boxes.
-Our analysis indicates that DETR mostly follows the third scheme; resulting in varying levels of reliability across predictions.
+DEtection TRansformer (DETR) and its variants have emerged as promising architectures for object detection, offering an end-to-end prediction pipeline.
+In practice, however, DETRs generate hundreds of predictions that far outnumber the actual objects present in an image.
+This raises a critical question: which of these predictions could be trusted?
+This is particularly important for safety-critical applications, such as in autonomous vehicles.
+Addressing this concern, we provide empirical and theoretical evidence that predictions within the same image play distinct roles, resulting in varying reliability levels.
+Our analysis reveals that DETRs employ an optimal “specialist strategy”: one prediction per object is trained to be well-calibrated, while the remaining predictions are trained to suppress their foreground confidence to near zero, even when maintaining accurate localization.
+We show that this strategy emerges as the loss-minimizing solution to the Hungarian matching algorithm, fundamentally shaping DETRs’ outputs.
+While selecting the well-calibrated predictions is ideal, they are unidentifiable at inference time.
+This means that any post-processing algorithm—used to identify trustworthy predictions—poses a risk of outputting a set of predictions with mixed calibration levels.
+Therefore, practical deployment necessitates a joint evaluation of both the model’s calibration quality and the effectiveness of the post-processing algorithm.
+However, we demonstrate that existing metrics like average precision and expected calibration error are inadequate for this task.
+To address this issue, we further introduce Object-level Calibration Error (OCE), which evaluates calibration by aggregating predictions per ground-truth object rather than per prediction.
+This object-centric design penalizes both retaining suppressed predictions and missed ground truth foreground objects, making OCE suitable for both evaluating models and identifying reliable prediction subsets.
+Finally, we present a post hoc uncertainty quantification (UQ) framework that predicts per-image model accuracy.
 I Introduction
@@
 This paradigm shift has led to the exploration of various DETR variants, positioning them as potential foundation models for object detection tasks.
-While notable progress has been made, the reliability of these predictions remains insufficiently investigated.
-The primary objective of this study is to determine how DETR can be used in a trustworthy manner on new downstream images/tasks and to establish how their predictions should be properly used.
-For example, when building an auto-labeling system using DETRs, it is crucial to consider that the model may not always provide accurate predictions.
-As a result, when the model’s reliability for a specific image is in question, the image might need to be reviewed by a human labeler.
-Therefore, evaluating reliability at the image level is often necessary to assess the model’s overall understanding of the given test image.
-However, owing to DETR’s unique set-prediction mechanism, quantifying its image-level reliability is not straightforward and remains largely underexplored.
-Moreover, there is an ongoing debate within the community about whether DETR can be considered an entirely end-to-end solution, free from any post-processing requirements in practice.
+While notable progress has been made, the reliability of these predictions remains under-investigated.
+Quantifying the reliability of these models’ predictions is critical across different scenarios.
+For instance, when building an auto-labeling system using DETRs, it is crucial to consider that the model may not always provide accurate predictions.
+Images where the model’s reliability is questionable may need to be reviewed by human labelers.
+In safety-critical applications, such as autonomous vehicles, knowledge of the reliability of a prediction in the perception pipeline for a given scene is crucial for the downstream decision-making.
+Without properly understanding this reliability, the system may, for example, overlook a pedestrian, leading to catastrophically wrong decisions.
+However, owing to DETR’s unique set-prediction mechanism, quantifying its reliability is not straightforward and remains largely underexplored.
 In particular, DETR outputs a fixed number of predictions, typically in the hundreds; consequently, the central concern is which predictions can be trusted and used.
-Practitioners often employ heuristic approaches to determine this subset, such as by setting a user-defined threshold to retain only a small subset of high-confidence (e.g., $>0.7$) outputs, as seen in the official demo.
+Practitioners often employ heuristic approaches to select a subset of predictions, such as by setting a user-defined threshold to retain only a small number of high-confidence (e.g., $0.7$) outputs, as seen in the official demo [detr_demo].
 Similarly, previous studies exploring model reliability, such as [8, 9, 10, 11, 12, 13], have also applied a user-defined confidence threshold (e.g., 0.3) to retain a subset of predictions for evaluating calibration quality, rather than using the entire set.
-On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR [14] and 300 out of 900 for DINO [15]) based on confidence score.
-Nonetheless, how different subset selection schemes affect model reliability, as well as the process of choosing appropriate configurations for each scheme, remain under-explored, and their significance has yet to be fully recognized.
-Our primary findings reveal that DETR predictions within the same image are interdependent, leading to significantly varying levels of reliability.
-Since DETR is trained using gradients derived exclusively from optimally matched predictions (i.e., the predictions most closely corresponding to each annotated object), it will output at least one well-calibrated prediction for each perceived object while having flexibility in handling the remaining unmatched predictions.
-Throughout the paper, we refer to those matched predictions as positive predictions and the others as negative predictions.
-In principle, to avoid an inaccurate match, DETR could assign those unmatched predictions low confidence with random bounding boxes (Figure 1a).
-Alternatively, it could generate similarly accurate bounding boxes with (nearly) equally calibrated confidence (Figure 1b).
-Otherwise, it may assign poorly calibrated confidence scores (Figure 1c).
-Our analysis indicates that a prediction generated by DETR mostly follows the third scheme.
-Building on this observation, this paper investigates the importance of distinguishing between positive and negative predictions, and addresses the following three research questions: (RQ1) Do all predictions generated for a given image exhibit comparable levels of reliability?
-(RQ2) If not, what is the appropriate way to identify reliable predictions across the entire set?
-and (RQ3) How can we accurately assess DETR’s image-level reliability?
-In addressing these questions, our specific contributions include:
-  1. 1.
-    We provide both qualitative and quantitative insights into how positive and negative predictions influence the model’s reliability.
-Our analysis reveals that, unlike positives, negative predictions are often poorly calibrated and therefore should be handled separately.
-Furthermore, their confidence scores are inversely correlated with image-level reliability, highlighting the need for a separation method to ensure DETR’s reliable use (Section IV).
-  2. 2.
-    Identifying ground-truth positives and negatives ideally requires a matching process based on ground-truth annotations, which are unavailable at test time.
-Consequently, an alternative framework for identifying positive predictions is needed.
-However, we show that existing performance and calibration metrics—such as average precision [16, 17, 18] and expected calibration error [8, 19, 20]—are inadequate for this purpose.
-To address this, we introduce a new calibration metric, object-level calibration error (OCE), which assesses calibration error along ground-truth objects rather than predictions.
-We demonstrate that OCE is more suitable for both ranking different models by their calibration qualities and identifying the positive predictions (Section V).
-  3. 3.
-    Based on our findings, we introduce a novel framework for quantifying image-level reliability by contrasting the average confidence scores between the positive and negative predictions identified by our OCE metric.
-We conduct numerical experiments, demonstrating the effectiveness of our approach across in-distribution, near out-of-distribution, and far out-of-distribution scenarios (Section VI).
-In addition to addressing the three primary research questions, we present a comparative analysis of standard post-processing methods for identifying positive predictions in DETR from an uncertainty quantification (UQ) perspective (Section VII).
-Our results indicate that thresholding on the predictive confidence score is more effective than widely used techniques such as top-$k$ selection or non-maximum suppression (NMS) for obtaining a reliable subset of predictions.
+On the other hand, the published implementations of several DETR variants select the top-$k$ outputs (e.g., 100 out of 300 for Deformable-DETR [14] and 300 out of 900 for DINO [15]) based on confidence scores.
+Nonetheless, how different subset-selection schemes affect model reliability, as well as the process of choosing appropriate configurations for each scheme, remains under-explored, and their significance has yet to be fully understood.
+To address this gap and move beyond heuristic subset selection, we first seek a fundamental characterization of DETRs’ predictions based on the training objective used.
+Our theoretical analysis shows that the Hungarian loss incentivizes a “specialist strategy”: having a single well-calibrated prediction (which we refer to as the primary prediction) to represent each ground-truth object, while suppressing all remaining predictions (referred to as the secondary predictions).
+This specialist strategy presents a fundamental challenge for practical deployment: while those primary predictions are trained to be reliable, they are unidentifiable at test time.
+This means that any post-processing algorithm poses a risk of outputting a set of predictions with mixed calibration levels.
+Consequently, evaluating model reliability requires jointly assessing the model’s calibration quality and the effectiveness of the post-processing algorithm in recovering the well-calibrated subset.
+To facilitate this joint evaluation, we introduce object-level calibration error (OCE), which evaluates calibration along ground-truth objects rather than predictions (Section V).
+OCE penalizes both missing ground truth objects and retaining unreliable secondary predictions, enabling proper assessment of model + post-processing combinations, something that common existing metrics are fundamentally incapable of.
+For instance, average precision (AP) [16, 17, 18] favors retaining all predictions, while expected calibration error (ECE) [8, 19, 20] favors discarding nearly all predictions to achieve artificially low error by ignoring missed objects.
+Finally, we leverage both our theoretical characterization of DETRs’ specialist strategy as well as OCE’s ability to identify primary and secondary predictions and develop a method for image-level uncertainty quantification.
+More specifically, we first demonstrate that the foreground confidence contrast between primary and secondary predictions is strongly correlated with image-level reliability.
+This motivates our novel framework for quantifying image-level reliability by measuring the confidence contrast between the “positive predictions” retained by an OCE post-processing and the remaining “negative” ones.
+We conduct numerical experiments across in-distribution, near out-of-distribution, and far out-of-distribution scenarios, and show the effectiveness of our approach (Section VI).
+In short, our key contributions are:
+  • We theoretically characterize and empirically validate the optimal strategy that DETR takes to minimize the Hungarian loss: it creates a single well-calibrated prediction per object while suppressing all others to near-zero foreground confidence while maintaining reasonable bounding box accuracy.
+  • We introduce OCE, a metric suitable for both evaluating model calibration and identifying reliable post-processing by measuring calibration per ground-truth object, avoiding pitfalls of existing metrics.
+  • We propose an algorithm to quantify image-level reliability by measuring the confidence contrast between positive and negative predictions identified by OCE.
+We demonstrate the effectiveness of our approach with numerical experiments on the COCO and Cityscapes datasets.
 II Related Work
@@
 In other words, low-confidence samples should exhibit low accuracy, and high-confidence samples should demonstrate high accuracy.
-For instance, if a model predicts an event with 80% confidence, it should be correct around 80% of the time for that event to be considered well-calibrated.
+For example, if among all instances to which the model assigns an 80% probability, the event occurs in about 80% of those instances, the model is well‑calibrated.
 Achieving good calibration is crucial because it improves trust in the model’s predictions, guides more informed decision-making, and enables better risk assessment in real-world applications.
-To evaluate the alignment, they measure the expected calibration error by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin.
+To evaluate the alignment, one can measure the expected calibration error by binning predictions based on their confidence scores and computing the mean absolute error between the average confidence and the corresponding accuracy within each bin.
 However, in object detection tasks, measuring accuracy is not straightforward because predictions comprise both class probabilities and bounding boxes.
-Additionally, due to the set prediction nature of object detection models, it is unclear which ground truth object corresponds to each prediction.
-To address these challenges, D-ECE [8] defines precision as the accuracy metric and matches each prediction to a ground truth object based on an intersection over union (IoU) threshold (commonly set at $0.5$ or $0.75$).
-On the other hand, LaECE [21] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
-However, as we demonstrate, DETR exhibits poor calibration on negative predictions, regardless of its actual reliability.
-In practice, users often focus only on positive samples.
-Yet, when model calibration is evaluated across the entire set of predictions, including the negatives, the resulting evaluation becomes highly biased and therefore unreliable.
+Additionally, due to the set-prediction nature of object detection models, it is unclear which ground-truth object corresponds to each prediction.
+To address these challenges, detection expected calibration error (D-ECE) [8] defines precision as the accuracy metric and matches each prediction to a ground-truth object based on an intersection over union (IoU) threshold (commonly set at $0.5$ or $0.75$).
+On the other hand, localisation-aware expected calibration error (LaECE) [19] defines accuracy as the product of precision and IoU, thereby accounting for localization errors as well.
+Yet, when model calibration is evaluated across the entire set of predictions, which includes the secondary predictions, the resulting evaluation becomes highly biased and therefore unreliable.
 To avoid this issue, previous studies have often measured D-ECE only on predictions with confidence scores exceeding $0.3$.
-As detailed in later sections, a threshold of $0.3$ approximates the optimal positive predictions in some DETR variants; thus, using such a fixed threshold may occasionally be acceptable.
-However, the optimal threshold is not guaranteed to be always $0.3$ for other models like UP-DETR, which introduces potential risks.
+As detailed in later sections, applying a threshold of $0.3$ selects the well-calibrated predictions in some DETR variants; thus, using such a fixed threshold may occasionally be acceptable.
+However, the optimal threshold is not guaranteed to be always $0.3$ for other models, like UP-DETR, which introduces potential risks.
 In response, our paper proposes OCE, which measures the model’s calibration quality alongside the employed post-processing scheme and thus can adaptively identify the reliable subset.
@@
 Several studies focus on out-of-distribution (OOD) identification in object detection models.
-For example, [22] propose a built-in OOD detector to isolate OOD data for human review, including those of unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty), by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions.
-[23] generate outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data.
-Similarly, [19] employ an auxiliary detection model capable of expressing its confidence.
+For example, [22] proposes a built-in OOD detector to isolate OOD data for human review, including those of unknown and uncertain classes (i.e., epistemic but not aleatoric uncertainty), by modeling the distribution of training data and assessing whether samples belong to any of the training class distributions.
+[23] generates outlier data from class-conditional distribution estimations derived from in-distribution data, training the model to assign high OOD scores to this generated data and low OOD scores to the original in-distribution data.
+Similarly, [19] employs an auxiliary detection model capable of expressing its confidence.
 Other works, including [24] and [25], investigate the latent representations generated by object detection models to identify the OOD nature of the input.
@@
 Moreover, they predominantly address CNN-based models and explore the methodological way to better quantify the uncertainty in object detection models.
-In contrast, our paper emphasizes the significance of identifying reliable sets within the entire set of predictions for uncertainty quantification, particularly in DETRs.
-Another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
+In contrast, our paper emphasizes the significance of identifying a reliable subset within the entire set of predictions for uncertainty quantification, particularly in DETRs.
+Further, another novelty of our work lies in investigating an appropriate methodology to integrate different predictions’ confidence estimates to quantify image-level reliability.
 III Preliminaries
@@
 This representation is then passed to the Transformer encoder, which extracts feature vectors.
-These feature vectors are sent to the decoder, which receives $M$ learned object queries together.
+These feature vectors are fed to the decoder, which receives $M$ learned object queries together.
 The decoder outputs $M$ prediction sets, each containing a bounding box and corresponding class probabilities.
 III-A Detection Transformer (DETR)
-We consider a test image $\bm{x}$ and denote $\mathcal{D}_{\bm{x}}$ as the set of ground truth objects present in the image.
-Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{D}}_{\theta}(\bm{x})$.
-Each prediction $d\in\hat{\mathcal{D}}_{\theta}(\bm{x})$ is characterized by a bounding box $\hat{\bm{b}}$ and an associated class label with a corresponding probability $\hat{p}$.
+Consider a test image $x$ and denote the set of ground-truth objects present in the image by $\mathcal{Y}$.
+Analogously, the set of predictions generated by DETR, parameterized by $\theta$, is denoted by $\hat{\mathcal{Y}}_{\theta}(x)$.
+Each prediction $\hat{y}_{i}\in\hat{\mathcal{Y}}_{\theta}(x)$ is characterized by a bounding box $\hat{b}_{i}$ and an associated class probability distribution $\hat{p}_{i}$ over $\mathcal{C}\cup\{\emptyset\}$, where $\mathcal{C}$ is the set of foreground classes and $\emptyset$ represents the background class.
 The structure of DETR is composed of two main components: the Transformer encoder, which extracts a collection of features from the given image; and the Transformer decoder, which uses these features to make predictions.
@@
 Together, these predictions form the final outputs, making DETR’s predictions essentially an $M$-element set.
-We refer to Figure 2 for an illustration.
-It is noteworthy that the encoder follows the common structure of standard computer vision models, whose reliability has been relatively widely explored [26, 27, 28, 29].
-This foundation further enables the use of prominent post hoc UQ techniques, such as Monte Carlo dropout [30].
-However, despite the decoder being the predominant component for object detection, there is a gap in understanding and quantifying its reliability due to its unique structural characteristic: set prediction.
-Therefore, this paper delves into the underlying characteristics of these predictions and presents a methodology to quantify the reliability of DETR.
-III-B Bipartite Matching
-Since the number of queries in DETR, $M$, is much higher than the number of annotated objects, DETR matches each object with the corresponding best model prediction during its training.
-To compute this optimal matching for the predictions in a given image, a bipartite matching algorithm is applied.
-More specifically, a matching cost between each pair of a given prediction and an object is defined as follows:
-$\mathcal{L}_{matching}=\mathcal{L}_{class}+\mathcal{L}_{box}$ (1)
-where $\mathcal{L}_{class}$ is the negative prediction confidence of the ground truth class and $\mathcal{L}_{box}$ is the linear combination of the $\ell_{1}$ loss between the corners of the bounding boxes and $\mathcal{L}_{iou}$.
-$\mathcal{L}_{iou}$ is the Generalized Intersection over Union (GIoU) [31] loss between bounding boxes.
-After computing this matching cost for every combination of prediction set and ground truth objects, DETR then efficiently calculates the permutation that minimizes the total matching cost using the Hungarian matching algorithm [32, 7].
-III-C Expected Calibration Error (ECE) Metrics
-D-ECE.
-For a given object detection model, detection expected calibration error (D-ECE) [8] quantifies how closely the model’s predicted confidences align with its observed precision.
-Specifically, let $\hat{\mathcal{D}}$ be the set of all detections that the model produces on the validation set.
-Each detection has an associated predicted confidence score $\hat{p}\in[0,1]$.
-To compute D-ECE, the confidence space $[0,1]$ is partitioned into $J$ bins, for instance via uniform intervals $\big{[}0,\tfrac{1}{J}\big{)},\big{[}\tfrac{1}{J},\tfrac{2}{J}\big{)},\dots,\big{[}\tfrac{J-1}{J},1\big{]}$.
-Let $\widehat{\mathcal{D}}_{j}\subseteq\widehat{\mathcal{D}}$ be the subset of detections falling into the $j$-th bin.
-Denote by $\bar{p}_{j}$ the average confidence of detections in bin $j$, and by $\mathrm{precision}(j)$ the empirical precision in bin $j$.
-A detection is considered a true positive if it has the correct predicted class label and its IoU with the ground-truth box exceeds a given threshold $\tau$.
-The empirical precision $\mathrm{precision}(j)$ is thus the fraction of detections in $\widehat{\mathcal{D}}_{j}$ that are true positives.
-D-ECE then aggregates the absolute difference between each bin’s average confidence and observed precision, weighted by bin size, and is defined as:
+See Figure 1 for an illustration.
+III-B Hungarian Loss in DETR
+Let $\hat{\mathcal{Y}}_{\theta}(x)=\{\hat{y}_{1},...,\hat{y}_{M}\}$, where $\hat{y}_{i}=(\hat{p}_{i},\hat{b}_{i})$.
+This set is compared against $\mathcal{Y}=\{y_{1},...,y_{N}\}$, where $N$ is the number of ground-truth objects.
+To evaluate the loss, a one-to-one correspondence between the two sets is first established.
+The ground-truth set is padded with MNM-N background objects, and an optimal matching permutation $\hat{\sigma}$ is found by minimizing a pairwise matching cost $\mathcal{L}_{\text{match}}$ using the Hungarian algorithm [32, 7].
 […]
-Following the standard protocol [20], two IoU thresholds ($\tau=0.5$ and $\tau=0.75$) are used, and the D-ECE scores obtained under these thresholds are averaged to yield the final metric.
-Because this process is carried out on a validation set, it provides an overall measure of the model’s calibration performance on that set.
-A lower D-ECE indicates that the confidence scores are better aligned with actual precision, while a higher D-ECE reveals a potential mismatch between confidence and true detection performance.
-LaECE.
-To account for both localization error and classification precision, localisation-aware expected calibration error (LaECE) [19] aligns model confidence with the product of the predicted bounding box’s precision and its IoU with the ground-truth box.
-Specifically, LaECE is computed in a class-wise manner to mitigate class imbalance.
-Let $c$ index the class, and let $\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$.
-Partition the confidence scores into $J$ bins, and let $\widehat{\mathcal{D}}_{j}^{c}\subseteq\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$ in bin $j$.
-Define $\bar{p}^{c}_{j}$ to be the average confidence in $\widehat{\mathcal{D}}_{j}^{c}$, $\mathrm{precision}^{c}(j)$ to be the corresponding precision, and $\bar{\mathrm{IoU}}^{c}(j)$ to be the average IoU of $\widehat{\mathcal{D}}_{j}^{c}$.
-The LaECE metric is then given by:
-$\text{LaECE}\triangleq\frac{1}{K}\sum_{c=1}^{K}\sum_{j=1}^{J}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}_{j}^{c}-\text{precision}^{c}(j)\times\bar{\text{IoU}}^{c}(j)\Bigr{|}$ (3)
-where $K$ is the total number of classes, and an IoU threshold $\tau$ is employed to determine whether a detection is considered a true positive.
-LaECE0 [20] sets the IoU threshold $\tau$ to zero and assigns an average IoU of zero to false positive detections, thereby reducing the original formulation.
-In particular, if a detection is not a true positive under $\tau=0$, its IoU contribution is taken to be zero.
-The resulting metric becomes:
-$\text{LaECE}_{0}=\frac{1}{K}\sum^{K}_{c=1}\sum^{J}_{j=1}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}^{c}_{j}-\bar{\text{IoU}}^{c}(j)\Bigr{|}.$ (4)
-By setting $\tau=0$, false positives have $\bar{\mathrm{IoU}}=0$, and thus the calibration criterion focuses more explicitly on the contribution of bounding box localization to the overall calibration error.
-III-D Image-Level Reliability
-We introduce a formal definition of image-level reliability by examining the model’s overall object detection performance on the image, extending [33, 29].
-Definition 1.
-We define image-level reliability as a measure of how accurately and confidently the predictions match the ground truth objects:
-$\mathsf{ImReli}(\bm{x};\theta)\triangleq\mathsf{Perf}\big{(}\hat{\mathcal{D}}_{\theta}(\bm{x}),~{}\mathcal{D}_{\bm{x}}\big{)}$ (5)
-where one can use any standard performance metrics, such as average precision and recall, for $\mathsf{Perf}$ depending on the user’s requirements.
-By its definition, image-level reliability directly addresses the model’s applicability to a given test instance.
-However, since image-level reliability requires ground truth annotations for its determination, obtaining it during inference is not feasible.
-Therefore, this paper’s objective is to develop a method that assigns a quantitative score to each image instance, closely aligning with the model’s image-level reliability.
-(a) Predictions within High-Reliability Image
-(b) Predictions within Low-Reliability Image
+$\mathcal{L}(\hat{y}_{\hat{\sigma}(j)},y_{j})=\sum_{j=1}^{M}w_{j}\left[-\log\hat{p}_{\hat{\sigma}(j)}(c_{j})+\mathbbm{1}_{\{c_{j}\neq\emptyset\}}\mathcal{L}_{\text{box}}(b_{j},\hat{b}_{\hat{\sigma}(j)})\right],$ (1)
+where $\hat{p}_{i}(c)$ is the predicted probability for class $c$ of the $i$-th prediction.
+$\mathcal{L}_{\text{box}}$ is a linear combination of the IoU and L1 loss between the two bounding boxes.
+This process naturally partitions the $M$ predictions into two groups for each image: the optimal positive predictions, which are matched to ground-truth foreground objects, and optimal negative predictions, which are matched to background objects.
+In practice, the log-probability term for negative predictions (where $c_{j}=\emptyset$) is down-weighted (e.g., by a factor of $w_{\emptyset}=0.1$):
+$w_{j}=\begin{cases}w_{\emptyset}&\text{if }c_{j}=\emptyset\\ 1&\text{if }c_{j}\neq\emptyset.\end{cases}$
+This structure, which penalizes high-confidence foreground predictions from negative predictions, forms the basis of our subsequent analysis.
+(a) Accurate boxes. Calibrated.
+(b) Random boxes. Low Confidence.
+(c) Accurate boxes. Uncalibrated.
+Figure 2: DETR generates hundreds of predictions for each image, resulting in multiple predictions per object, with at least one (i.e., blue boxes) being well-calibrated.
+This figure illustrates how DETR handles the remaining predictions (i.e., red and gray boxes).
+DETR could (a) assign equally high confidence scores with accurate bounding boxes to all predictions; however, this strategy fails to optimize the Hungarian loss, as high-confidence unmatched predictions incur substantial penalties.
+Instead, DETR could assign low confidence to unmatched predictions, either with (b) random bounding boxes or (c) accurate bounding boxes.
+Our analysis reveals that DETR follows the third strategy, which explains the varying levels of reliability observed across predictions.
+IV Optimal Prediction Strategies in DETR
+In this work, we evaluate a model’s reliability using the concept of calibration.
+We define a prediction as perfectly calibrated if its predicted probability distribution matches the true posterior probability distribution of the object it is matched to.
+We formalize this by considering an object whose true, and potentially ambiguous, class identity is represented by a ground-truth probability distribution $p_{\text{gt}}$.
+A prediction is perfectly calibrated with respect to this object if its predicted distribution $\hat{p}$ exactly matches the true distribution $p_{\text{gt}}$, i.e.,
+$\hat{p}(c)=p_{\text{gt}}(c)\quad\forall c\in\mathcal{C}\cup\{\emptyset\}.$ (2)
+IV-A Motivation and Scope
+During the inference stage at test-time, ground-truth annotations are unavailable, meaning the optimal positive predictions remain unknown.
+This raises the question of Which of the predictions can be trusted and used?
+If all of the model’s predictions were generated independently and were well-calibrated (e.g., Figure 2a), the large number of predictions would not be a concern.
+We could simply apply well-known algorithms like NMS to remove duplicates and resolve redundancy.
+However, empirical evidence suggests that such independence is unlikely to happen, as it appears to conflict with the minimization of the Hungarian loss in Equation (1).
+In practice, we observe that DETR tends to assign a well-calibrated confidence score to only a single prediction per object, while suppressing remaining predictions with low foreground confidence.
+Intriguingly, although the Hungarian loss imposes no penalty on the localization of these background-matched optimal negative predictions—implying they could potentially output random bounding boxes (Figure 2b)—we find that DETR exhibits a strong preference for accurate localization (Figure 2c), consistently producing precise boxes even for optimal negative predictions.
+In the following sections, we investigate this behavior in depth.
 Figure 3: Visualizations of the predictions generated by Cal-DETR.
@@
 For each prediction and layer, the maximum confidence score and its corresponding label are visualized.
-When the model is confident, the confidence score of the positive prediction either increases or remains high across the decoder layers, while those of negative predictions decrease or stay low despite its accurate bounding box predictions.
-On the other hand, when the model is uncertain, DETR assigns a confidence score to positive prediction based on its confidence, thereby maintaining good calibration.
-However, conversely, it slightly increases or maintains the confidence scores for negative predictions.
-IV Impact of Negative Predictions on DETR’s Reliability
-IV-A Motivation and Scope
-Notably, the number of predictions generated by DETR, $|\widetilde{\mathcal{D}}_{\theta}(\bm{x})|$, is fixed and often in the hundreds, far exceeding the number of ground truth objects.
-To address this issue during model training, a bipartite matching algorithm is used to find the optimal matching prediction [7] for each ground truth object based on the alignment of the class label and bounding box (as detailed in Section III-B).
-The training loss depends on the quality of this matching, and as a result, the parameters $\theta$ are primarily optimized to enhance the accuracy of these matched predictions.
-In this paper, we refer to the matched predictions as optimal positive predictions, while the remaining predictions are termed optimal negative predictions.
-At the inference stage in real-world scenarios, however, ground truth annotations are unavailable, meaning the optimal positive predictions remain unknown.
-This raises the question of whether all predictions are trustworthy, or should only a specific subset be chosen—and if so, which subset?
-If all of the model’s predictions were generated independently and were well-calibrated (e.g., Figure 1b), the large number of predictions would not be a concern.
-We could simply apply well-known algorithms like NMS to remove duplicates and resolve redundancy.
-Notably, we have observed that DETR assigns well-calibrated confidence scores to only a single positive prediction per object.
-Meanwhile, unmatched negative predictions, which often have accurate bounding boxes, receive uncalibrated low scores (e.g., Figure 1c).
-As a result, the overall calibration error becomes significantly high if we include those noisy predictions in the final inference results.
-Hence, as detailed in the following sections, improper handling of negative predictions can severely compromise DETR’s reliability, with potentially catastrophic consequences in safety-critical scenarios.
-For this reason, properly distinguishing the positives from the negatives is a crucial task for ensuring the reliable use of DETR.
+The model’s attention mechanism enables it to learn a “specialist strategy.” When an object is confidently identified, a single primary query (the intended optimal positive prediction) is assigned a (high) calibrated confidence score.
+In parallel, all other remaining queries (the intended optimal negative predictions) have their confidence scores suppressed to near-zero, even while they maintain accurate bounding box predictions.
+This specialist strategy—where the model designates one query as the sole reliable predictor and the rest as background specialists—is demonstrably favored by the loss function over having multiple competing, moderately-confident predictions for the same object.
 IV-B Exploring the Anatomy of DETR’s Predictions
-We begin our analysis by visualizing and examining the outputs generated from the DETR decoders.
+We begin our qualitative analysis by visualizing and examining the outputs generated from the DETR decoders.
 Since the Transformer decoder outputs only representation vectors, investigating their evolution across layers is not straightforward.
@@
 This is feasible due to the alignment of intermediate representations, facilitated by residual connections between decoder layers [34].
-Sample visualizations are in Figure 3.
+Sample visualizations are in Figures 3 and 4.
 In the first decoder layer, the model appears to explore the encoded image features, producing varied queries that result in various plausible predictions.
-In this early stage, the distinction between positive and negative queries can be ambiguous.
-However, the self-attentions through the subsequent decoder layers progressively refine these predictions.
-By the final layer, the model selects a single query (i.e., optimal positive prediction) and assigns a confidence score based on its understanding of the image and the object.
-In contrast, the confidence scores for neighboring queries (i.e., optimal negative predictions) do not increase to the same extent as the positives and even decrease in high-reliability images.
-In contrast, in low-reliability images, the confidence score of the positive query does not significantly increase, while the scores of the negative ones are either slightly raised or unchanged.
-Based on this observation, we present our claim:
-Claim 1.
-Predictions from DETR within a given image exhibit varying levels of reliability.
-For each object in the image, the optimal positive prediction is calibrated (i.e., reliable), while the remaining optimal negative predictions are not.
-IV-C Numerical Analysis
-(a) Object-level Calibration Error
-(b) Correlation to Image-level Reliability
-Figure 4: A visualization of the difference in calibration between positive and negative predictions on the COCO dataset.
-In Figure 4a, Object-Level Calibration Error (OCE) values of their confidence scores are shown, where a lower score represents better-calibrated predictions.
-In Figure 4b, Pearson Correlation Coefficient (PCC) values between the ground truth $\mathsf{ImReli}$ and the average confidence scores are shown, where a higher value is better.
+In this early stage, the distinction between optimal positive and negative queries can be ambiguous (e.g., Figure 3a and Figure 4a).
+However, the attentions through the subsequent decoder layers progressively refine these predictions.
+By the final layer, the model selects a single prediction and assigns a confidence score based on its understanding of the image and the object.
+In contrast, the confidence scores for remaining queries do not increase to the same extent as the positives and even decrease.
+On the other hand, in images of low reliability (where the model is uncertain), the confidence score of the optimal positive prediction does not significantly increase, while the scores of the optimal negative ones are either slightly raised or unchanged.
+Based on this observation, we present our main claim:
+DETR learns to assign multiple predictions for each object within a given image and to designate a single prediction that is calibrated, while the remaining predictions are suppressed to have low foreground confidence, even when maintaining accurate localization.
+When the model’s uncertainty about its prediction (and thus about the matching process) is high, the confidence separation between those predictions diminishes.
+(a) High-reliability Image
+(b) Low-reliability Image
+Figure 4: When the model is confident, it assigns a high confidence score to the optimal positive prediction by either increases or remains high across the decoder layers via cross-attention mechanism—thus be calibrated.
+While those of remaining optimal negative queries’ confidence scores are pushed to be near-zero while having accurate bounding box predictions.
+On the other hand, when the model is uncertain, DETR assigns a lower confidence score to the optimal positive prediction, thereby maintaining good calibration.
+However, conversely, it slightly increases or maintains the confidence scores for remaining predictions.
+IV-C Theoretical Analysis of the Optimal Prediction Strategy
+To provide theoretical support for the claim, we analyze the optimal prediction strategy in a stochastic setting.
+We define a random variable $y_{\text{gt}}=(c_{\text{gt}},b_{\text{gt}})$ representing the ground-truth object with distribution $Y$, where $c_{\text{gt}}\in\mathcal{C}$ represents the class that is sampled from the class distribution $p_{\text{gt}}$, and $b_{\text{gt}}$ represents the bounding box sampled from some distribution $\mathcal{B}$.
+The padded set of $M$ targets is denoted by $\{y_{j}\}_{j=1}^{M}$, where $y_{1}=y_{\text{gt}}$ (the sampled object) and $y_{j}=\emptyset$ for $j\neq 1$.
+The model provides a set of predictions $\{\hat{y}_{i}\}_{i=1}^{M}$.
+The objective is to find predictions that minimize the expected total loss over the distribution $Y$:
+$\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]=\mathbb{E}_{y_{\text{gt}}\sim Y}\left[\sum_{j=1}^{M}\mathcal{L}(\hat{y}_{\hat{\sigma}(j)},y_{j})\right].$
+To formalize the roles of the predictions, we define the index of the optimal positive query as $p=\hat{\sigma}(1)$, while the remaining M1M-1 indices, $\{n\}_{n\neq p}$, are the optimal negative queries.
+Given this definition, we can decompose the total loss into terms for the optimal positive and optimal negative indices:
+$\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]=\mathbb{E}_{y_{\text{gt}}\sim Y}\left[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})\right]+\sum_{n\neq p}\mathcal{L}(\hat{y}_{n},\emptyset).$
+The loss for the optimal positive prediction $\hat{y}_{p}=(\hat{p}_{p},\hat{b}_{p})$ matched to the $y_{\text{gt}}$ is:
+$\mathbb{E}\left[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})\right]$ $=\mathbb{E}_{y_{\text{gt}}\sim Y}\left[\mathcal{L}_{\text{box}}(\hat{b}_{p},b_{\text{gt}})-\log(\hat{p}_{p}(c_{\text{gt}}))\right]$
+$=\mathbb{E}_{b_{\text{gt}}\sim\mathcal{B}}\left[\mathcal{L}_{\text{box}}(\hat{b}_{p},b_{\text{gt}})\right]-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{p}_{p}(c)),$ (3)
+where $\hat{p}_{p}(c)$ is the predicted probability for class $c$, as defined before.
+The loss for optimal negative prediction $\hat{y}_{n}$ matched to $\emptyset$ (one of the optimal negatives) is the down-weighted negative log-probability of the background class:
+$\mathcal{L}(\hat{y}_{n},\emptyset)=-w_{\emptyset}\log(\hat{p}_{n}(\emptyset)).$ (4)
+Here, $w_{\emptyset}$ is the down-weighting factor (e.g., 0.1) for the background class. The following result establishes a lower bound on the expected total loss.
+Proposition 1.
+The minimum achievable expected total loss is lower bounded by the entropy of the ground-truth class distribution, i.e.,
+$\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]\geq\mathcal{H}(p_{\text{gt}})+\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})],$
+where $\mathcal{H}(p_{\text{gt}})=-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(p_{\text{gt}}(c))$.
+Proof.
+See Section B-A. ∎
+Next, we characterize the optimal prediction set that uniquely achieves this bound.
+Proposition 2 (Optimal Prediction Set).
+The unique globally optimal prediction set that minimizes $\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]$ is $\{(\hat{p}^{*}_{i},\hat{b}^{*}_{i})\}_{i=1}^{M}$, where
+  1. 1.
+    for the optimal positive prediction ($i=p$): $\hat{p}^{*}_{i}=p_{\text{gt}}$ (i.e., $\hat{p}^{*}_{i}(c)=p_{\text{gt}}(c)$ $\forall c\in\mathcal{C}$ and $\hat{p}^{*}_{i}(\emptyset)=0$) and the box $\hat{b}^{*}_{i}$ minimizes $\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b}_{i},b_{\text{gt}})]$, and
+  2. 2.
+    for any optimal negative prediction ($i\neq p$): $\hat{p}^{*}_{i}(c)=0$ $\forall c\in\mathcal{C}$ and $\hat{p}^{*}_{i}(\emptyset)=1$.
+Proof.
+See Section B-B. ∎
+From Optimal Set to Strategy under Ambiguity.
+The previous propositions formally show that the Hungarian loss incentivizes a deterministic specialization.
+The optimal state is not a symmetric hedge where multiple queries compete, but for one query to emerge as a specialist that becomes a perfectly calibrated predictor of the object’s true class distribution ($p_{\text{gt}}$).
+All other queries are simultaneously optimized to predict the background class, minimizing their own loss to zero.
+This finding implies that the model is incentivized to develop an internal strategy to achieve this state.
+Ideally, the model would learn to designate a specific object query (which we conceptually refer to as the primary query) to become the optimal positive prediction $p$, and the remaining ones (secondary queries) to become the optimal negatives.
+As we discussed in Section IV-B, this is architecturally feasible, as DETR can leverage its attention mechanism to orchestrate predictions: Throughout the decoder layers, the model can identify the most promising object query and progressively specialize it, while simultaneously suppressing the others.
+However, in practice, a critical ambiguity arises.
+While the Hungarian matching algorithm $\hat{\sigma}$ is deterministic for a deterministic annotation, the ground truth itself is often ambiguous.
+Consequently, the assignment of queries to targets becomes a stochastic process.
+We formalize this by defining $P_{i}$ as the probability that the $i$-th query is matched to the ground-truth object:
+$P_{i}\triangleq\mathbb{P}_{y_{\text{gt}}\sim Y}(\hat{\sigma}(1)=i).$ (5)
+In this setting, the extreme specialization is now suboptimal because it incurs an infinite loss if the primary prediction is matched to the background.
+Therefore, to minimize its expected loss under this uncertainty, each query must find an optimal prediction $\hat{p}_{i}^{*}$ that “hedges” against this matching uncertainty.
+The expected loss $J_{i}$ for query $i$ is decomposed as:
+$J_{i}$ $=\underbrace{P_{i}\cdot\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b}_{i},b_{\text{gt}})\mid\hat{\sigma}(1)=i]}_{\text{Expected Box Loss Term}}$
+$\quad\underbrace{-P_{i}\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{p}_{i}(c))-(1-P_{i})w_{\emptyset}\log(\hat{p}_{i}(\emptyset))}_{\text{Expected Classification Objective }J(\hat{p}_{i})}.$
+Minimizing $J_{i}$ with respect to $\hat{p}_{i}$ is equivalent to minimizing the expected classification objective $J(\hat{p}_{i})$.
+Proposition 3 characterizes the optimal strategy to minimize the objective.
+Proposition 3 (Optimal Strategy under Uncertainty).
+For a query with a matching probability $P_{i}$ to the ground-truth object whose class distribution is $p_{\text{gt}}$, the class prediction that minimizes the expected loss is $\hat{p}_{i}^{*}$, where
+  1. 1.
+    the prediction for any foreground class $c\in\mathcal{C}$ is:
+    $\hat{p}_{i}^{*}(c)=p_{\text{fg}}^{*}(P_{i})p_{\text{gt}}(c).$ (6)
+  2. 2.
+    the prediction for the background class is
+    $\hat{p}_{i}^{*}(\emptyset)=1-p_{\text{fg}}^{*}(P_{i}),$
+where $p_{\text{fg}}^{*}(P_{i})$ is the optimal foreground probability, defined as
+$p_{\text{fg}}^{*}(P_{i})=\frac{P_{i}}{P_{i}+w_{\emptyset}(1-P_{i})}.$ (7)
+Proof.
+See Section B-C. ∎
+Based on the result from Proposition 3, we analyze two characteristics in their prediction strategy:
+  • Perfect Specialization: Consider the scenario where the model is highly certain about the image context and thus about which query will be matched.
+In this scenario, the matching probability for the primary—which will eventually become the optimal positive—query index is perceived as 1 (i.e., $P_{p}\to 1$), while it vanishes for all others (i.e., $P_{i}\to 0$ for $i\neq p$).
+In this scenario, the strategy recovers the optimal prediction set (Proposition 2), where solely the $p$-th query becomes a perfectly calibrated predictor of the object ($p_{\text{gt}}$).
+  • Hedging: When the model is uncertain about the matching query (i.e., $P_{i}$ is dispersed), it optimally “hedges.” The foreground probability for the primary query is suppressed ($p_{\text{fg}}^{*}(P_{p})<1$), balancing the risk between foreground and background assignments.
+In practice, matching probabilities are skewed: $P_{p}$ is close to 1 for the primary query, and $P_{i}$ is close to 0 for secondary queries: $i\neq p$.
+Thus, DETRs prioritize specialization while retaining a hedging mechanism proportional to their uncertainty.
+We term this the Specialist Strategy.
+This jives with the empirical trends observed in Figures 3 and 4.
+Figure 5: Empirical cumulative distribution function (ECDF) plots for maximum confidence score over foreground classes and maximum IoU over ground truth objects, comparing optimal positive predictions (blue) and optimal negative predictions (red), in UP-DETR and Cal-DETR on the COCO dataset.
+(a, b) Confidence: The optimal negative prediction confidence is overwhelmingly clustered at zero.
+Optimal positive predictions span the full confidence range, acting as the system’s calibrated predictors.
+(c, d) Localization: Optimal positive predictions are highly accurate (high IoU), which is the necessary condition for the model to designate a primary query.
+Optimal negative predictions maintain decent localization, which is not a failure but the optimal risk-averse strategy.
+IV-D Numerical Analysis
 Setup.
-To provide quantitative support for our claim regarding the varying levels of reliability across predictions, we conducted experiments using four DETR variants: UP-DETR [35], Deformable-DETR (D-DETR), Cal-DETR [12], and DINO.
-Each model is trained on the COCO (train2017) dataset and we use $1,000$ images (i.e., $20\%$) of COCO (val2017) for the validation set and the remaining images for the test set.
-The model is tested on three datasets with varying levels of out-of-distribution (OOD) characteristics: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD).
-Further details can be found in the Appendix.
-Object-level Calibration Error.
-We evaluate and compare the object-level calibration errors on the optimal positive predictions and the optimal negative predictions.
-The optimal positives and negatives are determined using the previously mentioned bipartite matching algorithm.
-As illustrated in Figure 3, the optimal negative predictions exhibit a significantly high calibration error, driven by consistently low confidence scores regardless of their actual accuracy.
-Furthermore, as shown in Figure 4a, the optimal positive predictions are significantly better calibrated than the negative ones.
-This trend is evident across different state-of-the-art DETR variants and datasets, all of which exhibit poor calibration quality for optimal negative predictions.
-Correlation to Image-Level Reliability.
-We compute $\mathsf{ImReli}$ for each image by using the same COCO evaluator [18] to obtain the image-wise AP score for $\mathsf{Perf}$ in Equation 1, passing the predictions and annotations for each image individually rather than the entire image set.
-Then, we measure the Pearson correlation coefficient (PCC) between the $\mathsf{ImReli}$ and the average confidence score of the optimal positives and the optimal negatives, respectively.
-As shown in Figure 4b, the average confidence scores of the optimal positive predictions exhibit a moderately strong positive correlation with image-level reliability (i.e., the model’s actual average precision on that image) across different models.
-Notably, the average confidence scores for negative predictions are inversely correlated with image-level reliability.
-This finding reinforces our claim, highlighting the importance of distinguishing between positive and negative predictions, especially when conducting image-level UQ in DETR.
-Further visualizations for the other datasets are provided in the Appendix.
-V A Systematic Framework for Identifying Positive Predictions
-As illustrated so far, using positive predictions is crucial for the reliable use of DETR.
-Nonetheless, due to the lack of annotations, it is not feasible to identify either the ground truth or the optimal positive predictions during the inference stage.
-Therefore, an alternative systematic framework is essential not only for improving its interpretability but also for ensuring reliability.
-Given that the optimal matching achieves a small calibration error, we use calibration error to measure the quality of a separation scheme, i.e., a separation scheme with a lower calibration error is deemed a better scheme.
+To provide quantitative support for our claim, we conducted experiments using four DETR variants: UP-DETR [35], Deformable-DETR (D-DETR) [14], Cal-DETR [12], and DINO [15].
+We also investigate with Faster R-CNN [2] and YOLOv3 [redmon2018yolov3, 3].
+Each model is trained on the COCO (train2017) dataset and we use $1,000$ images (i.e., $20\%$) of COCO (val2017) for the validation set.
+Empirical Analysis of Prediction-Level Distributions.
+We now empirically analyze the distributions of both confidence and localization accuracy to validate our theoretical findings.
+Ideally, this analysis would compare the primary prediction (the query intended to be the optimal positive) against secondary predictions.
+However, since the model’s internal designation is latent and not directly observable, we utilize the outcomes of the Hungarian matching as effective proxies.
+Specifically, we analyze optimal positive predictions (those matched to the ground truth) as the observable realization of primary prediction, and optimal negative predictions (those matched to the background) as the proxy for secondary predictions.
+First, we investigate the confidence scores.
+We plot the empirical cumulative distribution function (ECDF) of the maximum foreground probability ($\max_{c\neq\emptyset}\hat{p}_{i}({c})$) in Figures 5a and 5b.
+This indicates that the optimal negative predictions’ confidence scores are heavily concentrated in the near-zero regime, effectively being suppressed, in contrast to the widely dispersed confidence distribution observed for optimal positive predictions.
+Next, we perform a parallel analysis on localization accuracy in Figures 5c and 5d.
+More specifically, for each prediction, we measure the maximum IoU with respect to the ground truth objects in the image.
+The results for optimal positive predictions confirm that the vast majority exhibit high accuracy.
+On the other hand, the analysis of optimal negative predictions reveals a critical distinction: despite their suppressed confidence, they frequently maintain substantial localization accuracy.
+It reflects a risk-averse strategy where the model maintains precise backup bounding boxes to mitigate the high penalty of potential matching errors, ensuring training stability.
+As shown in Figure 13, these observed trends are consistent across various DETR variants and other detectors.
+We remark that this is not a model failure, but rather a risk-averse strategy induced by the loss landscape.
+In a stochastic setting where $y_{\text{gt}}$ varies, there is always a non-zero probability that secondary queries get matched to the ground-truth object.
+If the secondary query predicts a random box, its matching to the foreground object elevates the expected box loss.
+Thus, to minimize the expected total loss, the model is incentivized to set secondary queries to predict a decent bounding box rather than a random box.
+Figure 6: Flowchart of our evaluation framework.
+Two components vary across configurations: (1) the model ($\theta$), which generates a full set of predictions ($\hat{\mathcal{Y}}_{\theta}(x)$), and (2) the post-processing algorithm ($\mathcal{A}$), which selects a positive set ($\hat{\mathcal{S}}_{\theta}(x)$) from these predictions.
+In addition, we can obtain the optimal positive sets ($\mathcal{S}_{\theta}(x)$) by passing the full prediction set ($\hat{\mathcal{Y}}_{\theta}(x)$) and ground-truth object annotations ($\mathcal{Y}$) to the Hungarian Algorithm.
+Figure 7: Flowchart of our method for evaluating calibration error.
+A calibration metric ($\mathcal{M}$) compares either given positive sets ($\hat{\mathcal{S}}_{\theta}(\mathcal{X}_{\text{test}})$) or optimal positive sets (${\mathcal{S}}_{\theta}(\mathcal{X}_{\text{test}})$) with ground-truth object annotations ($\mathcal{Y}_{\text{test}}$) to get calibration error ($\mathcal{M}(\hat{\mathcal{S}}_{\theta}(\mathcal{X}_{\text{test}}),\mathcal{Y}_{\text{test}})$) or optimal calibration error ($\mathcal{M}({\mathcal{S}}_{\theta}(\mathcal{X}_{\text{test}}),\mathcal{Y}_{\text{test}})$), respectively.
+V Quantifying Calibration in DETR
+V-A Implications of the Specialist Strategy for Evaluation
+Our analysis reveals that DETR’s training objective incentivizes a specialist strategy. This strategy assigns two fundamentally different roles to the predictions:
+  1. 1.
+    A small subset of primary predictions (intended to be matched with foreground objects) is trained to be calibrated predictors, matching the true class distribution $p_{G}$.
+  2. 2.
+    The vast majority of secondary predictions are trained to be background classifiers, intentionally suppressing their foreground confidence to near-zero.
+This specialist strategy introduces a fundamental ambiguity in downstream decision-making.
+When a prediction exhibits low confidence, it becomes indistinguishable: is it a positive prediction correctly expressing high uncertainty about an ambiguous object, or is it a secondary prediction whose confidence has been artificially suppressed?
+This distinction is paramount, particularly in safety-critical domains.
+Safety-Critical Implication For an autonomous vehicle, a low-confidence “person” prediction could be a truly ambiguous object (like a person depicted on a billboard or on a vehicle), which the model is correctly uncertain about.
+However, it could also be a definite pedestrian identified by a secondary prediction, whose confidence was artificially suppressed to near-zero as part of the model’s behavior, resulting from optimizing the Hungarian loss (as discussed in Section IV).
+In the latter case, relying on this suppressed score could lead the system to catastrophically misjudge the risk posed by a real person.
+Ideally, at test time, we aim to filter out the secondary predictions and retain only the primary ones.
+However, identifying the primary predictions is non-trivial because the model’s internal designation is latent.
+Consequently, practitioners rely on a post-processing algorithm (e.g., confidence thresholding) to select a subset of predictions serving as an estimate of the primary set.
+While one might consider evaluating calibration directly on the optimal positive predictions, identifying this set requires ground-truth labels, making it infeasible at test time.
+Evaluating against this oracle set is useful for analysis, but it does not perfectly reflect the actual deployment conditions, where the system must rely solely on a post-processed subset.
+These issues create a critical challenge for practical deployment: we must evaluate calibration quality using the same post-processing algorithm that will be applied at test time.
+Therefore, practitioners should assess the entire (model + post-processing) pipeline jointly to identify configurations that produce reliable positive sets, as illustrated in Figures 6 and 7, rather than evaluating the model in isolation.
+To address this, we propose a new calibration metric that takes as input any positive set produced by a post-processing algorithm, enabling such joint evaluation.
+V-B Proposed Metric: Object-Level Calibration Error
+Notation.
+For each image $x_{l}$ having $N_{l}$ objects, we consider a set of ground-truth object annotations $\mathcal{Y}_{l}=\{y_{l,j}=(c_{l,j},b_{l,j})\}_{j=1}^{N_{l}}$, and a set of DETR predictions: $\hat{\mathcal{Y}}_{\theta}(x_{l})=\{\hat{y}_{l,i}=(\hat{p}_{l,i},\hat{b}_{l,i})\}_{i=1}^{M}$ where $M$ represents the number of object queries.
+Let $\mathcal{Y}_{\text{test}}=\big(\mathcal{Y}_{l}\big)_{l=1}^{N_{\text{test}}}$ denote the set of all ground-truth objects across all $N_{\text{test}}$ test images $\mathcal{X}_{\text{test}}=\big(x_{l}\big)_{l=1}^{N_{\text{test}}}$.
+Definition 1.
+Consider a subset of predictions, $\hat{\mathcal{S}}_{\theta}(x_{l})\subseteq\hat{\mathcal{Y}}_{\theta}(x_{l})$, that is generated by post-processing algorithm $\mathcal{A}$ from the full prediction set: $\hat{\mathcal{S}}_{\theta}=\mathcal{A}\circ\hat{\mathcal{Y}}_{\theta}$.
+We define an object-level calibration error (OCE) as the average Brier score per object:
+$\mathrm{OCE}_{\tau}\big(\hat{\mathcal{S}}_{\theta},\mathcal{Y}_{\text{test}}\big)=\frac{1}{|\mathcal{Y}_{\text{test}}|}\sum_{y_{l,j}\in\mathcal{Y}_{\text{test}}}\mathrm{Brier}_{\tau}\big(\hat{\mathcal{S}}_{\theta}(x_{l}),y_{l,j}\big),$ (8)
+$\mathrm{Brier}_{\tau}\big(\hat{\mathcal{S}}_{\theta}(x_{l}),y_{l,j}\big)=\sum_{c=1}^{C}\big(\mathbbm{1}(c=c_{l,j})-\bar{p}^{(\tau)}_{l,j}(c)\big)^{2},$
+$\bar{p}^{(\tau)}_{l,j}(c)=\frac{1}{|\mathcal{Q}^{(\tau)}_{l,j}|}\sum_{q\in\mathcal{Q}^{(\tau)}_{l,j}}\hat{p}_{l,i}(c),$
+where $\mathcal{Q}^{(\tau)}_{l,j}$ is a set of prediction indices that has bounding box overlap with the ground-truth object $y_{l,j}$:
+$\mathcal{Q}^{(\tau)}_{l,j}$ $\triangleq\{i\mid\text{IoU}(b_{l,j},\hat{b}_{l,i})\geq\tau\}.$
+When $|\mathcal{Q}^{(\tau)}_{l,j}|=0$, we consider the predicted probability to be zero, thus the corresponding Brier score is estimated as $1.0$.
+Following [18, 20], we use IoU thresholds of $\tau=0.5,0.75$ and report the average score as $\mathrm{OCE}(\cdot)=\mathrm{OCE}_{0.5}(\cdot)+\mathrm{OCE}_{0.75}(\cdot)$.
+The introduced object-level calibration error (OCE) has two desirable characteristics derived from its object-centric design.
+Unlike prediction-level metrics (e.g., D-ECE) that average errors over the selected predictions, OCE averages errors over the ground-truth objects.
+Consequently, the metric effectively penalizes two critical failure modes in recovering the primary predictions.
+(1) Retaining secondary predictions: OCE penalizes subsets containing secondary predictions—those that maintain systematically low confidence despite accurate localization—as they fail to contribute correctly to the assigned objects.
+(2) Missing primary predictions: OCE penalizes subsets that are too aggressive and discard primary predictions, leaving ground-truth objects unmatched and incurring a maximum penalty.
+Thus, unlike existing prediction-level ECE metrics, OCE effectively measures how well the post-processed subset approximates the set of reliable predictions.
+Remark (complexity).
+Computing OCE only requires IoU checks between GT boxes and predicted boxes; precomputing an $N_{i}\times M$ IoU matrix per image makes the additional cost negligible relative to standard evaluation.
+V-C Numerical Analysis
+Table 1: The Pearson correlation coefficient between the proposed OCE and established calibration metrics, D-ECE and LA-ECE0, on the optimal positive predictions () across six models—UP-DETR, D-DETR, Cal-DETR, DINO, Faster R-CNN, and YOLOv3—and three data distributions.
+The consistently very strong correlations confirm that OCE effectively captures the same underlying concept of calibration quality as D-ECE and LA-ECE0.
+Predictions Metrics COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD) Average $\pm$ Std.
+[…]
+Optimal Positives D-ECE 1.000 0.983 0.791 1.000 0.963 0.967 1.000 0.890 0.989 0.954 $\pm$ 0.066
+[…]
+$\mathsf{OCE}^{*}$ 0.791 0.859 1.000 0.967 0.873 1.000 0.989 0.854 1.000 0.926 $\pm$ 0.076
+V-C1 Validating OCE as a Calibration Metric
+Our first step is to validate that OCE is a rightful measure of calibration.
+We first perform a sanity check to ensure OCE fundamentally measures the same concept of calibration as established metrics.
+To do this, we isolate the metrics’ behaviors from the confounding effects of different post-processing decisions.
+We therefore choose the optimal positive set.
+This set best represents the predictions that each model was specifically trained to produce as a calibrated predictor.
+Hypothesis.
+If OCE is a valid calibration metric, its ranking of models on this oracle optimal positive set should strongly correlate with the rankings produced by established metrics (D-ECE, LA-ECE0).
+A lack of correlation would suggest OCE is measuring a completely different, unrelated property.
+Method.
+We evaluate the calibration errors on the optimal positive set from our six models—UP-DETR, D-DETR, Cal-DETR, DINO, Faster-RCNN, and YOLOv3—using OCE, D-ECE, and LA-ECE0.
+We then evaluate the models using each metric and compute the Pearson correlation coefficient (PCC) between metrics.
+As shown in the first three rows in Table 1, the correlations are all extremely high.
+This robust finding provides a critical validation: OCE captures the same underlying concept of calibration quality as established metrics.
+This confirms OCE is a legitimate calibration metric, providing the necessary foundation for the next section, where we will demonstrate its superiority in the practical scenario where optimal positive set is not accessible.
+V-C2 Pitfalls of Existing Metrics as a Reliable Set Selection Criteria
+We now analyze the behavior of existing metrics and the proposed OCE in a practical scenario: selecting a confidence threshold (i.e., post-processing configuration) that achieves the lowest calibration error for a given model.
+To this end, this section analyzes the effectiveness of existing calibration metrics (D-ECE, LA-ECE0, and our OCE) as well as performance metrics (AP [16, 17, 18] and Localization recall precision (LRP) [21, 36]) for this task.
+Specifically, for each of the four different DETR models, we sweep the confidence score threshold from $0.0$ to $1.0$.
+This generates different positive subsets of predictions.
+We then plot the score of each metric for these subsets to determine the threshold that each metric prefers (i.e., the one that yields its optimal score).
+Results on Cal-DETR are illustrated in Figure 8.
+For definitions of metrics and additional results on other DETR models, please refer to the Appendix C.
+Our empirical findings reveal the following behaviors:
+Pitfall 1: Average Precision Prefer No Filtering.
+As noted in several studies [21, 19, 20, 36], the optimal AP is achieved when the threshold is set close to $0.0$.
+This is because AP does not harshly penalize low-confidence predictions.
+However, as discussed so far, using the full prediction set (hundreds of predictions) is practically non-sensical.
+It carries a high risk of including uncalibrated negative predictions, diminishes interpretability, and leads to unreliable downstream decisions.
+Pitfall 2: ECEs Prefer Discarding Almost All Predictions.
+More importantly, the optimal ECEs (D-ECE and LA-ECE0) are achieved when the threshold is set close to $1.0$.
+This behavior, which favors retaining only a tiny subset of high-confidence predictions, is a structural pitfall of prediction-level ECEs [20].
+The cause is simple: ECEs do not penalize missed predictions (i.e., false negatives).
+As a result, they can achieve a near-zero (perfect) error by evaluating a set containing only a few highly accurate and confident predictions, while ignoring all the ground-truth objects the model missed at that high threshold.
+Therefore, this pitfall makes ECEs fundamentally unsuitable for the goal we defined: finding a reliable (model + post-processing) combination.
+If a practitioner uses D-ECE to select a threshold, the metric will misleadingly guide them to select a threshold near 1.0, which is nonsensical as it discards almost all predictions to achieve a low error.
 Figure 8: Impact of confidence threshold selection on various performance and error metrics in Cal-DETR evaluated on COCO.
 Higher scores are preferable ($\uparrow$), while lower scores are preferable ($\downarrow$).
-V-A Limitations of Existing Metrics
-To this end, this section analyzes the effectiveness of existing metrics, including AP, D-ECE, LA-ECE0, and localization recall precision (LRP) [21, 36], for identifying positive predictions.
-Specifically, we sweep the threshold for the confidence score to generate different subsets of DETR predictions (i.e., the predicted positives).
-We then assess the performance of each metric for these subsets and determine the threshold value that yields the highest performance.
-Analytical results using Cal-DETR are illustrated in Figure 5.
-For definitions of AP and LRP and additional results on the other DETR models, please refer to the Appendix.
-Primarily, as noted in several studies [21, 19, 20, 36] and our empirical findings, the optimal AP is achieved when the threshold is set to $0.0$ because AP does not penalize harshly for having low confidence predictions.
-However, as discussed earlier, using the entire prediction set carries a high risk of including uncalibrated negatives, leading to unreliable decisions in practical applications.
-Furthermore, using hundreds of predictions diminishes the interpretability of the model.
-In contrast, the optimal ECEs are often achieved when the threshold is set close to $1.0$, meaning ECEs favor retaining fewer predictions with high confidence.
-This is a structural pitfall that prediction-level ECEs commonly face [20].
-Since ECEs do not penalize missed detections (i.e., false negatives), they could achieve near-zero error when the evaluation positive set consists solely of highly accurate and confident predictions.
-Therefore, unless the model is trained to be excessively overconfident—which is unlikely given that DETR is trained on large datasets using various auxiliary loss functions—ECEs can result in very small error values when paired with a large confidence threshold.
-On the other hand, LRP, a localization-focused performance metric, could be used instead.
-However, LRP is not designed as a calibration metric so does not explicitly consider calibration error.
-As a result, the model ranking scored by LRP across the different models is not necessarily aligned with the model’s calibration qualities, as shown in the following numerical analysis.
-V-B Proposed Metric: Object-Level Calibration Error
-Notation.
-For each image $\bm{x}_{i}$ having $N_{i}$ objects, we consider a set of ground truth object annotations $\mathcal{D}_{i}=\{d_{i,j}=(l_{i,j},\bm{b}_{i,j})\}_{j=1}^{N_{i}}$, and a set of DETR predictions: $\hat{\mathcal{D}}_{\theta}(\bm{x}_{i})=\{\hat{d}_{i,q}=(\hat{\bm{p}}_{i,q},\hat{\bm{b}}_{i,q})\}_{q=1}^{M}$ where $M$ represents the number of object queries.
-Here, $l_{i,j}\in[1,C]$ and $\hat{\bm{p}}_{i,q}\in\{0,1\}^{C}$ represents the ground-truth label and predicted probability distribution over $C$ classes, respectively, while $\bm{b}_{i,j}$ and $\hat{\bm{b}}_{i,q}\in\{0,1\}^{4}$ corresponds to the ground truth and predicted scaled bounding box, respectively.
-Definition 2.
-Consider a subset of predictions, $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\subseteq\hat{\mathcal{D}}_{\theta}(\bm{x}_{i})$, that is generated by post-processing algorithm $\mathcal{A}$ from the entire prediction set: $\hat{\mathcal{S}}_{\theta}=\mathcal{A}\circ\hat{\mathcal{D}}_{\theta}$.
-We define an object-level calibration error (OCE) as the average Brier score per object:
-$\mathsf{OCE}\big{(}\hat{\mathcal{S}}_{\theta};\mathcal{I}\big{)}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(i,j)\in\mathcal{I}}\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});d_{i,j}\big{)}$ (6)
-$\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});d_{i,j}\big{)}=\sum_{c=1}^{C}\big{(}\mathbbm{1}(c=l_{i,j})-\bar{p}_{i,j}[c]\big{)}^{2}$ (7)
-$\bar{p}_{i,j}[c]=\frac{1}{|\mathcal{Q}_{i,j}|}\sum_{q\in\mathcal{Q}_{i,j}}\hat{\bm{p}}_{i,q}[c]$ (8)
-where $\mathcal{I}=\cup_{i}\{(i,j)\}_{j=1}^{N_{i}}$ is a set of all objects indices and $\hat{\bm{p}}(\cdot)[c]$ outputs the probability of $c$-th class; $\mathcal{Q}_{i,j}$ is a set of query indices that matches to the ground truth object $d_{i,j}$ and we propose two variants:
-$\mathcal{Q}_{i,j}$$\triangleq\{q\mid\text{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\geq\delta\}~{}$ $(\mathsf{OCE_{ENS,\delta}})$ (9)
-$\mathcal{Q}_{i,j}$$\triangleq\{q=\text{argmax}_{q}\text{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\}~{}$ $(\mathsf{OCE_{MAX}})$ (10)
-The difference is that $\mathsf{OCE_{ENS,\delta}}$ ensembles the overlapping predictions, while $\mathsf{OCE_{MAX}}$ selects the prediction with the best bounding-box matching.
-When $|\mathcal{Q}_{i,j}|=0$, we consider the predicted probability to be zero, thus the corresponding Brier score is estimated as $1.0$.
-Following [18, 20], we use IoU thresholds of $\delta=0.5,0.75$ and report the average score as $\mathsf{OCE_{ENS}}$ or simply $\mathsf{OCE}$.
-The introduced calibration error has two desirable characteristics.
-First, the prediction set achieves the lowest calibration error when the predictions are well-calibrated to the respective closest ground truth objects.
-Second, it penalizes the subset, $\hat{\mathcal{S}}_{\theta}(\cdot)$, that includes missing ground truth objects.
-This is achievable due to the primary difference between our OCE and ECEs: while ECEs are evaluated based on selected predictions, OCE is assessed along the ground truth objects.
-This ensures that subsets containing a small set of highly precise predictions are not assigned an artificially low error, unlike D-ECE and LA-ECE metrics.
-Thus, this metric can effectively assess not only whether the given subsets $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})$ are reliable, but also whether they comprehensively capture all ground truth objects, providing richer information compared to existing metrics.
-V-C Numerical Analysis: Effectiveness of OCE
-Table 1: Comparison of the effectiveness of the proposed OCE and exisiting metrics, focusing on their correlation with the models’ overall calibration quality on optimal positive predictions (denoted by the superscript *).
-For each metric, different post-processing schemes (i.e., methods for identifying positive predictions) are applied and specified in parentheses.
-These include using the top 100 predictions or choosing predictions with confidence scores above the threshold that optimizes the respective metric.
-Correlations exceeding $0.9$ are highlighted in blue, while negative correlations are highlighted in red.
-Our OCE consistently provides strong correlations with optimal calibration errors and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.
-Methods COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD) Average
-[…]
-References D-ECE 1.000 0.995 0.969 1.000 0.996 0.993 1.000 0.873 0.996 0.980 $\pm$ 0.039
-[…]
-$\mathsf{OCE_{ENS{}}}^{*}$ 0.969 0.948 1.000 0.993 0.990 1.000 0.996 0.836 1.000 0.970 $\pm$ 0.050
-Baselines AP (Top-100) [17, 18] 0.552 0.583 0.343 0.779 0.784 0.700 0.597 0.843 0.523 0.634 $\pm$ 0.149
-AP@50 (Top-100) [18] 0.380 0.409 0.164 0.687 0.710 0.606 -0.360 0.079 -0.440 0.249 $\pm$ 0.402
-AP@75 (Top-100) [18] 0.652 0.680 0.458 0.830 0.829 0.757 0.765 0.930 0.705 0.734 $\pm$ 0.127
-LRP (LRP) [21] 0.103 0.051 0.338 -0.268 -0.272 -0.150 -0.257 -0.610 -0.170 -0.137 $\pm$ 0.257
-D-ECE (D-ECE) -0.757 -0.700 -0.772 0.365 0.284 0.337 0.309 0.436 0.336 -0.018 $\pm$ 0.514
-D-ECE (Top-100) -0.873 -0.833 -0.966 0.608 0.576 0.511 -0.254 0.250 -0.316 -0.144 $\pm$ 0.614
-D-ECE (0.3) [8] 0.905 0.939 0.827 0.982 0.967 0.990 0.964 0.799 0.981 0.928 $\pm$ 0.067
-D-ECE (LRP) 0.802 0.855 0.659 0.987 0.984 0.999 0.983 0.793 0.996 0.895 $\pm$ 0.116
-D-ECE ($\mathsf{OCE_{ENS}}$) 0.856 0.901 0.727 0.987 0.983 0.999 0.983 0.793 0.996 0.914 $\pm$ 0.095
-LA-ECE0 (LA-ECE0) -0.960 -0.930 -0.968 -0.913 -0.874 -0.907 -0.826 -0.839 -0.835 -0.895 $\pm$ 0.051
-LA-ECE0 (Top-100) -0.944 -0.913 -0.994 -0.431 -0.482 -0.517 -0.786 -0.386 -0.822 -0.697 $\pm$ 0.228
-LA-ECE0 (0.3) -0.789 -0.729 -0.842 -0.780 -0.818 -0.830 -0.451 -0.041 -0.468 -0.639 $\pm$ 0.254
-LA-ECE0 (LRP) [20] -0.908 -0.864 -0.948 -0.833 -0.877 -0.846 -0.901 -0.597 -0.913 -0.854 $\pm$ 0.097
-LA-ECE0 ($\mathsf{OCE_{ENS}}$) -0.901 -0.855 -0.943 -0.723 -0.780 -0.739 -0.884 -0.568 -0.897 -0.810 $\pm$ 0.112
-Ours $\mathsf{OCE_{ENS}}$ ($\mathsf{OCE_{ENS}}$) 0.962 0.940 1.000 0.987 0.983 0.999 0.964 0.737 0.984 0.951 $\pm$ 0.078
-$\mathsf{OCE_{MAX}}$ ($\mathsf{OCE_{MAX}}$) 0.954 0.929 0.998 0.954 0.958 0.983 0.955 0.697 0.976 0.934 $\pm$ 0.086
-To showcase OCE’s effectiveness in assessing calibration quality, we analyze how different metrics rank calibration errors across four DETR models.
-As baselines, we measure ECEs using the top $100$ detections  a standard confidence threshold of 0.3, and the threshold that minimizes the different metrics on the validation set.
-In a similar manner, we assess our OCEs using thresholds that minimize each OCE metric on the validation set.
-For a fair comparison, we execute the optimal matching process on each test dataset to obtain the optimal positive set.
-We then evaluate D-ECE, LA-ECE0, and OCE using these sets as reference scores and calculate the PCC correlation to the other scoring schemes across different datasets.
-Table 1 presents the correlations between the three calibration metrics on the optimal positive set and various methods.
-The first three rows demonstrate that the calibration metrics—D-ECE, LA-ECE, and our proposed OCE—are highly correlated.
-This indicates that all three metrics effectively capture the notion of calibration quality.
-Additionally, these metrics show a decent correlation with AP metrics.
-It is important to note that AP primarily accounts for accuracy rather than calibration quality, which may explain why the correlation, particularly for AP@50 on the out-of-distribution dataset, is not perfectly aligned.
-LRP, D-ECE, and LA-ECE0 exhibit inverse correlations with calibration quality when used alone.
-Furthermore, LA-ECE0 with a non-optimal positive set also shows an inverse correlation with the overall calibration errors estimated on the optimal positive set.
-This supports our assertion that these metrics are inadequate for measuring models’ calibration quality in conjunction with the employed post-processing scheme, unlike our OCE.
-Another important attribute of OCE is that it can not only compare different models but also identify the most reliable configuration of the same model.
-As illustrated in Figure 5, the confidence threshold vs.
-OCE curve exhibits a bell shape, unlike ECE metrics, and the optimal OCE is achieved around a threshold of $0.3$.
-This is particularly notable because it reflects the practical choices in many calibration studies to date [12, 20], affirming the reliability of using OCE.
-The D-ECE and OCE scores measured at the optimal LRP threshold—denoted by D-ECE (LRP) and OCE (LRP), respectively—show a strong correlation to those computed from the optimal positive set; thus, the method of using LRP to identify a positive set [20] appears acceptable.
-Experimentally, we confirmed that OCE and LRP provide thresholds in a similar range across different settings.
-Nonetheless, the biggest advantage of using OCE over other metrics is that OCE measures the model’s calibration quality alongside the post-processing scheme employed.
-While measuring the calibration error on the optimal positive set would ideally assess quality, this optimal set is not accessible during test-time inference.
-Consequently, the score measured on the optimal set may not correspond to the quality experienced in real-world applications, such as those employing confidence thresholding schemes.
+A Note on LRP: Good for Localization, Not Calibration.
+On the other hand, LRP, a localization-focused performance metric, does appear to find a sensible, non-extreme optimal threshold.
+This is rational: optimal positives are trained to be well-localized predictors, while optimal negatives are relatively not.
+LRP, by focusing purely on localization quality, thus correctly identifies a threshold that filters out the poorly-localized predictions.
+However, LRP is not designed to measure calibration and does not consider the confidence quality of the predictions at all.
+As a result, a model ranking scored by LRP reflects localization performance, which is not necessarily aligned with the model’s confidence calibration qualities.
+V-C3 The effectiveness of OCE
+The biggest advantage of using OCE over other metrics is that OCE is capable of measuring the model’s calibration quality alongside the post-processing scheme employed.
+Resultingly, as illustrated in Figures 8 and 14, the confidence threshold vs.
+OCE curve is U-shaped and the optimal OCE is achieved around a threshold of $0.3$.
+This is particularly notable because it reflects the practical choices in many calibration studies to date [8, 9, 10, 11, 12, 13], confirming the reliability of using OCE.
 Using a fixed separation threshold of $0.3$ might serve as a reasonable approximation for distinguishing between optimal positives and negatives.
@@
 Consequently, employing a fixed threshold poses potential risks.
-Similarly, using a fixed number of top confident samples (e.g., Top-$100$) is inadequate, as it may include more predictions than the actual number of ground truth positives.
+(a) Threshold Optimizing AP
+(b) Threshold Optimizing OCE
+(c) Threshold Optimizing D-ECE
+Figure 9: Exemplary visualization demonstrating the impact of confidence threshold ($\rho$) selection on the final prediction subset in Cal-DETR across different post-processing configurations.
+Optimal positive and negative predictions are highlighted with blue and red boxes, respectively.
+The threshold selected by AP (e.g., $\rho=0.1$) retains redundant predictions, mixing calibrated positives with uncalibrated negatives.
+The threshold selected by D-ECE (e.g., $\rho=0.9$) discards nearly all detections, including the optimal positive prediction with moderate confidence, resulting in a missed detection.
+The proposed OCE identifies a practical threshold (e.g., $\rho=0.27$) that balances precision and recall while maintaining calibrated predictions.
+Figure 10: A visualization of the difference in calibration between optimal positive and negative predictions on the COCO dataset.
+Object-level Calibration Error (OCE) value of each set is shown, where a lower score represents better-calibrated predictions.
+As shown, OCE varies significantly between optimal positive and negative predictions, indicating that OCE is a useful metric for assessing the reliability of post-processing schemes.
+Figure 11: Pearson Correlation Coefficient (PCC) between per-image AP scores and per-image average foreground confidence scores on optimal positive and negative predictions (higher values indicate stronger correlation).
+As expected, optimal positive predictions show high positive correlation between confidence and performance.
+More interestingly, optimal negative predictions exhibit an inverse correlation with image-level reliability.
 VI Quantifying Image-level Reliability
-As demonstrated, positive and negative predictions exhibit varying levels of reliability.
-Interestingly, having predictions with low confidence scores does not necessarily imply low reliability.
-Our empirical observations show that confidence scores in negative predictions are actually inversely correlated with image-level reliability.
-More specifically, for a reliable instance, we observe that the confidence of positive predictions increases across the decoder layers, while that of negative predictions remains low; this results in a large gap between positive and negative predictions.
-In contrast, for unreliable instances (e.g., Figure 3a), the confidence score of the positive prediction does not increase across the layers, whereas the scores of negative predictions are either slightly elevated or remain unchanged.
-Consequently, there is a small gap between positive and negative predictions (e.g., Figure 3b).
-VI-A Proposed Method: Quantifying Reliability by Contrasting
-Based on the finding, we propose a post hoc UQ approach by contrasting the confidence scores of positives and negatives:
-$\mathsf{ContrastiveConf}(\bm{x})=\mathsf{Conf}^{+}(\bm{x})-\mathsf{Conf}^{-}(\bm{x})$ (11)
-$\mathsf{Conf}^{+}(\bm{x})=\frac{1}{|\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})|}\sum\nolimits_{(\hat{\bm{p}},\hat{\bm{b}})\in\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})}\max_{c}\hat{\bm{p}}[c]$ (12)
-$\mathsf{Conf}^{-}(\bm{x})=\frac{1}{|\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})|}\sum\nolimits_{(\hat{\bm{p}},\hat{\bm{b}})\in\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})}\max_{c}\hat{\bm{p}}[c]$ (13)
-where $\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})$ and $\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})$ are predicted sets of positive and negative predictions, respectively, and $\lambda$ is a scaling factor that can be determined on the validation set.
-We include an ablation study of the scaling factor in the subsequent section.
-Our results demonstrate that the proposed method is robust to the choice of scaling factor and consistently outperforms baseline approaches, achieving the robust performance with a scaling factor of $5.0$ $10.0$.
-Identifying a $\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})$ from $\hat{\mathcal{D}}_{\theta}(\bm{x})$ is a crucial factor in the success of this approach.
-In practice, however, neither the ground truth nor the optimal positives are available during the test time.
-Instead, we approximate the ground truth separation by applying a post-processing algorithm $\mathcal{A}^{*}$ that minimizes the calibration error on the validation set:
-$\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})=\mathcal{A}^{*}\circ\hat{\mathcal{D}}_{\theta}(\bm{x})$ (14)
-$\mathcal{A}^{*}=\text{argmin}_{\mathcal{A}}\mathsf{OCE}(\mathcal{A}\circ\hat{\mathcal{D}}_{\theta};\mathcal{I}_{val})$ (15)
-where $\mathcal{I}_{val}=\cup_{i\in\mathcal{V}}\{(i,j)\}_{j=1}^{N_{i}}$ is the set of all objects indices in the validation dataset $\mathcal{V}$.
+In this section, we revisit the importance of separation between primary and secondary predictions at test-time, now from the perspective of image-level reliability estimation.
+Recall that our theoretical analysis (Proposition 3) shows that when the model is uncertain about its prediction—expressed by matching probability $P_{i}<1$—it suppresses the predictive distribution.
+Reflecting this observation to the secondary prediction, we find that its foreground confidence score is no longer zero.
+Consequently, we expect the separation between these predictions to become blurred under uncertainty, and our empirical observations confirm this.
+Specifically, for reliable images, we observe that the confidence of optimal positive predictions increases across decoder layers while optimal negative predictions remain low, resulting in a large gap between optimal positive and negative predictions (e.g., Figure 4a).
+In contrast, for unreliable instances, optimal positive prediction confidence fails to increase across layers, while optimal negative prediction scores are either slightly elevated or unchanged.
+This produces a small gap between optimal positive and negative predictions (e.g., Figure 4b).
+To further empirically verify this, we first obtain the image-wise AP score by using the same COCO evaluator [18], passing the predictions and annotations for each image individually rather than the entire image set.
+We use this value as a proxy of image-level reliability (See Definition 2 for more formal interpretation.) Then, we measure the Pearson correlation coefficient between the image-wise AP and the average confidence score of the optimal positives and the optimal negatives, respectively.
+As shown in Figure 11, the average confidence scores of the optimal positive predictions exhibit a moderately strong positive correlation with image-level reliability (i.e., the model’s actual average precision on that image) across different models.
+Notably, the average confidence scores for negative predictions are inversely correlated with image-level reliability.
+This finding reinforces our claim, highlighting the importance of identifying primary and secondary predictions, when conducting image-level uncertainty quantification (UQ) in DETR.
+VI-A Image-Level Reliability
+We introduce a formal definition of image-level reliability by examining the model’s overall object detection performance on the image, extending [33, 29].
+Definition 2.
+We define image-level reliability as a measure of how closely the predictions align with the ground-truth annotations for a given test image $x$:
+$\mathsf{ImReli}(x;\theta)\triangleq\mathsf{Perf}\big(\hat{\mathcal{Y}}_{\theta}(x),~\mathcal{Y}\big),$
+where one can use any standard performance metric, such as average precision, for $\mathsf{Perf}$ depending on the user’s requirements.
+By its definition, image-level reliability directly addresses the model’s applicability to a given test instance.
+However, since image-level reliability requires ground-truth annotations for its determination, obtaining it during inference is not feasible.
+Therefore, this paper’s objective is to develop a method that assigns a quantitative score to each image instance, closely aligning with the model’s image-level reliability.
+VI-B Proposed Method: Quantifying Reliability by Contrasting
+Based on our findings, we propose a post hoc UQ approach by contrasting the confidence scores of positives and negatives:
+$\mathsf{ContrastiveConf}(x)=\mathsf{Conf}^{+}(x)-\lambda\mathsf{Conf}^{-}(x),$
+$\mathsf{Conf}^{+}(x)=\frac{1}{|\hat{\mathcal{S}}_{\theta}(x)|}\sum\nolimits_{(\hat{p},\hat{b})\in\hat{\mathcal{S}}_{\theta}(x)}\max_{c\in\mathcal{C}}\hat{p}(c),$
+$\mathsf{Conf}^{-}(x)=\frac{1}{|\hat{\mathcal{S}}_{\theta}^{-}(x)|}\sum\nolimits_{(\hat{p},\hat{b})\in\hat{\mathcal{S}}_{\theta}^{-}(x)}\max_{c\in\mathcal{C}}\hat{p}(c),$
+where $\hat{\mathcal{S}}_{\theta}(x)$ and $\hat{\mathcal{S}}_{\theta}^{-}(x)\triangleq\hat{\mathcal{Y}}_{\theta}(x)\setminus\hat{\mathcal{S}}_{\theta}(x)$ are predicted sets, which we refer to positive and negative predictions, respectively.
+$\lambda$ is a scaling factor that can be determined on the validation set; we include an ablation study of the scaling factor in the subsequent section.
+Our results demonstrate that the proposed method is robust to the choice of scaling factor and consistently outperforms baseline approaches, achieving the robust performance with a scaling factor of $5.0$ - $10.0$.
+Selecting positives at test time.
+Ground-truth matches are unavailable at test time, so we approximate the optimal positive set $\mathcal{S}_{\theta}(x)$ by applying a post-processing operator $\mathcal{A}$ (e.g., confidence thresholding, Top-$k$, or NMS).
+We choose $\mathcal{A}^{*}$ on the validation set to minimize OCE:
+$\mathcal{A}^{*}=\arg\min_{\mathcal{A}}~\mathrm{OCE}\big(\mathcal{A}\!\circ\!\hat{\mathcal{Y}}_{\theta}(\mathcal{X}_{\text{val}}),\,\mathcal{Y}_{\text{val}}\big),$
+where $\mathcal{Y}_{\text{val}}$ is the set of all ground-truth objects across the images $\mathcal{X}_{\text{val}}$ in validation dataset.
 Table 2: Comparison of the proposed method ($\mathsf{ContrastiveConf}$) with baseline methods on their Pearson correlation coefficient with the image-level reliability in Equation (2), with different models.
@@
 $\mathsf{ContrastiveConf}$ (OCE) 0.656 0.588 0.628 0.613 0.359 0.407 0.441 0.440 0.275 0.350 0.317 0.327
-VI-B Numerical Analysis
+VI-C Numerical Analysis
+The model is tested on three datasets with varying levels of out-of-distribution (OOD) characteristics: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD).
 We compare different methods for quantifying per-image reliability based on different separation methods.
 For separation methods, we apply the optimal matching (as a reference) as well as the practical post-processing schemes such as fixed and adaptive confidence thresholding.
-We measure the Pearson correlation coefficient of each method with the $\mathsf{ImReli}$ computed based on AP metric (see Appendix for the detail).
+We measure the Pearson correlation coefficient of each method with the $\mathsf{ImReli}$ computed using the standard AP metric.
 Table 2 shows that the proposed $\mathsf{ContrastiveConf}$ consistently achieves the best correlation with image-level reliability.
@@
 Figure 12: Impact of the scaling factor ($\lambda$) on image-level UQ performance of $\mathsf{ContrastiveConf}$ (OCE).
-Pearson correlation coefficient (PCC) using various scaling factors is reported.
+Pearson correlation coefficient (PCC) using various scaling factors is reported on (a) In-distribution (COCO), (b) Near OOD (Cityscapes), and (c) Far OOD (Foggy Cityscapes).
 The optimal scaling factor lies within the range of $5.0$ to $10.0$, while this range generalizes well across out-of-distribution datasets.
@@
 Furthermore, the results demonstrate that $\mathsf{ContrastiveConf}$ with $1.0\leq\lambda\leq 10.0$ consistently outperforms $\mathsf{Conf}^{+}$ (i.e., $\lambda=0.0$, the leftmost point on each line).
-However, we also observe a rapid drop in performance when the scaling factor is excessively large, particularly on the OOD dataset, Foggy Cityscapes.
-Thus, caution is needed when applying to substantially different datasets.
+That said, we also observe a rapid drop in performance when the scaling factor is excessively large, particularly on the OOD dataset, Foggy Cityscapes.
+Thus, we do not recommend using large scaling factors or relying solely on the negative predictions.
+VII Conclusion
+The main contribution of our work is an in-depth analysis of the reliability of DETR frameworks.
+In particular, we reveal that DETR’s predictions for a given image exhibit varying calibration quality, highlighting the importance of identifying well-calibrated predictions.
+To address this challenge, we introduced a systematic framework that leverages our object-level calibration error metric to discern these reliable predictions effectively.
+Furthermore, we proposed a novel uncertainty quantification method for estimating image-level reliability in DETR and conducted comparative studies of various post-processing schemes regarding their impact on DETR’s reliability.
+We hope our efforts expand the scope of DETR applications by enabling more precise and reliable deployment, sparking further research into the area of reliable positive prediction identification.
+Acknowledgments
+The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper.
+This work was supported in part by the MIT Summer Research Program (MSRP), the MIT-Google Program for Computing Innovation, the MIT-Amazon Science Hub, the MIT-IBM Watson AI Lab, and MathWorks.
+Appendix A Dataset
+For our experiments, we used the Cityscapes and Foggy Cityscapes datasets, which each have 500 images of first-person driving footage in realistic environments.
+Foggy Cityscapes has the same base images as Cityscapes, but with fog simulated and added to create a further out of distribution set.
+Since the DETR models were trained on COCO, the Cityscapes and Foggy Cityscapes annotations were converted to correspond to the labels of COCO.
+More specifically, the person, bicycle, car, motorcycle, bus, train, and truck classes were transferred directly.
+In addition, the rider class of Cityscapes and Foggy Cityscapes was mapped to the person class of COCO.
+The other classes present in Cityscapes and Foggy Cityscapes are largely focused on image segmentation, and thus were omitted (e.g.
+building, sky, sidewalk).
+The pre-trained model weights were obtained from their respective official implementations.
+Appendix B Detailed Theoretical Proofs
+B-A Proof of Proposition 1
+Proposition 1.
+The minimum achievable expected total loss is lower bounded by the entropy of the ground-truth class distribution, i.e.,
+$\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]\geq\mathcal{H}(p_{\text{gt}})+\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})],$
+where $\mathcal{H}(p_{\text{gt}})=-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(p_{\text{gt}}(c))$.
+Proof.
+The expected total loss is the sum of the expected loss for the optimal positive query matched to $y_{\text{gt}}$ (denoted as index $p$) and the expected losses for the M1M-1 remaining queries matched to $\emptyset$ (denoted as indices $\{n\neq p\}$):
+$\mathbb{E}[\mathcal{L}_{\text{total}}]=\mathbb{E}[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})]+\sum_{n\neq p}\mathcal{L}(\hat{y}_{n},\emptyset).$
+From Equation 4, $\mathcal{L}(\hat{y}_{n},\emptyset)=-w_{\emptyset}\log(\hat{p}_{n}(\emptyset))$; and since $\log(\hat{p}_{n}(\emptyset))\leq 0$: $\mathcal{L}(\hat{y}_{n},\emptyset)\geq 0$.
+Therefore, the total expected loss is bounded below by:
+$\mathbb{E}[\mathcal{L}_{\text{total}}]\geq\mathbb{E}[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})].$
+From Equation IV-C, we have:
+$\mathbb{E}[\mathcal{L}(\hat{y}_{p},y_{\text{gt}})]=\underbrace{-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{p}_{p}(c))}_{\mathbb{E}[\mathcal{L}_{\text{cls}}]}+\mathbb{E}_{b_{\text{gt}}\sim\mathcal{B}}[\mathcal{L}_{\text{box}}(\hat{b}_{p},b_{\text{gt}})].$
+We minimize the two terms on the right-hand side independently.
+First, we find the minimum classification loss $\mathbb{E}[\mathcal{L}_{\text{cls}}]$.
+We define $S=\sum_{c\in\mathcal{C}}\hat{p}_{p}(c)=1-\hat{p}_{p}(\emptyset)$ and $\hat{q}_{p}(c)=\hat{p}_{p}(c)/S$.
+Then:
+$\mathbb{E}[\mathcal{L}_{\text{cls}}]$ $=-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)(\log(S)+\log(\hat{q}_{p}(c)))$
+$=-\log(S)+\mathrm{CE}(p_{\text{gt}},\hat{q}_{p}),$
+where $\mathrm{CE}(p_{\text{gt}},\hat{q}_{p})=\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{q}_{p}(c))$ is the cross-entropy.
+By Gibbs’ inequality, $\mathrm{CE}(p_{\text{gt}},\hat{q}_{p})\geq\mathcal{H}(p_{\text{gt}})$ and equality holds if and only if $\hat{q}_{p}=p_{\text{gt}}$.
+Additionally, $\log(S)\geq 0$ and equality holds if and only if $S=1$ (i.e., $\hat{p}_{p}(\emptyset)=0$).
+Denoting the minimum box loss as $\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})]$, a lower bound of $\mathbb{E}_{b_{\text{gt}}\sim\mathcal{B}}[\mathcal{L}_{\text{box}}(\hat{b}_{p},b_{\text{gt}})]$, we have:
+$\mathbb{E}[\mathcal{L}_{\text{total}}]\geq\mathcal{H}(p_{\text{gt}})+\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})].$
+B-B Proof of Proposition 2
+Proposition 2 (Optimal Prediction Set).
+The unique globally optimal prediction set that minimizes $\mathbb{E}\left[\mathcal{L}_{\text{total}}\right]$ is $\{(\hat{p}^{*}_{i},\hat{b}^{*}_{i})\}_{i=1}^{M}$, where
+  1. 1.
+    for the optimal positive prediction ($i=p$): $\hat{p}^{*}_{i}=p_{\text{gt}}$ (i.e., $\hat{p}^{*}_{i}(c)=p_{\text{gt}}(c)$ $\forall c\in\mathcal{C}$ and $\hat{p}^{*}_{i}(\emptyset)=0$) and the box $\hat{b}^{*}_{i}$ minimizes $\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b}_{i},b_{\text{gt}})]$, and
+  2. 2.
+    for any optimal negative prediction ($i\neq p$): $\hat{p}^{*}_{i}(c)=0$ $\forall c\in\mathcal{C}$ and $\hat{p}^{*}_{i}(\emptyset)=1$.
+Proof.
+We evaluate the expected total loss for the proposed strategy $\{(\hat{p}^{*}_{i},\hat{b}^{*}_{i})\}_{i=1}^{M}$:
+  1. 1.
+    For the optimal positive prediction: The strategy sets $\hat{p}^{*}_{p}=p_{\text{gt}}$ and $\hat{b}^{*}_{p}$ to minimize the expected box loss. The resulting expected loss is:
+    $\mathbb{E}[\mathcal{L}(\hat{y}^{*}_{p},y_{\text{gt}})]$ $=\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})]-\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(p_{\text{gt}}(c))$
+    $=\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})]+\mathcal{H}(p_{\text{gt}}).$
+  2. 2.
+    For any optimal negative prediction: The strategy sets $\hat{p}^{*}_{n}(\emptyset)=1$. The resulting expected loss is:
+    $\mathcal{L}(\hat{y}^{*}_{n},\emptyset)=-w_{\emptyset}\log(1)=0.$
+Summing these terms, the strategy achieves the lower bound derived in Proposition 1:
+$\mathbb{E}[\mathcal{L}_{\text{total}}]=\min_{\hat{b}}\mathbb{E}[\mathcal{L}_{\text{box}}(\hat{b},b_{\text{gt}})]+\mathcal{H}(p_{\text{gt}}).$
+Any deviation from this strategy results in a strictly higher loss:
+  • If $\hat{p}_{n}(\emptyset)<1$, then $w_{\emptyset}\log(\hat{p}_{n}(\emptyset))>0$.
+  • If $\hat{b}_{p}$ is not the minimizer, the box loss term increases.
+  • If $\hat{p}_{p}\neq p_{\text{gt}}$, then either $\log(S)>0$ or $\mathrm{CE}(p_{\text{gt}},\hat{q}_{p})>\mathcal{H}(p_{\text{gt}})$ (by Gibbs’ inequality).
+Thus, the proposed prediction set is the unique global minimizer. ∎
+B-C Proof of Proposition 3
+Proposition 3 (Optimal Strategy under Uncertainty).
+For a query with a matching probability $P_{i}$ to the ground-truth object whose class distribution is $p_{\text{gt}}$, the class prediction that minimizes the expected loss is $\hat{p}_{i}^{*}$, where
+  1. 1.
+    the prediction for any foreground class $c\in\mathcal{C}$ is:
+    $\hat{p}_{i}^{*}(c)=p_{\text{fg}}^{*}(P_{i})p_{\text{gt}}(c).$ (9)
+  2. 2.
+    the prediction for the background class is
+    $\hat{p}_{i}^{*}(\emptyset)=1-p_{\text{fg}}^{*}(P_{i}),$
+where $p_{\text{fg}}^{*}(P_{i})$ is the optimal foreground probability, defined as
+$p_{\text{fg}}^{*}(P_{i})=\frac{P_{i}}{P_{i}+w_{\emptyset}(1-P_{i})}.$ (10)
+Proof.
+We find the optimal probability distribution $\hat{p}_{i}$ by minimizing the expected classification objective:
+$J(\hat{p}_{i})=-P_{i}\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)\log(\hat{p}_{i}(c))-w_{\emptyset}(1-P_{i})\log(\hat{p}_{i}(\emptyset)).$
+This is a constrained optimization problem, as we must satisfy the constraint that:
+$g(\hat{p}_{i})=\sum_{c\in\mathcal{C}}\hat{p}_{i}(c)+\hat{p}_{i}(\emptyset)-1=0.$
+We form the Lagrangian $\mathcal{J}$ with a Lagrange multiplier $\lambda$:
+$\mathcal{J}=J(\hat{p}_{i})-\lambda g(\hat{p}_{i}).$
+We find the optimum by setting the partial derivatives of $\mathcal{J}$ with respect to each probability $\hat{p}_{i}(c)$ to zero.
+For any foreground class $c\in\mathcal{C}$:
+$\frac{\partial\mathcal{J}}{\partial\hat{p}_{i}(c)}=0\implies\hat{p}_{i}(c)=-\frac{P_{i}\cdot p_{\text{gt}}(c)}{\lambda}.$
+For the background class $\emptyset$:
+$\frac{\partial\mathcal{J}}{\partial\hat{p}_{i}(\emptyset)}=0\implies\hat{p}_{i}(\emptyset)=-\frac{w_{\emptyset}(1-P_{i})}{\lambda}.$
+We solve for $\lambda$ by substituting these back into the constraint $g(\hat{p}_{i})=0$:
+$\sum_{c\in\mathcal{C}}\left(-\frac{P_{i}\cdot p_{\text{gt}}(c)}{\lambda}\right)+\left(-\frac{w_{\emptyset}(1-P_{i})}{\lambda}\right)=1.$
+Since $\sum_{c\in\mathcal{C}}p_{\text{gt}}(c)=1$ by definition:
+$\frac{P_{i}}{\lambda}-\frac{w_{\emptyset}(1-P_{i})}{\lambda}=1\implies\lambda=-(P_{i}+w_{\emptyset}(1-P_{i})).$
+Finally, we substitute this $\lambda$ back to find the optimal predictions $\hat{p}_{i}^{*}$:
+$\hat{p}_{i}^{*}(c)=-\frac{P_{i}\cdot p_{\text{gt}}(c)}{-(P_{i}+w_{\emptyset}(1-P_{i}))}=\left(\frac{P_{i}}{P_{i}+w_{\emptyset}(1-P_{i})}\right)p_{\text{gt}}(c),$
+and
+$\hat{p}_{i}^{*}(\emptyset)=-\frac{w_{\emptyset}(1-P_{i})}{-(P_{i}+w_{\emptyset}(1-P_{i}))}=1-\left(\frac{P_{i}}{P_{i}+w_{\emptyset}(1-P_{i})}\right).$
+Appendix C Omitted Experimental Results
+This section provides the omitted experimental results. Figure 14 extends Figure 8 across different DETR variants and datasets. Figure 13 extends Figure 5 across different DETR variants and datasets.
+Figure 13: Empirical cumulative distribution function (ECDF) plots for maximum confidence score and maximum IoU, comparing optimal positive predictions (blue) and optimal negative predictions (red) on the COCO dataset.
 Table 3: Comparison of post-processing schemes based on calibration quality (i.e., OCE).
@@
 We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding).
-For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses, and is applied on the test set.
+For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses and is applied on the test set.
 The strongest correlations are highlighted in bold.
@@
 NMS -0.227 (0.10) -0.157 (0.90) -0.402 (0.10) -0.140 (0.90) -0.044 (0.80) -0.034 (0.75) -0.203 (0.10) -0.080 (0.80) 0.035 (0.75) -0.035 (0.75) -0.149 (0.90) -0.075 (0.70)
-(a) Thresholding
-(b) Top-$k$
-(c) NMS
-Figure 7: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{Conf}^{+}$ and image-level reliability in Cal-DETR on COCO for different post-processing schemes.
-VII Comparative Study on Post-Processing
+Appendix D Comparative Study on Post-Processing
 In this section, we provide a comprehensive analysis regarding the impact of post-processing on the model’s overall calibration quality and UQ performance.
@@
 Therefore, we confirm that using an excessively large number (e.g., 100) for top-$k$ is inadequate for achieving well-calibrated predictions.
-On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive queries are included, a substantial number of negative queries are retained, often outnumbering the positive ones by several times.
+On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive predictions are included, a substantial number of negative predictions are retained, often outnumbering the positive ones by several times.
 Similarly, we compare the image-level uncertainty quantification performance, and the results are shown in Table 4.
@@
 Figure 15 highlight that carefully identifying a reliable subset is crucial for achieving both high object-level calibration quality and effective image-level uncertainty quantification performance.
-We also present this with exemplary visualizations in Figure 8 and in the Appendix.
-(a) Thresholding (0.1)
-(b) Thresholding (0.3)
-(c) Thresholding (0.9)
-(d) Top-$1$
-(e) Top-$20$
-(f) Top-$100$
-(g) NMS ($\tau=0.1$)
-(h) NMS ($\tau=0.5$)
-(i) NMS ($\tau=0.9$)
-Figure 8: Exemplary visualization demonstrating the impact of parameter selection on the final subset of predictions in Cal-DETR for different post-processing schemes.
-Optimal positive and negative predictions are highlighted with green and red boxes, respectively.
-As shown, the top-$k$ and NMS approaches often include too many negative predictions, degrading the calibration quality.
-Confidence thresholding with too low of a threshold faces a similar issue, while too high of a threshold risks omitting positive predictions.
-Therefore, accurately identifying a reliable set of predictions significantly affects the reliability of DETR for downstream applications.
-VIII Conclusion
-The main contribution of our work is an in-depth analysis of the reliability of DETR frameworks.
-In particular, we reveal that DETR’s predictions for a given image exhibit varying calibration quality, highlighting the importance of identifying well-calibrated positive predictions.
-To address this challenge, we introduced a systematic framework that leverages our object-level calibration error metric to discern these positive predictions effectively.
-Furthermore, we proposed a novel uncertainty quantification method for estimating image-level reliability in DETR and conducted comparative studies of various post-processing schemes regarding their impact on DETR’s reliability.
-We hope our efforts expand the scope of DETR applications by enabling more precise and reliable deployment, sparking further research into the area of positive prediction identification.
-Acknowledgments
-The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper.
-This work was supported in part by the MIT Summer Research Program (MSRP), the MIT-Google Program for Computing Innovation, the MIT-Amazon Science Hub, the MIT-IBM Watson AI Lab, and MathWorks.
-Appendix A Dataset
-For our experiments, we used the Cityscapes and Foggy Cityscapes datasets, which each have 500 images of first-person driving footage in realistic environments.
-Foggy Cityscapes has the same base images as Cityscapes, but with fog simulated and added to create a further out of distribution set.
-Since the DETR models were trained on COCO, the Cityscapes and Foggy Cityscapes annotations were converted to correspond to the labels of COCO.
-More specifically, the person, bicycle, car, motorcycle, bus, train, and truck classes were transferred directly.
-In addition, the rider class of Cityscapes and Foggy Cityscapes was mapped to the person class of COCO.
-The other classes present in Cityscapes and Foggy Cityscapes are largely focused on image segmentation, and thus were omitted (e.g.
-building, sky, sidewalk).
-The pre-trained model weights were obtained from their respective official implementations.
-Appendix B Performance Metrics
-Average precision (AP) [17] is the most common accuracy metric for object detectors.
-Before calculating AP, the top-$k$ (e.g., typically $100$) class predictions for a given image are kept and sorted, while the others are discarded.
-Then, in order of this sorting, the predictions are compared against against the ground-truth bounding boxes, and the predicted label is compared with the ground truth label if the IoU is large enough.
-The precision/recall curve is calculated over the sorted predictions, and AP averages the precision across a set of 101 evenly spaced recall thresholds:
-$\text{AP}=\frac{1}{101}\sum_{r\in[0:0.01:1]}p(r)$ (16)
-where $p(r)$ is calculated by the maximum precision at recall level $r$.
-[…]
-$\text{LRP}_{\tau}=\frac{1}{N_{\text{FP}}+N_{\text{FN}}+N_{\text{TP}}}\left(N_{\text{FP}}+N_{\text{FN}}+\sum_{i=1}^{N_{\text{TP}}}\frac{1-\text{IoU}_{i}}{1-\tau}\right)$ (17)
-where $\tau$ is IoU threshold, $i$ denotes the index of each true positive prediction, and IoUi represents the IoU between that prediction and the best-matching ground-truth object.
-LRP is then computed by averaging LRP0.5 and LRP0.75.
-Appendix C Omitted Experimental Results
-This section provides the omitted experimental results.
-Figure 10 extends Figure 5 in Section V-A across different DETR variants and datasets.
-Figure 11 extends Figure 8 in Section VII across different DETR variants and post-processing schemes.
-Appendix D Contrasting with Confounding Negatives for Image-level UQ
-In Table 2, we observe an interesting empirical result: $\mathsf{Conf}^{+}$ using the top 100 predictions (i.e., the fourth row) exhibits a stronger negative correlation compared to $\mathsf{Conf}^{-}$ using the optimal negative predictions (i.e., the second row).
-While a stronger negative correlation might not be desirable when used alone for UQ, this phenomenon can be leveraged within our contrasting framework, similar to how negative predictions are utilized.
-For instance, rather than contrasting the average confidence scores of positive predictions against those of negative (non-positive) predictions, we can instead contrast them against the top 100 predictions.
-However, since positive predictions are included within this top 100 set, an alternative and potentially clearer approach is to contrast the positive predictions specifically against the top 100 non-positive predictions.
-Nevertheless, employing a fixed value of $k=100$ might introduce unintended consequences, particularly depending on the number of ground truth objects present in an image.
-We hypothesize that among non-positive predictions, those with bounding boxes significantly overlapping the positive predictions play a more meaningful role.
-Consequently, negative predictions that have random bounding boxes and very low confidence scores (as illustrated in Figure 1a) can be safely disregarded during the contrastive evaluation.
-To formalize this intuition, we introduce the concept of confounding negatives:
-Definition 3.
-Confounding negative predictions are defined as non-positive predictions whose bounding boxes significantly overlap with positive predictions, specifically exceeding a threshold overlap $\delta$:
-$\hat{\mathcal{D}}_{\theta}^{-\delta}(\bm{x})=\Big{\{}d^{-}\in\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})\mid\max_{d^{+}\in\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})}\mathrm{IoU}(d^{-},d^{+})\geq\delta\Big{\}}.$ (18)
-We conduct an ablation study comparing the original approaches, which use the entire set of negative predictions, against variants that employ different subsets of negatives described above.
-Our findings are summarized in Table 5.
-Overall, we see that restricting attention to negatives with significant overlap can sharpen the negative correlation for $\mathsf{Conf}^{-}$, especially on in-distribution data (COCO), supporting our hypothesis that misleading or “confouding” boxes are more informative for contrastive scoring.
-On the other hand, using entire negatives sometimes proves more robust in near-OOD and OOD scenarios, particularly for $\mathsf{ContrastiveConf}$.
-Table 5: Comparison of correlation results under $\mathsf{Conf}^{-}$ (rows 1–5) and $\mathsf{ContrastiveConf}$ (rows 6–10) when using different negative sets on three datasets: COCO (in-distribution), Cityscapes (near OOD), and Foggy Cityscapes (OOD).
-“Entire Negatives” refers to all non-positive boxes, while “Top-100 Negatives” and “Top-100 Positives” limit the selection to the 100 highest-confidence predictions in each category.
-Confounding negatives restricts the negative set to boxes overlapping positive boxes by at least $\delta$.
-Stronger (more negative) correlation values under $\mathsf{Conf}^{-}$ appear when negatives more closely resemble the positive boxes, and $\mathsf{ContrastiveConf}$ tends to benefit from using such confounding negatives, especially in in-distribution scenarios.
-Methods Negatives COCO (in-distribution) Cityscapes (near OOD) Foggy Cityscapes (OOD)
-UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO UP-DETR D-DETR Cal-DETR DINO
-$\mathsf{Conf}^{-}$ Entire Negatives -0.640 -0.575 -0.571 -0.585 -0.349 -0.382 -0.379 -0.394 -0.256 -0.322 -0.253 -0.283
-Top-100 Positives -0.619 -0.603 -0.581 -0.601 -0.409 -0.374 -0.372 -0.391 -0.262 -0.270 -0.239 -0.229
-Top-100 Negatives -0.640 -0.607 -0.607 -0.620 -0.349 -0.385 -0.415 -0.404 -0.256 -0.315 -0.272 -0.279
-Confounding Negatives ($\delta=0.5$) -0.641 -0.640 -0.574 -0.647 -0.330 -0.351 -0.364 -0.386 -0.250 -0.260 -0.254 -0.261
-Confounding Negatives ($\delta=0.75$) -0.635 -0.628 -0.561 -0.602 -0.385 -0.388 -0.384 -0.390 -0.285 -0.279 -0.269 -0.254
-$\mathsf{ContrastiveConf}$ Entire Negatives 0.656 0.588 0.628 0.613 0.359 0.407 0.441 0.440 0.275 0.350 0.317 0.327
-Top-100 Positives 0.632 0.610 0.604 0.619 0.412 0.379 0.388 0.404 0.270 0.277 0.257 0.247
-Top-100 Negatives 0.656 0.612 0.629 0.633 0.359 0.401 0.439 0.431 0.275 0.332 0.299 0.308
-Confounding Negatives ($\delta=0.5$) 0.653 0.641 0.623 0.654 0.340 0.371 0.434 0.418 0.268 0.284 0.314 0.296
-Confounding Negatives ($\delta=0.75$) 0.647 0.630 0.613 0.619 0.388 0.406 0.448 0.418 0.298 0.301 0.326 0.286
-(a) Object-level Calibration Error
-(b) Object-level Calibration Error
-(Foggy Cityscapes)
-(c) Corr. to Image-level Reliability
-(d) Corr. to Image-level Reliability
-(Foggy Cityscapes)
-Figure 9: A visualization of the difference in calibration between positive and negative predictions on the Cityscapes and Foggy Cityscapes datasets.
-In Figures 9a and 9b, Object-Level Calibration Error (OCE) values of their confidence scores are shown, where a lower score represents better-calibrated predictions.
-In Figures 9c and 9d, Pearson Correlation Coefficient (PCC) values between the ground truth $\mathsf{ImReli}$ and the average confidence scores are shown, where a higher value is better.
+We also present this with exemplary visualizations in Figure 16.
 (a) UP-DETR (COCO)
@@
 Figure 15: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{Conf}^{+}$ and image-level reliability across different DETR models on COCO for different post-processing schemes.
+(a) Thresholding (0.1)
+(b) Thresholding (0.27)
+(c) Thresholding (0.9)
+(d) Top-11
+(e) Top-2020
+(f) Top-$100$
+(g) NMS ($\tau=0.1$)
+(h) NMS ($\tau=0.5$)
+(i) NMS ($\tau=0.9$)
+Figure 16: Exemplary visualization demonstrating the impact of parameter selection on the final subset of predictions in Cal-DETR for different post-processing schemes.
+Optimal positive and negative predictions are highlighted with green and red boxes, respectively.
+As shown, the top-$k$ and NMS approaches could include too many negative predictions, degrading the calibration quality.
+Confidence thresholding with too low of a threshold faces a similar issue, while too high of a threshold risks omitting positive predictions.
+Therefore, accurately identifying a reliable set of predictions significantly affects the reliability of DETR for downstream applications.